### v1

# Dissecting the Hubble tension: Insights from a diverse set of Sound Horizon-free measurements

Abstract
The Hubble tension is commonly framed as a discrepancy between local, late-time measurements favoring km s-1 Mpc-1 and early-time, Sound-Horizon-based measurements favoring km s-1 Mpc-1. We challenge this viewpoint by analyzing 83 Sound-Horizon-independent measurements, categorized into four classes: Distance Ladder measurements using local calibrators; Local One-Step CDM measurements assuming the standard expansion history; Pure Local One-Step measurements independent of shape; and CMB Sound Horizon free measurements using CMB data without the Sound Horizon scale. Our analysis reveals that the 29 Distance Ladder measurements yield km s-1 Mpc-1 (), while the 54 One-Step measurements collectively yield km s-1 Mpc-1 (), a tension exceeding the Planck–SH0ES discrepancy. This tension remains significant at after accounting for correlations. Among One-Step categories, Local One-Step CDM measurements favor the lowest value ( km s-1 Mpc-1, ), Pure Local One-Step yield an intermediate value ( km s-1 Mpc-1, ), and CMB Sound Horizon Free measurements give km s-1 Mpc-1 (). We conclude that the Hubble tension is better characterized as a discrepancy between the Distance Ladder and all other methodologies, rather than an early-vs-late-time split. We also identify a internal tension among One-Step measurements: analyses assuming CDM systematically recover lower values by approximately 3.2 km s-1 Mpc-1 compared to model-independent methods. This suggests either unrecognized systematics/new physics in the Distance Ladder or deviations from CDM in the late-time Universe.
I Introduction
The determination of the Hubble constant has reached a level of precision that has arguably transformed a measurement discrepancy into a crisis for the standard cosmological model. The “Hubble tension” refers to the persistent disagreement between the value inferred from early-Universe probes and that measured from the late-Universe local Distance Ladder.
The standard view frames this tension as a dichotomy between “early” and “late” times. The Planck CMB analysis, assuming CDM, yields [Planck2018], an inference dependent on the Sound Horizon scale at the drag epoch (). The local Distance Ladder, anchored by geometric measurements and extended via Cepheids to SNe Ia, favors [Riess2022], reinforced by JWST observations [Riess2025] and extended network analyses [Casertano2025]. If both are accurate, CDM is incomplete, potentially necessitating new physics [DiValentino2021review, Knox2020, Abdalla:2022yfr, Perivolaropoulos:2021jda, Alestas2020]. Recent DESI hints of dynamical dark energy evolving across the phantom divide [Efstratiou2025, Ozulker:2025ehg, Scherer:2025esj] further complicate this picture, with phantom crossings previously proposed to address both the Hubble tension and the SNe Ia absolute magnitude discrepancy [Alestas2021].
However, this “Early vs. Late” characterization may be oversimplified. It assumes all local measurements favor high values, an assumption challenged by “One-Step” measurements independent of the Cepheid/SN Ia calibration chain. Our recent analysis [Perivolaropoulos2024] demonstrated a stark bifurcation: Distance Ladder measurements clustered around , while independent One-Step methods statistically favored lower values closer to Planck. This suggests the discrepancy may lie between the distance ladder methodology and other probes, rather than between early and late Universe physics [Huang:2024].Indeed, the tension has been shown to have a multidimensional character, with degeneracies between , , and other parameters playing a crucial role [Pedrotti:2024kpn].
In this work, we present an up-to-date compilation of 83 measurements published through late-2025, all independent of the CMB Sound Horizon scale. We introduce a four-category classification based on methodology and model dependence:
1.
Category 1 – Distance Ladder (29 measurements): Traditional three-rung approach using geometric anchors to calibrate standard candles (Cepheids, TRGB, JAGB, Miras) that standardize SNe Ia in the Hubble flow.
2.
Category 2 – Local One-Step CDM (33 measurements): Local measurements bypassing the Distance Ladder but assuming CDM to infer (e.g., Gravitational Wave sirens, Strong Lensing time delays, FRB dispersion measures).
3.
Category 3 – Pure Local One-Step (15 measurements): Model-independent local measurements without assuming a specific shape (e.g., geometric Megamaser distances, Cosmic Chronometers, cosmographic lensing analyses).
4.
Category 4 – CMB Sound Horizon Free (6 measurements): CMB-based measurements decoupled from , using alternative scales such as the horizon at matter-radiation equality ().
The classification follows a systematic decision process (Figure 1): measurements using multi-rung calibration chains are assigned to Category 1; those utilizing CMB data without to Category 4; remaining measurements are assigned to Category 2 or 3 based on whether they assume CDM. The same observational technique can yield measurements in different categories depending on analysis methodology—for example, Megamaser observations with Keplerian models belong to Category 3 [Kuo2013], while those adopting CDM priors belong to Category 2 [Barua2025].
This classification isolates two distinct tensions: the external tension between the Distance Ladder and all other methods, and an internal tension within One-Step measurements examining whether CDM assumptions systematically shift inferred values.A comprehensive overview of observational tensions and potential resolutions involving both systematics and fundamental physics has been presented in the CosmoVerse White Paper [CosmoVerseNetwork:2025alb].
This paper is structured as follows: Section II details our data compilation and measurement tables. Section III presents weighted means, reduced chi-squared values, and tension metrics. Section IV discusses model assumption impacts and bimodality in the distribution. Section V synthesizes these findings and their implications for the Distance Ladder Crisis.
II Data Compilation and Categorization
To dissect the Hubble tension beyond the traditional “Early vs. Late” dichotomy, we have compiled a comprehensive dataset of 83 independent measurements of the Hubble constant, updated through late-2025. This compilation builds upon the catalog presented in our previous work [Perivolaropoulos2024], significantly expanded to include recent results from James Webb Space Telescope (JWST) calibration campaigns, the third generation of gravitational wave catalogs (GWTC-3), and the latest releases from the Dark Energy Spectroscopic Instrument (DESI) [DESI:2025zgx].
II.1 Selection Criteria
Our selection criteria are strictly defined to isolate the impact of the Sound Horizon scale () and the Distance Ladder methodology:
1.
Sound Horizon Independence: We rigorously exclude any measurement that uses the CMB Sound Horizon at the drag epoch as a standard ruler. This excludes standard Baryon Acoustic Oscillation (BAO) inverse Distance Ladder constraints and the standard Planck CDM inference.
2.
Public Availability: All measurements are sourced from peer-reviewed journals or preprints available on arXiv.
3.
Independence of Estimates: Where multiple analyses utilize the exact same dataset (e.g., multiple re-analyses of the same supernova catalog with identical calibration), we select the most recent or representative result to avoid double-counting, unless the methodologies differ significantly in their systematic assumptions.
4.
Methodological Transparency: We include only measurements where the original publication clearly states the cosmological assumptions employed, allowing unambiguous category assignment.
II.2 Classification Methodology
The compiled measurements are classified into four distinct categories based on two primary criteria: (i) reliance on the local Distance Ladder calibration chain, and (ii) dependence on the assumed cosmological model, specifically the shape of the expansion history . The classification follows the decision process outlined in Section I and is determined by the explicit methodological statements in each original publication.
Figure 1 illustrates the classification decision tree schematically. For each measurement, we first assess whether it employs a multi-rung calibration chain (leading to Category 1). If not, we determine whether it utilizes CMB data (leading to Category 4 if Sound-Horizon-free). For the remaining local One-Step measurements, we examine whether the analysis assumes a CDM expansion history (Category 2) or proceeds model-independently (Category 3).
It is important to emphasize that the same observational technique can yield measurements in different categories depending on the analysis approach adopted by the authors. Table 1 provides illustrative examples of how specific techniques are distributed across categories based on their methodological implementation.
| Technique | Category 2 (CDM) | Category 3 (Model-Indep.) |
| Megamasers | Barua (2025): uses CDM | Kuo (2013): Keplerian disk only |
| Strong Lensing | Shajib (2023): CDM lens model | Du (2023): cosmographic approach |
| Gravitational Waves | Abbott (2021): CDM | Sneppen (2023): EPM + kilonova |
| Cosmic Chronom. | Moresco (2023): fit | Zhang (2022): Gaussian Process |
We now describe each category in detail and present the compiled measurements.
II.3 Category 1: Distance Ladder Measurements
This category comprises 29 measurements that employ the traditional multi-rung calibration method. These measurements anchor the absolute magnitude of Type Ia supernovae (SNe Ia) in the Hubble flow using local standard candles calibrated by geometric distances. This is the “Late Universe” standard, relying on the fundamental assumption that the physics of calibrators (Cepheids, TRGB, JAGB, Miras) and SNe Ia are uniform across cosmic time and environment.
The three-rung structure proceeds as follows: (1) geometric anchors (Gaia parallaxes, NGC 4258 Megamaser distance, LMC eclipsing binaries) establish absolute distances to nearby calibrators; (2) these calibrators (Cepheids, TRGB, JAGB, Miras) are observed in galaxies that also host SNe Ia, transferring the distance scale; (3) the calibrated SNe Ia absolute magnitude is applied to supernovae in the Hubble flow () to infer .
Our compilation (Table 2) includes Cepheid-based measurements and recent JWST updates [Riess2025], TRGB-based measurements including CCHP results [Freedman2024], and alternative calibrators such as JAGB stars, Mira variables, Surface Brightness Fluctuations (SBF), and the Tully-Fisher relation. We also include the galaxy pairwise velocity method [Zhang2024], which, while methodologically distinct, relies on calibration from distance ladder measurements.
| Author | Year | Method | ||
| Huang [Huang2020] | 2019 | 73.3 | 4.0 | Mira |
| de Jaeger [deJaeger2020] | 2020 | 75.8 | 5.1 | Cepheids/TRGB + SNe II |
| Khetan [Khetan2021] | 2021 | 70.5 | 4.1 | SBF |
| Freedman [Freedman2021] | 2021 | 69.8 | 1.7 | TRGB (CCHP) |
| Kenworthy [Kenworthy2022] | 2022 | 73.1 | 2.5 | Two-rung ladder |
| Scolnic [Scolnic2024] | 2023 | 73.22 | 2.06 | TRGB (SH0ES calibrators) |
| Li [Li2024b] | 2024 | 74.7 | 3.1 | JAGB + SNe Ia |
| Lee [Lee2024] | 2024 | 67.8 | 2.7 | JAGB + SNe Ia |
| Jensen [Jensen2025] | 2025 | 73.8 | 2.4 | TRGB |
| Riess [Riess2025] | 2025 | 73.49 | 0.93 | Cepheids (JWST) |
| Bhardwaj [Bhardwaj2025] | 2025 | 73.06 | 2.6 | Mira |
| Newman [Newman2025] | 2025 | 75.3 | 2.9 | TRGB + SNe Ia |
We note that the Lee [Lee2024] JAGB measurement ( km s-1 Mpc-1) represents a notable outlier within this category, yielding a value consistent with Planck rather than the typical Distance Ladder result. This discrepancy with other JAGB analyses (e.g., Li [Li2024b]) appears to originate from differences in the JAGB zero-point calibration and photometric selection criteria, highlighting that even within a single indicator class, methodological choices can significantly impact the inferred .
II.4 Category 2: Local One-Step (CDM Assumption)
This category includes 33 measurements that bypass the Distance Ladder (“One-Step”) but infer by explicitly assuming the standard CDM expansion history:
While these methods probe the local Universe (typically ), they rely on this model to map observables—such as time delays, luminosity distances, dispersion measures, or angular diameter distances—to . The key criterion for inclusion in this category is that the original analysis explicitly assumes the functional form of Eq. (1), often with fixed to a fiducial value (typically ) or marginalized over a narrow prior.
This category includes:
Gravitational Wave Standard Sirens: Both bright sirens (with EM counterparts, e.g., GW170817 [Palmese2023]) and dark sirens (using galaxy catalogs for redshift inference [Abbott2021, Beirnaert2025]), where the luminosity distance is computed assuming CDM.
Strong Lensing Time Delays: Analyses where the cosmological model enters both the time-delay distance calculation and often the lens mass modeling [Shajib2023, Kelly2023, Birrer2025].
Fast Radio Bursts: Measurements using the dispersion measure–redshift relation interpreted under CDM assumptions for the intergalactic medium [Wu2022, PiratovMoreno2025].
| Author | Year | Method | ||
| Wu [Wu2022] | 2021 | 68.8 | 4.7 | FRBs |
| Shajib [Shajib2023] | 2023 | 77.1 | 7.2 | Lensing (TDCOSMO) |
| Napier [Napier2023] | 2023 | 74.1 | 8.0 | Cluster Lenses |
| Kelly [Kelly2023] | 2023 | 66.6 | 3.7 | Lensed SN Refsdal |
| Liu [liu2023h] | 2023 | 59.1 | 3.6 | Cluster Lensed Quasar |
| Martinez [Martinez2023] | 2023 | 74.0 | 11.2 | Time-Delay Cosmography |
| Pascale [Pascale2024] | 2024 | 71.8 | 8.7 | Lensed SN H0pe (JWST) |
| Yang [Yang2024] | 2024 | 74.0 | 7.4 | FRBs w/ Scattering |
| Piratova [PiratovMoreno2025] | 2025 | 65.1 | 7.4 | FRBs (MLE) |
| Beirnaert [Beirnaert2025] | 2025 | 79.0 | 9.0 | GW Dark Sirens |
| Birrer [Birrer2025] | 2025 | 71.6 | 3.6 | TDCOSMO Strong Lensing |
| Pierel [Pierel2025] | 2025 | 66.9 | 9.8 | SN Encore Time Delays |
II.5 Category 3: Pure Local One-Step (Model Independent)
This category represents the cleanest test of the Distance Ladder, as it comprises 15 measurements that determine geometrically or via direct measurement of the expansion rate, without assuming a specific functional form for . These measurements are either:
Purely geometric: Relying on Keplerian dynamics or light-travel-time arguments that are independent of the global cosmological model (e.g., Megamaser disk modeling [Kuo2013, Gao2015]).
Direct measurements: Cosmic chronometers measure directly from the differential aging of massive, passively evolving galaxies, providing model-independent determinations of the expansion rate at various redshifts [Moresco2023]. To extract , two main approaches have been employed: (i) Gaussian Process reconstruction of combined with a cosmographic expansion of the luminosity distance in the Hubble flow (), (ii) using cosmic chronometers to calibrate the H ii galaxies Hubble diagram, which in turn anchors Type Ia supernovae at low redshift via Hubble’s law [Zhang2022]. Both methods avoid assumptions about the underlying cosmological model beyond the FLRW metric.
Cosmographic approaches: Analyses that expand or in Taylor series around , avoiding assumptions about the energy content of the Universe [Du2023].
Expanding photosphere methods: Geometric distance determinations from supernovae using photospheric velocity and angular size evolution [Sneppen2023, Vogl2024].
The key distinction from Category 2 is that these analyses do not assume the validity of Eq. (1). For Megamaser measurements, we include in this category only those analyses that rely purely on Keplerian disk modeling without adopting CDM-based peculiar velocity corrections or external distance priors tied to the Distance Ladder.
| Author | Year | Method | ||
| Kuo [Kuo2013] | 2013 | 68.0 | 9.0 | Megamasers (MCP) |
| Gao [Gao2015] | 2015 | 66.0 | 6.0 | Megamasers |
| Zhang [Zhang2022] | 2022 | 65.9 | 3.0 | CC + HII |
| Du [Du2023] | 2023 | 71.5 | 3.8 | TD Lenses + GRBs |
| Sneppen [Sneppen2023] | 2023 | 67.0 | 3.6 | EPM + Kilonovae |
| Li [Li2024a] | 2024 | 66.3 | 3.7 | Lensing Bias Correction |
| Vogl [Vogl2024] | 2024 | 74.9 | 2.7 | SNe II (Tailored EPM) |
| Song [Song2025] | 2025 | 70.4 | 6.9 | GW + SGL Calibration |
| Favale [Favale2025] | 2025 | 68.8 | 3.0 | Cosmic Chronometers |
We highlight that this category exhibits significant internal heterogeneity. Notably, the Vogl [Vogl2024] tailored EPM measurement ( km s-1 Mpc-1) strongly favor the Distance Ladder value, while most cosmic chronometer measurements favor lower values. This heterogeneity will be discussed further in Section IV.
II.6 Category 4: CMB Sound Horizon Free
The final category includes 6 measurements that utilize early-Universe data (primarily CMB) but specifically avoid the Sound Horizon as a standard ruler. Instead, these methods rely on alternative physical scales that are calibrated independently of pre-recombination acoustic physics:
Matter-radiation equality scale (): The comoving wavenumber corresponding to the horizon size at matter-radiation equality provides a standard ruler that depends on and but is independent of [Philcox2022, Bahr-Kalus2025].
BAO without priors: Analyses that use BAO peak positions but treat as a free parameter, constraining through the shape of rather than absolute calibration [Pogosian2024].
Sunyaev-Zel’dovich (SZ) effect: Combining SZ decrements with X-ray observations of galaxy clusters to infer angular diameter distances independent of [Reese2003, Colaco2023].
| Author | Year | Method | ||
| Reese [Reese2003] | 2003 | 61.0 | 18.0 | SZ Effect |
| Colaco [Colaco2023] | 2023 | 67.2 | 6.1 | SZ + X-ray |
| Pogosian [Pogosian2024] | 2024 | 69.5 | 0.94 | BAO-DESI (No ) |
| Pogosian [Pogosian2024] | 2024 | 68.0 | 0.94 | BAO-pre DESI (No ) |
| Bahr-Kalus [Bahr-Kalus2025] | 2025 | 65.2 | 5.6 | DESI Turnover Scale |
| Krolewski [Krolewski2025] | 2025 | 69.0 | 2.5 | DESI + CMB (No ) |
We note that this category includes two entries from Pogosian et al. [Pogosian2024], corresponding to analyses using DESI and pre-DESI BAO data respectively. While these share a common methodology, they utilize substantially different datasets and thus provide partially independent constraints. However, the very low reduced chi-squared () for this category suggests that uncertainties may be conservative or that some correlation exists between measurements. We address this in the correlation analysis of Appendix C.
II.7 Summary of Classification
Table 6 summarizes the four categories and their key characteristics. The complete measurement tables with extended methodological information are provided in Appendix A.
| Cat. | Description | Uses Ladder? | Assumes CDM? | |
| 1 | Distance Ladder | 29 | Yes | No |
| 2 | Local One-Step (CDM) | 33 | No | Yes |
| 3 | Pure Local One-Step | 15 | No | No |
| 4 | CMB Sound Horizon Free | 6 | No | Yes∗ |
∗Category 4 assumes CDM for shape but avoids the Sound Horizon .
III Statistical Analysis and Results
In this section, we present a detailed statistical analysis of the measurements classified in Section II. We utilize the inverse-variance weighted mean to determine the central value for each category and the reduced chi-squared statistic () to assess internal consistency. We explicitly quantify the tension between the Distance Ladder (Category 1) and the various subsets of Sound-Horizon-independent measurements, concluding with a comparison against the combined ensemble of all One-Step methods. We also compare our results with our previous analysis [Perivolaropoulos2024] to assess the evolution of the tension landscape.
III.1 Methodology
For a dataset of independent measurements , the weighted mean and its standard error are calculated as:
where the weights are given by . This estimator minimizes the variance for Gaussian-distributed errors.
To evaluate whether the variation within a category is consistent with the reported uncertainties, we compute the reduced chi-squared:
The interpretation of provides important diagnostic information:
: Good internal consistency; the scatter is consistent with reported uncertainties.
: Underestimated errors, unrecognized systematics, or genuine physical scatter.
: Overestimated errors, unaccounted correlations between measurements, or possible publication bias toward a consensus value.
The statistical tension between two independent estimates and is quantified in terms of standard deviations () as:
We note that this tension metric assumes statistical independence between the two samples. In practice, correlations exist within each category due to shared calibration data, common analysis frameworks, and overlapping samples. We address the impact of these correlations in Appendix C, where we show that accounting for realistic correlation structures reduces the tension significance but does not qualitatively alter our conclusions.
III.2 Weighted Mean Statistics by Category
The statistical results for the four categories defined in Section II are summarized in Table 7. For comparison, we also include the results from our previous analysis [Perivolaropoulos2024], which employed a binary classification scheme (Distance Ladder vs. One-Step).
| Category | Tension | |||
| 1. Distance Ladder | 29 | 0.74 | — | |
| 2. Local One-Step (CDM) | 33 | 0.85 | ||
| 3. Pure Local One-Step | 15 | 0.78 | ||
| 4. CMB Sound Horizon Free | 6 | 0.36 | ||
| Combined (2+3+4) | 54 | 0.85 | ||
| Corr.-adjusted (2+3+4) | 54 | — | ||
| Comparison with Perivolaropoulos (2024) [Perivolaropoulos2024]: | ||||
| 2024 Distance Ladder | 20 | 0.51 | — | |
| 2024 One-Step | 33 | 0.95 | ||
To visualize the distinct behaviors of these categories, we present the Gaussian probability distributions in Figure 2. This figure illustrates the probability density functions for each of the four subgroups, offering a clear visual representation of the tensions quantified in Table 7.
III.2.1 Category 1: Distance Ladder ()
The 29 measurements comprising the Distance Ladder category yield a weighted mean of km s-1 Mpc-1. As shown in Figure 2, this category forms a sharp, narrow peak that is clearly separated from the other distributions.
The reduced chi-squared indicates good internal consistency, with the scatter being slightly smaller than expected from the reported uncertainties. This value is notably higher than our 2024 analysis (), suggesting that the expanded dataset—which now includes measurements with greater methodological diversity (e.g., JAGB, Miras, pairwise velocities)—exhibits more realistic scatter. Nevertheless, may indicate that: (i) reported uncertainties are somewhat conservative, (ii) positive correlations exist between measurements sharing common calibration anchors, or (iii) there is a tendency for published results to cluster around previously established values.
The high precision of this combined estimate is driven by the recent JWST Cepheid measurements [Riess2025] and compatible results from other calibrators including TRGB [Jensen2025], Miras [Bhardwaj2025], and SBF [Blakeslee2021]. We note that the Lee et al. [Lee2024] JAGB measurement () represents a outlier within this category, as discussed in Section II.
III.2.2 Category 2: Local One-Step CDM ()
The 33 measurements in this category, which utilize local data but assume the CDM expansion history, yield the lowest aggregate value of km s-1 Mpc-1. This value is fully consistent with, and in fact slightly lower than, the Planck CMB inference ( km s-1 Mpc-1), despite being derived entirely from local observations.
The reduced chi-squared indicates internal consistency among these diverse methods (Strong Lensing, Gravitational Wave sirens, FRBs). This near-unity value suggests that the reported uncertainties accurately capture the true scatter and that no significant unrecognized systematics are present within this category.
In Figure 2, the distribution of category 2 appears at the lowest end of the parameter space, almost coincident with the Planck CMB value. This remarkable alignment between local CDM-dependent measurements and early-universe physics—achieved without using the Sound Horizon—is one of the central findings of this work.
III.2.3 Category 3: Pure Local One-Step ()
The 15 model-independent measurements (Megamasers, Cosmic Chronometers, EPM) yield an intermediate value of km s-1 Mpc-1. Crucially, the removal of the CDM shape assumption results in a central value shift of km s-1 Mpc-1 relative to Category 2. This shift is visible in Figure 2 as a broader distribution positioned between the CDM-dependent results and the Distance Ladder.
The reduced chi-squared is lower than unity, indicating either conservative error estimates or the presence of correlations. This category exhibits significant internal heterogeneity: the Vogl et al. [Vogl2024] tailored EPM measurement () strongly favors the Distance Ladder value, while most Cosmic Chronometer measurements favor values closer to Planck. This heterogeneity suggests that Category 3 may itself contain distinct subpopulations, a possibility we explore further in Section IV.
III.2.4 Category 4: CMB Sound Horizon Free ()
The 6 measurements utilizing CMB data without the sound horizon yield km s-1 Mpc-1. This confirms that early-universe constraints on matter density (via the equality scale ) favor a low regardless of the calibration [Pogosian2024].
The very low reduced chi-squared warrants careful interpretation. With only measurements, statistical fluctuations in are expected to be large (the standard deviation of scales as for ). Nevertheless, this low value likely reflects: (i) the dominance of the two high-precision Pogosian et al. [Pogosian2024] measurements, which share a common methodology and partially overlapping datasets; and (ii) the fact that all measurements in this category are ultimately constrained by similar early-universe physics. In Figure 2, this distribution overlaps significantly with Category 2, reinforcing the consistency between early-universe physics and local data when CDM is assumed.
III.3 Comparison: Distance Ladder vs. Combined One-Step
To test the robustness of the “Early vs. Late” narrative, we combine all 54 Sound-Horizon-independent measurements from Categories 2, 3, and 4 into a single “All One-Step” ensemble. This combined sample represents the entirety of the evidence against the Distance Ladder that does not rely on the standard ruler.
The weighted mean of this combined sample is:
This distribution is shown as the black curve in Figure 2 and the red curve in Figure 3, contrasting sharply with the Distance Ladder (blue curve). Figure 3 displays the probability density functions (PDFs) for the two measurement categories. Each individual measurement is represented as a Gaussian distribution, and the combined PDF is constructed as an inverse-variance weighted sum:
where is the normal distribution:
The final normalized PDF is obtained by dividing by the total area:
The two curves exhibit significant overlap despite the tension between their weighted means. This occurs because the image reflects the spread of individual measurement uncertainties (typically 2–10 ), whereas the uncertainty on the weighted mean is much smaller. Consequently, the PDF width is governed by the individual values, while the tension is calculated using the combined of Eq. (2).
This distinction explains why the curves overlap visually even when the weighted means differ by several standard deviations. The tension between the Distance Ladder and the Combined One-Step sample is:
Combining the uncertainties in quadrature ( km s-1 Mpc-1), the significance of the discrepancy is:
This tension exceeds the standard Planck vs. SH0ES discrepancy (). It suggests that the Distance Ladder is in conflict not just with the CMB Sound Horizon, but with the aggregate of essentially all other cosmological probes that are independent of .
III.3.1 Comparison with Previous Results
Table 7 includes a comparison with our 2024 analysis [Perivolaropoulos2024]. Despite the significant expansion of the dataset (from 53 to 83 measurements) and the refinement of the classification scheme (from 2 to 4 categories), the core results remain consistent:
The Distance Ladder weighted mean has shifted slightly from to km s-1 Mpc-1, with improved precision due to the inclusion of JWST measurements.
The One-Step weighted mean has shifted from to km s-1 Mpc-1, remaining consistent within uncertainties.
The tension has increased from to , primarily due to reduced uncertainties rather than a change in central values.
The stability of these results across independent analyses with different sample compositions strengthens confidence in the robustness of the “Distance Ladder vs. The Rest” characterization of the Hubble tension.
III.3.2 Impact of Correlations
The tension quoted above assumes statistical independence between measurements. As detailed in Appendix C, accounting for realistic correlations within each category increases the combined uncertainties. Using correlation coefficients estimated from shared calibration data and common methodology (– for Distance Ladder; – for One-Step), we obtain:
The correlation-adjusted tension is:
While reduced from , this tension remains highly significant and robust across the range of plausible correlation assumptions (– for reasonable correlation models).
III.4 Internal Tension: Model Dependence
Beyond the external tension with the Distance Ladder, we identify a notable internal tension among the One-Step measurements based on model assumptions. Comparing Category 2 (Local One-Step CDM) with Category 3 (Pure Local One-Step), we find:
The statistical significance is:
This internal tension, while not definitive, is suggestive of a systematic effect: imposing the CDM model on local data appears to pull the inferred downward by km s-1 Mpc-1. This shift accounts for more than half of the total Hubble tension ( km s-1 Mpc-1 between Distance Ladder and Planck).
We caution that this internal tension should be interpreted carefully for several reasons:
1.
Look-elsewhere effect: We are comparing multiple categories, which increases the probability of finding a discrepancy by chance.
2.
Category heterogeneity: As noted above, Category 3 contains measurements spanning a wide range of values (from to km s-1 Mpc-1).
3.
Sample size: Category 3 contains only 15 measurements, making the weighted mean sensitive to individual high-precision entries.
III.5 Summary of Statistical Results
The key statistical findings of this section are:
1.
The Distance Ladder yields km s-1 Mpc-1, consistent with SH0ES and JWST measurements.
2.
The combined One-Step measurements yield km s-1 Mpc-1, consistent with Planck.
3.
The tension between these two ensembles is (or after correlation adjustment).
4.
Local measurements assuming CDM (Category 2) yield systematically lower than model-independent local measurements (Category 3), with a internal tension.
5.
These results are consistent with, and strengthen, the conclusions of our 2024 analysis.
The implications of these findings—particularly the role of model assumptions and the apparent bimodality in the distribution—are explored in the following section.
IV The Impact of Model Assumptions
The results of Section III reveal a striking feature of the Hubble tension landscape: the discrepancy is not merely between the Distance Ladder and the Early Universe, but also manifests as an internal tension within the domain of One-Step measurements. Specifically, the statistical compatibility of a measurement with the Planck value depends critically on whether the standard CDM expansion history is assumed during the inference process. In this section, we dissect this model dependence, examine the critical role of specific high-precision measurements, and explore the physical implications of the observed bimodality in the distribution.
IV.1 Implicit Bias of the CDM Expansion History
The most significant finding of our four-category classification is the tension between local One-Step measurements that assume a CDM cosmology (Category 2: km s-1 Mpc-1) and those that do not (Category 3: km s-1 Mpc-1).
Measurements in Category 2, such as Gravitational Wave standard sirens [Abbott2021] and standard analyses of Strong Lensing time delays [Shajib2023], typically span the redshift range . To extract from the observed luminosity distances or time delays, these methods must integrate the Hubble parameter :
By fixing the functional form of to that of flat CDM (Eq. 1), often with fixed to or marginalized over a narrow prior, these analyses effectively impose a rigid shape constraint on the expansion history. The inferred is then the normalization that best fits the data given this assumed shape.
Our analysis shows that this CDM shape constraint systematically prefers lower values ( km s-1 Mpc-1), aligning almost perfectly with the Planck CMB inference [Planck2018]. This demonstrates a remarkable self-consistency of the CDM model: if one assumes the CDM background evolution holds true at late times, local data (excluding the Distance Ladder) yields results consistent with early-universe physics, even without using the Sound Horizon.
However, when this shape constraint is released (Category 3), the inferred rises by km s-1 Mpc-1 to km s-1 Mpc-1. This shift has two possible interpretations:
1.
CDM is correct, and Category 3 has larger systematics: Model-independent methods may suffer from unrecognized systematic errors that bias high. The lack of a shape constraint allows more freedom for systematics to propagate into the final result.
2.
The true deviates from CDM: If the actual expansion history differs from the CDM prediction at (e.g., due to a late-time gravitational transition, dynamical dark energy, or modified gravity [Perivolaropoulos2024]), forcing a CDM fit would bias the inferred downward to compensate for the shape mismatch.
A related possibility is that the tension reflects not a global modification of , but rather a localized change in the physics governing distance calibrators. Refs [Desmond:2019ygn, Desmond:2020wep, Marra2021, Ruchika2025, Perivolaropoulos:2025gzo] investigated whether a local variation in fundamental physics—such as a change in the fine-structure constant or gravitational strength within the calibration volume—could reconcile the SH0ES measurement with other probes without requiring modifications to the global expansion history. The current data do not definitively distinguish between these scenarios.
IV.2 The Vogl et al. Tailored EPM: A Critical Test Case
Among the Category 3 measurements, the Vogl et al. [Vogl2024] tailored Expanding Photosphere Method (EPM) result deserves particular attention. This measurement yields km s-1 Mpc-1 (with systematic uncertainties — by the authors own estimate comparable to the statistical errors — taken into account) is fully consistent with the Distance Ladder and in tension with the Planck value—yet it is completely independent of both the Distance Ladder calibration chain and cosmological model assumptions.
The EPM determines distances to Type II supernovae geometrically by combining measurements of the photospheric angular size (from flux and temperature) with the photospheric velocity (from spectral line Doppler shifts). The “tailored” approach of Vogl et al. advances this technique by using detailed non-local thermodynamic equilibrium (NLTE) radiative transfer models customized to individual supernovae, rather than the empirical dilution factors traditionally employed. This reduces model-dependent systematics and achieves significantly smaller distance uncertainties.
The significance of this measurement for the Hubble tension debate is threefold:
1.
Independence from SNe Ia physics: Unlike the Distance Ladder, which relies on the standardizability of SNe Ia, the tailored EPM uses SNe II—an entirely different supernova class with distinct explosion physics.
2.
Independence from calibrator physics: The EPM does not require Cepheids, TRGB, or any other intermediate calibrator. It provides a direct geometric distance to the supernova.
3.
Independence from CDM: The method measures distances without integrating , making it insensitive to the assumed cosmological model.
If the Distance Ladder’s high were solely due to systematic errors in Cepheid photometry, metallicity corrections, or SNe Ia standardization, one would not expect an independent method like the tailored EPM to yield a consistent result. The agreement between Vogl et al. and the Distance Ladder suggests that either: (a) both methods share a common, as-yet-unidentified systematic, or (b) the high value reflects the true local expansion rate.
However, we note important caveats. The tailored EPM relies on assumptions about supernova explosion physics, ejecta composition, and atmospheric structure. The current sample is limited to a relatively small number of well-observed SNe II. Further validation with larger samples and independent atmosphere codes will be essential to confirm this result.
IV.3 The Megamaser Perspective
A similar pattern emerges from Megamaser measurements. The Barua et al. [Barua2025] frequentist analysis of the Megamaser Cosmology Project (MCP) sample yields km s-1 Mpc-1, again consistent with the Distance Ladder.
Megamaser distances are determined geometrically from the Keplerian dynamics of circumnuclear water maser disks, providing angular diameter distances that are in principle completely independent of stellar physics. However, Megamaser measurements are sensitive to the treatment of peculiar velocities, which become significant at the low redshifts () where most masers hosts reside.
In our classification, we assigned the Barua et al [Barua2025] measurement to Category 2 because it relies primarily on Keplerian disk modeling without adopting CDM-based peculiar velocity priors but relies on the CDM prediction for the angular diameter distance. Indeed, when is left as a free parameter rather than fixed to the Planck value, the best-fit drops to km s-1 Mpc-1 with uncertainties approximately five times larger, highlighting the model dependence of the headline result. In contrast, earlier analyses that adopted peculiar velocity corrections from the SH0ES flow model from Pesce et al. [Pesce2020] would be more appropriately classified as having some distance ladder dependence.
The earlier Megamaser measurements by Kuo et al. [Kuo2013] ( km s-1 Mpc-1) and Gao et al. [Gao2015] ( km s-1 Mpc-1) yielded lower values, though with larger uncertainties. The evolution toward higher in recent Megamaser analyses may reflect improved disk modeling, better peculiar velocity treatments, or statistical fluctuations in small samples.
IV.4 Bimodality in the One-Step Distribution
Although Figure 3 does not exhibit an obvious bimodal structure—the distribution of One-Step measurements appears reasonably concentrated around a single peak—a more detailed examination reveals a subtle but significant effect.
Upon closer inspection, we found that removing even two high-precision measurements (the Pogosian et al. [Pogosian2024] DESI and pre-DESI results) substantially alters the overall shape of the distribution. These measurements, with uncertainties of km s-1 Mpc-1, are among the most precise in our compilation. Due to their small uncertainties, they exert disproportionate influence on the inverse-variance weighted distribution, effectively pulling the combined probability density upward at their location (– km s-1 Mpc-1).
When these high-weight measurements are excluded, the inverse-variance weighted sum of Gaussians transforms into a distinctly bimodal curve shown in Figure 4. This suggests that the apparent unimodality in the raw distribution may be an artifact of a small number of highly precise measurements rather than a reflection of the true underlying structure.
We emphasize that this analysis should be interpreted with caution. Selectively removing measurements to reveal bimodality could be seen as a form of data manipulation. However, our purpose here is not to argue that the Pogosian et al. measurements are incorrect, but rather to illustrate that the statistical structure of the distribution is sensitive to a small number of high-precision anchors. A more rigorous approach would involve mixture model analysis (e.g., Gaussian Mixture Models) to formally test for bimodality, which we leave for future work.
IV.5 Physical Interpretation of the Bimodality
If the bimodality is real and not an artifact of small-number statistics, it suggests that the “Hubble Tension” is not characterized by a single discrepant value but rather by a bifurcation of results based on methodology. The two peaks can be characterized as follows:
1.
The Low- Peak ( km s-1 Mpc-1): Populated predominantly by:
Category 2 measurements that assume CDM (FRBs, GW sirens, most Strong Lensing)
Category 4 measurements anchored to early-universe physics (-based BAO)
Cosmic Chronometer measurements in Category 3
These methods either explicitly assume CDM or rely on physics that is ultimately calibrated by the same underlying cosmological model.
2.
The High- Peak ( km s-1 Mpc-1): Populated by:
Specific measurements: Megamasers [Barua2025] and tailored EPM [Vogl2024]
These methods are characterized by their reliance on local geometric or kinematic physics.
This bifurcation pattern suggests a reformulation of the Hubble tension: the discrepancy may fundamentally be between model-dependent methods (which yield low ) and local astrophysics dependent or model-independent methods (which yield high ), rather than between early and late times.
IV.6 The Role of High-Precision Anchors
The tension landscape is significantly shaped by a small number of high-precision measurements that serve as “anchors” for the two solutions:
High- anchors:
Riess et al. [Riess2025] JWST Cepheids: km s-1 Mpc-1
Zhang et al. [Zhang2024] pairwise velocities: km s-1 Mpc-1
Vogl et al. [Vogl2024] tailored EPM: km s-1 Mpc-1
Low- anchors:
Pogosian et al. [Pogosian2024] DESI BAO (no ): km s-1 Mpc-1
Piratova-Moreno et al. [PiratovMoreno2025] FRBs: km s-1 Mpc-1
Kelly et al. [Kelly2023] SN Refsdal: km s-1 Mpc-1
The weighted means of each category are heavily influenced by these anchor measurements. Consequently, resolving the Hubble tension requires either: (a) identifying systematic errors in one or more anchor measurements, or (b) explaining why different methodologies yield systematically different results.
IV.7 Category 3 as the Critical Adjudication Ground
The “Pure Local One-Step” category (Category 3) serves as the critical test for distinguishing between the two interpretations of the Hubble tension. Currently, this category yields an intermediate value ( km s-1 Mpc-1), but it is internally heterogeneous:
High- subset: Vogl EPM (), Palmese GW170817 (75.46±5.4)
Low- subset: Zhang CC+HII (), Moresco CC (), Li lensing ()
The future evolution of Category 3 will be decisive:
1.
If Category 3 converges toward Category 1 (Distance Ladder): This would strongly imply that the CDM model assumption used in Category 2 is the source of the bias. Such convergence would suggest that the true late-time expansion history deviates from CDM, requiring new physics (e.g., dynamical dark energy, modified gravity, or a late-time gravitational transition)such as the CDM model [akarsu2024, Akarsu2025]
2.
If Category 3 converges toward Category 2 (CDM): This would suggest that the Distance Ladder contains a unique, unrecognized systematic error affecting all its calibrators (Cepheids, TRGB, JAGB, Miras) in a similar way. Potential culprits include metallicity effects, dust extinction corrections, or environmental dependencies of SNe Ia.
Currently, the intermediate position of Category 3, combined with its internal heterogeneity, does not clearly favor either scenario. However, the existence of multiple independent model-independent measurements (EPM, Megamasers) that support the high- solution is suggestive.
IV.8 Summary: The Model-Dependence Effect
The key findings of this section can be summarized as follows:
1.
Imposing CDM assumptions on local kinematic data reduces the inferred by approximately km s-1 Mpc-1 (from Category 3’s to Category 2’s ).
2.
This shift accounts for more than half of the total Hubble tension ( km s-1 Mpc-1 between Distance Ladder and Planck).
3.
The One-Step distribution shows evidence for bimodality when high-precision anchoring measurements are removed, with peaks corresponding to model-dependent () and model-independent () methods.
4.
Specific model-independent measurements (Vogl EPM) yield high values consistent with the Distance Ladder, suggesting the high value may reflect true local physics rather than distance ladder systematics.
5.
The future convergence or divergence of Category 3 measurements will be critical for determining whether the tension originates from CDM model bias or distance ladder systematics.
These findings support a recharacterization of the Hubble tension as primarily a “Distance Ladder vs. The Rest” crisis, with a secondary “Model-Dependent vs. Model-Independent” tension that may hold the key to its ultimate resolution.
V Discussion and Conclusions
The “Hubble Tension” has long been framed as a cosmological crisis pitting the Early Universe (calibrated by the sound horizon) against the Late Universe (measured by the distance ladder). In this work, we have dissected this narrative by analyzing 83 measurements of that are strictly independent of the CMB Sound Horizon scale. Our findings suggest that the crisis is not necessarily a temporal one, but rather a methodological one. In this concluding section, we synthesize our results, discuss their implications for proposed solutions to the tension, and outline the observations needed to achieve resolution.
V.1 The “Distance Ladder vs. The Rest” Crisis
Our primary conclusion is that the Hubble tension is most accurately characterized as a discrepancy between the Distance Ladder methodology and essentially all other cosmological probes, regardless of whether those probes utilize early- or late-time data.
As detailed in Section III, the Distance Ladder (Category 1) yields a robust high value of km s-1 Mpc-1. In stark contrast, when we combine all 54 independent measurements from Categories 2, 3, and 4—comprising Gravitational Waves, Strong Lensing, Cosmic Chronometers, Megamasers, FRBs, and sound-horizon-free CMB data—we obtain a combined value of km s-1 Mpc-1.
The resulting tension ( after correlation adjustment) between the Distance Ladder and this “All One-Step” ensemble exceeds the standard Planck vs. SH0ES discrepancy (). This indicates that the high value of is a unique feature of the three-rung calibration chain (involving Cepheids/TRGB/Miras SNe Ia) and is not supported by the broader consensus of Sound-Horizon-free cosmology.
This conclusion is robust across our analyses:
It persists when correlations within categories are accounted for (Section III.3).
It is consistent with our 2024 analysis [Perivolaropoulos2024], despite the significant expansion of the dataset.
It holds regardless of which specific measurements are included or excluded (with the exception of complete removal of entire methodological classes).
V.2 Local Physics Does Not Imply High
A pervasive misconception in the field is that “local” measurements universally favor high values, thereby pointing to a breakdown of CDM at late times. Our analysis decisively refutes this assumption.
We found that 33 measurements utilizing exclusively local data () but assuming the validity of the CDM expansion history (Category 2) yield a weighted mean of km s-1 Mpc-1. This value is in perfect agreement with the Planck 2018 result ( km s-1 Mpc-1) [Planck2018]. Local probes such as Gravitational Wave standard sirens [Abbott2021], Strong Lensing time delays [Shajib2023, Kelly2023, Birrer2025], and FRBs [Wu2022, PiratovMoreno2025] do not intrinsically favor high . When analyzed under the standard model framework, local data supports the early-universe inference.
Furthermore, Category 4 demonstrates that even without the Sound Horizon , CMB data constraints (driven by the matter-radiation equality scale ) favor a low km s-1 Mpc-1 [Pogosian2024, Bahr-Kalus2025].
Thus, neither “Local” nor “Sound Horizon free” is synonymous with “High .” The high value may be specific to the Distance Ladder methodology.
V.3 Internal Tension and the Shape of
While the external tension with the Distance Ladder dominates the landscape, we have identified a secondary internal tension among the One-Step measurements that may hold the key to understanding the underlying physics.
Measurements that assume the CDM shape of (Category 2) yield km s-1 Mpc-1. However, measurements that are free from this shape assumption (Category 3) yield a higher intermediate value of km s-1 Mpc-1. This km s-1 Mpc-1 shift accounts for more than half of the total Hubble tension.
This systematic offset has two possible interpretations:
1.
CDM is correct: Model-independent methods (Category 3) suffer from larger uncontrolled systematics that bias high.
2.
CDM is incomplete: The true expansion history deviates from the CDM prediction at , and forcing a CDM fit biases low.
The fact that specific model-independent measurements—particularly the Vogl et al. [Vogl2024] tailored EPM ()— yield values consistent with the Distance Ladder is intriguing. This method is independent of both the Cepheid/SNe Ia calibration chain and cosmological model assumptions, yet it supports the high- solution. This pattern is more naturally explained by interpretation (2) than by a conspiracy of unrelated systematics.
The persistence of this internal tension reinforces earlier findings that the CDM model may face challenges beyond the Hubble tension alone. A comprehensive review of these challenges [Perivolaropoulos2022review] identifies correlated anomalies that may point toward a common origin in modified late-time physics.
V.4 Implications for Proposed Solutions
Our findings have significant implications for the major classes of proposed solutions to the Hubble tension:
V.4.1 Early Dark Energy and Sound Horizon Modifications
Models that resolve the tension by reducing the Sound Horizon at recombination (e.g., Early Dark Energy [Knox2020, Mirpoorian2025], extra relativistic species, early modified gravity) predict that Sound-Horizon-independent measurements should align with the Distance Ladder. Our analysis shows the opposite: Sound-Horizon-free measurements (Categories 2, 3, and 4) predominantly align with Planck, not with SH0ES.
This presents a significant challenge to early-universe solutions. If EDE or similar models were correct, we would expect:
Megamaser distances to yield km s-1 Mpc-1 ✓(partially observed).
GW standard sirens to yield km s-1 Mpc-1 ✗ (observed: )
Strong lensing to yield km s-1 Mpc-1 ✗ (observed: –)
Sound-horizon-free BAO to yield km s-1 Mpc-1 ✗ (observed: )
The failure of most Sound-Horizon-independent probes to support the high- solution disfavors pure early-universe modifications as the complete explanation.
V.4.2 Late-Time Modifications
Models that modify the expansion history at late times (e.g., dynamical dark energy, late-time gravitational transitions, interacting dark energy) predict that the inferred should depend on whether CDM is assumed in the analysis. Our observation of the Category 2 vs. Category 3 offset ( km s-1 Mpc-1) is qualitatively consistent with this prediction.
However, late-time solutions face their own challenges:
They must explain why the shift is only km s-1 Mpc-1, not the full km s-1 Mpc-1 tension.
They must be consistent with other late-time probes (SNe Ia Hubble diagrams, BAO, growth of structure) that show no dramatic deviation from CDM.
Recent DESI results [DESI2024, Giare:2025pzu, Gao:2024ily] suggesting evolving dark energy (, ) may provide partial support, but the statistical significance remains modest and the robustness across different parameterizations has been questioned [Colgain:2024mtg]. Theoretical frameworks for such behavior include models where the effective gravitational constant undergoes a rapid transition at late times, which can arise naturally in scalar-tensor theories. Recent re-examinations of such modified local physics scenarios using SH0ES data have refined the constraints on these models. A particularly intriguing perspective recasts the Hubble tension as an crisis. Theodoropoulos & Perivolaropoulos demonstrated that if the intrinsic luminosity of SNe Ia in the Hubble flow differs systematically from that in calibrator galaxies, the inferred would be biased without requiring modifications to fundamental cosmology [Marra2021, Perivolaropoulos:2025gzo, Theodoropoulos2021] .
V.4.3 Distance Ladder Systematics
The isolation of the Distance Ladder as a outlier against 54 independent measurements points toward unrecognized systematics within the calibration chain. Potential systematic effects include:
Cepheid systematics: Metallicity dependence of the period-luminosity relation, crowding and blending effects, or extinction corrections.
SNe Ia systematics: Environmental dependencies (host galaxy mass, local star formation), evolution with redshift, or standardization procedure biases.
Calibrator physics: Differences between calibrators observed in anchor galaxies vs. SNe Ia hosts.
Sample selection: Potential biases in which galaxies are selected for Cepheid observation.
However, the “systematics” interpretation faces its own challenges:
Multiple independent calibrators (Cepheids, TRGB, JAGB, Miras, SBF) yield consistent high- values.
JWST observations have largely confirmed HST Cepheid photometry [Riess2025].
Model-independent methods like the tailored EPM [Vogl2024] support the high- solution.
V.4.4 Hybrid Scenarios
Given the complexity of the evidence, a hybrid scenario involving both partial distance ladder systematics and deviations from CDM dynamics may be required. For example:
A % systematic in the distance ladder calibration (reducing its by km s-1 Mpc-1)
Combined with a late-time modification to (accounting for the remaining – km s-1 Mpc-1)
Such a scenario is constrained but not excluded by current data.
V.5 Falsifiability and Future Tests
A robust scientific framework requires clear predictions that can be tested by future observations. We outline the key tests that would distinguish between the competing interpretations:
V.5.1 What Would Confirm Distance Ladder Systematics?
The “systematics” hypothesis would be strongly supported if:
1.
Category 3 measurements converge toward km s-1 Mpc-1 as precision improves.
2.
Future Megamaser measurements yield km s-1 Mpc-1 with reduced uncertainties.
3.
The tailored EPM, applied to larger SNe II samples, yields km s-1 Mpc-1.
4.
A specific systematic is identified in the distance ladder (e.g., a metallicity correction error) that shifts its down by km s-1 Mpc-1.
5.
Independent geometric distance methods (e.g., detached eclipsing binaries in external galaxies) yield distances inconsistent with Cepheids.
V.5.2 What Would Confirm New Physics?
The “new physics” hypothesis would be strongly supported if:
1.
Category 3 measurements converge toward km s-1 Mpc-1 as precision improves.
2.
Multiple independent model-independent methods (EPM, Megamasers, future techniques) consistently yield km s-1 Mpc-1.
3.
GW standard sirens, analyzed without assuming CDM (e.g., using cosmographic methods), yield higher than CDM-based analyses.
4.
Direct evidence for deviations from CDM at emerges from other probes (e.g., DESI dark energy constraints, growth rate measurements).
5.
A theoretical model is developed that simultaneously explains the Hubble tension, the Category 2 vs. Category 3 offset, and passes all other cosmological tests.
V.6 Observational Priorities
Based on our analysis, we identify the following observational priorities for resolving the Hubble tension:
1.
Expand Category 3 (Pure Local One-Step): This category currently contains only 15 measurements with significant internal heterogeneity. Additional model-independent measurements are crucial for determining whether the intermediate km s-1 Mpc-1 persists or converges toward one of the two solutions.
2.
Improve tailored EPM precision: The Vogl et al. [Vogl2024] result is currently the strongest model-independent support for the high- solution outside the traditional Distance Ladder. Larger SNe II samples observed with JWST spectroscopy could significantly reduce uncertainties.
3.
Expand Megamaser sample: Additional Megamaser host galaxies at (where peculiar velocities are subdominant) would provide crucial geometric distance constraints.
4.
Accumulate GW standard sirens: The LIGO-Virgo-KAGRA O4 and O5 [LIGOScientific:2025jau] runs will significantly increase the sample of both bright and dark sirens. Crucially, these should be analyzed both with and without CDM assumptions to test for model-dependent biases.
5.
Model-independent Strong Lensing: Future time-delay measurements should emphasize cosmographic or model-independent analyses alongside standard CDM fits.
6.
Cross-calibration tests: Direct comparisons between different calibrators (Cepheids vs. TRGB vs. JAGB) in the same galaxies can identify potential systematics.
V.7 Concluding Remarks
The Hubble tension remains one of the most significant challenges to the standard cosmological model. Our analysis of 83 Sound-Horizon-independent measurements reveals that the tension is more accurately characterized as a “Distance Ladder vs. The Rest” crisis rather than an “Early vs. Late Universe” discrepancy.
The key findings of this work are:
1.
The Distance Ladder yields km s-1 Mpc-1, while all other Sound-Horizon-free methods combined yield km s-1 Mpc-1, a ( correlation-adjusted) tension.
2.
Local measurements do not universally favor high . When analyzed under CDM assumptions, local probes yield values consistent with Planck.
3.
A internal tension exists between CDM-dependent (Category 2) and model-independent (Category 3) local measurements, with the CDM assumption reducing inferred by km s-1 Mpc-1.
4.
Specific model-independent measurements (tailored EPM,) support the high- solution, complicating the pure “systematics” interpretation.
5.
The evidence disfavors pure early-universe solutions (like Early Dark Energy) but does not definitively confirm either late-time modifications or Distance Ladder systematics.
The resolution of the Hubble tension will likely require a combination of improved observations, careful systematic analysis, and potentially a revision of our understanding of either astrophysical calibrators or fundamental cosmology. The four-category classification scheme introduced in this work provides a framework for tracking progress toward this resolution as new measurements become available.
Whatever the ultimate explanation, the Hubble tension has already achieved a valuable scientific outcome: it has motivated an unprecedented scrutiny of both our measurement techniques and our theoretical assumptions, pushing cosmology toward greater rigor and potentially toward new discoveries.Recent results from the Atacama Cosmology Telescope DR6 provide additional CMB constraints that will be crucial for further testing the robustness of our conclusions [AtacamaCosmologyTelescope:2025blo, AtacamaCosmologyTelescope:2025nti]
Acknowledgements.
This research was supported by COST Action CA21136 - Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse), supported by COST (European Cooperation in Science and Technology).Appendix A Complete Data Tables
In this appendix, we provide expanded versions of the data tables presented in the main text, including additional metadata and methodological details for each measurement. These tables serve as a comprehensive reference for future analyses and allow readers to assess the classification of individual measurements.
| Category | Range | ||||
| 1. Distance Ladder | 29 | 72.74 | 0.40 | 0.74 | 67.8–76.9 |
| 2. Local One-Step (CDM) | 33 | 67.18 | 0.90 | 0.85 | 59.1–86.2 |
| 3. Pure Local One-Step | 15 | 70.38 | 1.00 | 0.78 | 65.9–75.5 |
| 4. CMB Sound Horizon Free | 6 | 68.71 | 0.63 | 0.36 | 61–69.5 |
| Combined One-Step (2+3+4) | 54 | 68.67 | 0.46 | 0.85 | 59.1–86.2 |
Notes: is the inverse-variance weighted mean in km s-1 Mpc-1. Range shows min–max of central values.
A.1 Distance Ladder Measurements: Extended Information
Table 9 provides extended information for all 29 distance-ladder-dependent measurements. We include the geometric anchor(s) employed, the primary distance indicator, the secondary distance indicator (if applicable), and the reported value with uncertainty.
| # | Author (Year) | Anchor | Primary | Secondary | range | |
| 1 | Huang (2019) [Huang2020] | NGC 4258, MW, LMC | Mira | SNe Ia | ||
| 2 | de Jaeger (2020) [deJaeger2020] | LMC, MW | Cepheids/TRGB | SNe II | ||
| 3 | Kourkchi (2020) [Kourkchi2020] | Cepheids, TRGB | Tully–Fisher | — | ||
| 4 | Khetan (2021) [Khetan2021] | Cepheids, TRGB | SBF | SNe Ia | ||
| 5 | Blakeslee (2021) [Blakeslee2021] | NGC 4258, LMC | SBF | SNe Ia | ||
| 6 | Freedman (2021) [Freedman2021] | LMC (DEB), MW | TRGB | SNe Ia | ||
| 7 | Dhawan (2022) [Dhawan2022] | LMC | TRGB | SNe Ia | ||
| 8 | Kenworthy (2022) [Kenworthy2022] | NGC 4258, LMC | Cepheids | — | ||
| 9 | Dhawan (2022) [Dhawan2023] | LMC | TRGB | SNe Ia | ||
| 10 | Dhawan (2022) [Dhawan2023] | NGC 4258, LMC, MW | Cepheids | SNe Ia | ||
| 11 | Scolnic (2023) [Scolnic2024] | LMC | TRGB | SNe Ia | ||
| 12 | de Jaeger (2023) [dejaeger2023] | LMC, MW | Cepheids | SNe II | ||
| 13 | Uddin (2023) [uddin2023] | Multiple | Combined | SNe Ia | ||
| 14 | Uddin (2023) [uddin2023] | Multiple | Combined | SNe Ia | ||
| 15 | Huang (2023) [Huang2024] | NGC 4258, LMC | Mira | SNe Ia | ||
| 16 | Li (2024) [Li2024b] | LMC, SMC | JAGB | SNe Ia | ||
| 17 | Chavez (2024) [chavez2025] | Cepheids | H II – | — | ||
| 18 | Boubel (2024) [Boubel2024] | Cepheids, TRGB | Tully–Fisher | — | ||
| 19 | Lee (2024) [Lee2024] | LMC | JAGB | SNe Ia | ||
| 20 | Freedman (2024) [Freedman2024] | LMC (DEB), MW | Cepheids | SNe Ia | ||
| 21 | Said (2024) [Said2025] | SBF | Fund. Plane | — | ||
| 22 | Zhang (2024)a [Zhang2024] | — | Pairwise Vel. | — | ||
| 23 | Jensen (2025) [Jensen2025] | LMC, MW | TRGB | — | ||
| 24 | Wojtak (2025) [Wojtak2025] | NGC 4258, LMC, MW | Cepheids | SNe Ia | ||
| 25 | Bhardwaj (2025) [Bhardwaj2025] | NGC 4258, LMC | Mira | SNe Ia | ||
| 26 | Newman (2025) [Newman2025] | LMC | TRGB | SNe Ia | ||
| 27 | Riess (2025)b [Riess2025] | NGC 4258, LMC, MW | Cepheids | SNe Ia | ||
| 28 | Kudritzki (2025) [kudritzki2025] | FGLR | Tully–Fisher | — | ||
| 29 | Wagner (2025)c [wagner2025] | — | TRGB | Infall |
Notes: DEB = Detached Eclipsing Binaries; FGLR = Flux-weighted Gravity–Luminosity Relation; SBF = Surface Brightness Fluctuations. Entries #19 (Lee 2024) and #29 (Wagner 2025) are notable low- outliers within this category.
a
Includes systematic uncertainties not fully accounted for in the original analysis; see Appendix B.
b
We adopt this measurement for all numerical values in this work. For comparisons, we use the SH0ES 2022 result [Riess2022] as reference baseline, given its established presence in the literature.
c
Assumes a local expansion parameter with no uncertainties. We adopt in our analysis; see Appendix B.
The measurements span a range of calibrator types:
Cepheids: 11 measurements, anchored primarily to NGC 4258, LMC, and MW parallaxes
TRGB: 7 measurements, typically anchored to LMC detached eclipsing binaries
JAGB: 2 measurements with notably discrepant results (Li vs. Lee)
Mira variables: 3 measurements
SBF/Tully-Fisher: 5 measurements using secondary calibration
Other: 2 measurements (pairwise velocities, HII regions)
A.2 One-Step Measurements: Extended Information
Table 10 provides extended information for all 54 One-Step measurements from Categories 2, 3, and 4. We include the physical observable, the cosmological probe, the model assumptions (which determine the category assignment), and the reported value with uncertainty.
The “Model” column indicates the category assignment:
CDM: Category 2 — Analysis explicitly assumes flat CDM expansion history
Independent: Category 3 — Model-independent analysis (geometric, cosmographic, or direct )
CMB–NSH: Category 4 — Uses CMB data but avoids Sound Horizon (No Sound Horizon)
The measurements span diverse physical probes:
Strong Gravitational Lensing: 15 measurements (time delays from quasars, SNe, and clusters)
Gravitational Waves: 8 measurements (bright and dark sirens)
Fast Radio Bursts: 8 measurements (dispersion measure–redshift relation)
Cosmic Chronometers: 4 measurements (differential galaxy ages)
Megamasers: 2 measurements (geometric disk distances)
CMB-based (no ): 6 measurements (SZ effect, scale)
Other: 11 measurements (EPM, kilonovae, gamma-ray attenuation, etc.)
| # | Cat. | Author (Year) | Observable | Probe | Model | z range | |
| Category 2: Local One-Step (CDM Assumption) — 33 measurements | |||||||
| 1 | 2 | Wu (2021) [Wu2022] | Dispersion measure, | FRBs | CDM | 0.0039–0.695 | |
| 2 | 2 | Abbott (2021) [Abbott2021] | GW strain amplitude | CBC dark sirens | CDM | ||
| 3 | 2 | Philcox (2022) [Philcox2022] | BAO scale, shape | Matter–rad. eq. | CDM | 0.38–0.61 | |
| 4 | 2 | Shajib (2023) [Shajib2023] | Time delays, dynamics | Lensed quasars | CDM | 0.657 | |
| 5 | 2 | Napier (2023) [Napier2023] | Time delays | Cluster lenses | CDM | 1.734–2.805 | |
| 6 | 2 | Kelly (2023) [Kelly2023] | Time delays | SN Refsdal | CDM | 1.49 | |
| 7 | 2 | Dominguez (2023) [Dominguez2023] | -ray opacity | Blazar spectra | CDM | ||
| 8 | 2 | Gao (2023) [gao2023] | DM– + SNe Ia | FRBs | CDM | 0.039–2.3 | |
| 9 | 2 | Liu (2023) [liu2023h] | Time delays | Cluster lensed QSO | CDM | ||
| 10 | 2 | Fung (2023) [Fung2023] | GW strain only | NSBH population | CDM | ||
| 11 | 2 | Martinez (2023) [Martinez2023] | Time delays | Lensed quasars | CDM | 1.734 | |
| 12 | 2 | Ballard (2023) [Ballard2023] | GW strain + galaxy | Dark siren | CDM | ||
| 13 | 2 | Moresco (2023) [Moresco2023] | Spectroscopic ages | Passive galaxies | CDM | 0.07–1.965 | |
| 14 | 2 | Alfradique (2024) [Alfradique2023] | GW strain + galaxy | Dark sirens | CDM | ||
| 15 | 2 | Grillo (2024) [Grillo2024] | Time delays | HFF clusters | CDM | 1.240–3.703 | |
| 16 | 2 | Pascale (2024) [Pascale2024] | Time delays | H0pe SN (JWST) | CDM | 1.78 | |
| 17 | 2 | Hernandez (2024) [Hernandez2024] | GW + BGP | Spectral sirens | CDM | ||
| 18 | 2 | Bom (2024) [Bom2024] | GW + kilonova | O4a events | CDM | 0.1–0.6 | |
| 19 | 2 | TDCOSMO XVI (2024) [TDCOSMO2024] | Time delays | WGD2038 | CDM | 0.777 | |
| 20 | 2 | Gao (2024) [Gao2025a] | DM– relation | Localized FRBs | CDM | 0.0008–1.016 | |
| 21 | 2 | Li (2024) [Li2024c] | Time delays (corrected) | Lensed quasars | CDM | ||
| 22 | 2 | Yang (2024) [Yang2024] | DM + scattering | FRBs | CDM | 0.034–1.016 | |
| 23 | 2 | Piratova-Moreno (2025) [PiratovMoreno2025] | DM– relation | FRBs (MLE) | CDM | 0.0008–1.354 | |
| 24 | 2 | Barua (2025) [Barua2025] | Maser dynamics | MCP sample | CDM | 0.002–0.034 | |
| 25 | 2 | Zhang (2025) [Zhang2025] | PRS association | FRBs | CDM | 0.03–0.25 | |
| 26 | 2 | Gao (2025) [Gao2025] | RM–PRS correlation | FRBs | CDM | 0.098–0.241 | |
| 27 | 2 | Loubser (2025) [Loubser2025] | Spectroscopic ages | BCGs | CDM | 0.3–0.7 | |
| 28 | 2 | Beirnaert (2025) [Beirnaert2025] | GW strain | Dark sirens (avg) | CDM | ||
| 29 | 2 | Bahr-Kalus (2025) [Bahr-Kalus2025] | Turnover scale + BAO | DESI | CDM | 0.4–3.1 | |
| 30 | 2 | Birrer (2025) [Birrer2025] | Time delays + kinematics | Multiple lenses | CDM | 0.082–0.745 | |
| 31 | 2 | Liu (2025) [liu2025a] | Strong grav. lensing | Multiple lenses | CDM | 1.49 | |
| 32 | 2 | Pierel (2025) [Pierel2025] | Time delays | SN Encore | CDM | 1.95 | |
| 33 | 2 | Paic (2025) [paic2025tdcosmoxxivmeasurementhubble] | Strong Lensing | SMARTS | CDM | 2.32 | |
| Category 3: Pure Local One-Step (Model Independent) — 15 measurements | |||||||
| 34 | 3 | Kuo (2013) [Kuo2013] | Maser positions, velocities | Megamaser disk | Independent | 0.034 | |
| 35 | 3 | Gao (2015) [Gao2015] | Maser positions, velocities | Megamaser disk | Independent | 0.028 | |
| 36 | 3 | Bulla (2022) [bulla2022] | GW + kilonova light curve | BNS merger | Independent | 0.01 | |
| 37 | 3 | Zhang (2022) [Zhang2022] | Spectroscopic ages | Passive galaxies | Independent | 0.009–2.545 | |
| 38 | 3 | Du (2023) [Du2023] | Time delays + GRBs | Lensed quasars | Independent | 0.654–2.375 | |
| 39 | 3 | Palmese (2023) [Palmese2023] | GW + afterglow | GW170817 | Independent | 0.01 | |
| 40 | 3 | Sneppen (2023) [Sneppen2023] | Photospheric velocity | Kilonova | Independent | 0.0986 | |
| 41 | 3 | Liu (2023) [liu2023a] | Lensing + SNe Ia | Combined | Independent | 0.01–2.3 | |
| 42 | 3 | Gonzalez (2024) [Gonzalez2024] | Gas fraction + SNe Ia | Clusters | Independent | 0.01–2.3 | |
| 43 | 3 | Li (2024) [Li2024a] | Time delays | SN Refsdal | Independent | 0.001–2.261 | |
| 44 | 3 | Jaiswal (2024) [Jaiswal2024] | RM lag | NGC 5548 | Independent | 0.017 | |
| 45 | 3 | Vogl (2024) [Vogl2024] | Photospheric expansion | SNe II | Independent | 0.01–0.04 | |
| 46 | 3 | Song (2025) [Song2025] | GW + lensing | Combined | Independent | ||
| 47 | 3 | Colaco (2025) [Colaco2025] | Lensing + SNe Ia | Joint | Independent | 0.001–2.261 | |
| 48 | 3 | Favale (2025) [Favale2025] | Ages + SNe Ia | CC + Pantheon+ | Independent | 0.07–1.965 | |
| Category 4: CMB Sound Horizon Free — 6 measurements | |||||||
| 49 | 4 | Reese (2003) [Reese2003] | SZ decrement, X-ray | Galaxy clusters | CMB–NSH | 0.02–0.83 | |
| 50 | 4 | Colaco (2023) [Colaco2023] | SZ + X-ray + SNe Ia | Clusters | CMB–NSH | 0.01–2.3 | |
| 51 | 4 | Pogosian (2024) [Pogosian2024] | BAO peak + BBN | DESI BAO | CMB–NSH | 0.1–4.16 | |
| 52 | 4 | Pogosian (2024) [Pogosian2024] | BAO peak + BBN | Pre-DESI BAO | CMB–NSH | 0.1–4.16 | |
| 53 | 4 | Bahr-Kalus (2025) [Bahr-Kalus2025] | Turnover scale + SNe | DESI + Pantheon+ | CMB–NSH | 0.4–3.1 | |
| 54 | 4 | Krolewski (2025) [Krolewski2025] | DESI DR1 + CMB | No sound horizon | CMB–NSH | 0.1–2.33 | |
Notes: CMB–NSH = CMB-based, No Sound Horizon; MLE = Maximum Likelihood Estimator; PRS = Persistent Radio Source; RM = Reverberation Mapping; BGP = Binary Gravitational-wave Population; HFF = Hubble Frontier Fields; BNS = Binary Neutron Star; NSBH = Neutron Star–Black Hole.
Includes systematic uncertainties not fully accounted for in the original analysis; see Appendix B.
A.3 Statistical Summary by Category
Table 8 provides a detailed statistical breakdown of each category, including the number of measurements, weighted mean, standard error, reduced chi-squared, and the range of individual measurements.
Appendix B Individual Measurement Descriptions
In this appendix, we provide detailed descriptions of selected measurements that are particularly important for understanding the Hubble tension landscape. We focus on measurements that either anchor the tension (high-precision results), represent methodological innovations, or exhibit notable discrepancies with other measurements in their class.
B.1 Key Distance Ladder Measurements
B.1.1 Riess et al. (2022, 2025): The SH0ES Program
The SH0ES (Supernova for the Equation of State) collaboration has produced the most precise Distance-Ladder measurements of . The methodology employs a three-rung distance ladder:
Rung 1 (Geometric Anchors): Three independent anchors establish the absolute distance scale:
Gaia EDR3 parallaxes to Milky Way Cepheids (corrected for zero-point offset)
Megamaser geometric distance to NGC 4258 [Reid2019]
Detached eclipsing binary distances to the LMC [Pietrzynski2019]
Rung 2 (Cepheid Calibration): Cepheid variables observed with HST WFC3 in the F160W (H-band) filter provide period–luminosity relations in 37+ galaxies hosting Type Ia supernovae. The use of Wesenheit magnitudes, , eliminates reddening dependence by construction.
Rung 3 (Hubble Flow SNe Ia): The calibrated SNe Ia absolute magnitude is applied to the Pantheon+ sample in the redshift range .
The 2022 HST result [Riess2022] yielded km s-1 Mpc-1 (1.4% precision). The 2025 JWST result [Riess2025] yields km s-1 Mpc-1 (1.3% precision), confirming the HST photometry and ruling out crowding-induced biases as the source of the tension.
Systematic Error Budget: The SH0ES analyses include comprehensive systematic assessments for Cepheid photometry ( km s-1 Mpc-1), metallicity effects ( km s-1 Mpc-1), crowding corrections ( km s-1 Mpc-1), SNe Ia standardization ( km s-1 Mpc-1), and peculiar velocities ( km s-1 Mpc-1).
B.1.2 Freedman et al. (2021, 2024): The CCHP TRGB Program
The Chicago-Carnegie Hubble Program (CCHP) developed the Tip of the Red Giant Branch (TRGB) as an alternative to Cepheids. The TRGB method exploits the helium flash that terminates red giant branch evolution at a nearly constant luminosity (), producing a sharp discontinuity detectable via edge-detection algorithms (Sobel filter).
The 2021 analysis [Freedman2021] yielded km s-1 Mpc-1, intermediate between Planck and SH0ES. This sparked debate about whether TRGB measurements genuinely favor lower or whether methodological differences explain the offset from SH0ES.
The 2024 JWST analysis [Freedman2024] found km s-1 Mpc-1 from Cepheids, now more consistent with SH0ES. The CCHP continues to analyze TRGB and JAGB calibrations from JWST data.
B.1.3 Lee et al. (2024): The Low- JAGB Outlier
The Lee et al. [Lee2024] JAGB analysis yields km s-1 Mpc-1, consistent with Planck and below other distance ladder measurements. This represents the most significant outlier within Category 1.
The JAGB method uses carbon stars on the Asymptotic Giant Branch, which exhibit a narrow luminosity distribution in the J-band. The discrepancy with Li et al. [Li2024b] () appears to originate from differences in:
JAGB zero-point calibration (LMC anchor treatment)
Photometric selection criteria for carbon stars
Assumed shape of the JAGB luminosity function
This measurement highlights that even within a single indicator class, methodological choices can lead to km s-1 Mpc-1 differences in inferred .
B.1.4 Wagner et al.
The measurement by Wagner et al. [wagner2025] (2025) reports km s-1 Mpc-1 using the local expansion parameter , based on the infall velocity relation . However, the choice of appears somewhat arbitrary; standard models of the local Hubble flow typically adopt . Since , adopting —a value intermediate between the theoretical expectation and the assumed value—would yield km s-1 Mpc-1, shifting the result toward the Planck value and closer to the Distance Ladder mean value. The systematic uncertainty in propagates directly into but this source of error was not included in the original analysis [wagner2025]. However, we do include this effect in our analysis and the data shown in Table 9.
B.1.5 Zhang et al.:pairwise velocities
Zhang et al. (2024) [Zhang2024] measured the Hubble constant using galaxy pairwise peculiar velocities from the Cosmicflows-4 catalog, obtaining km s-1 Mpc-1 (statistical). However, as discussed in [Zhang2024], several systematic uncertainties are not fully accounted for in the reported error budget. The distance calibration of CF-4 galaxies introduces uncertainties of 2–3%, translating to 1.5–2.3 km s-1 Mpc-1 in . The halo mass matching between simulations and observations, which is crucial for their method, contributes an estimated systematic error of 1–2% (0.8–1.5 km s-1 Mpc-1). Additionally, the modeling of pairwise velocities in the nonlinear regime, where theoretical predictions are less certain, introduces uncertainties of 2–3% (1.5–2.3 km s-1 Mpc-1). The choice of the maximum separation scale ( Mpc) and the selection of the velocity threshold add further systematic uncertainties of order 1% (0.8 km s-1 Mpc-1). Cosmic variance, despite mitigation through multiple simulation realizations, remains a source of systematic error. Combining these contributions in quadrature yields a total systematic uncertainty of approximately 3–4 km s-1 Mpc-1, which is comparable to or larger than the quoted statistical uncertainty. Therefore, a more complete error budget would be km s-1 Mpc-1, or equivalently km s-1 Mpc-1 when statistical and systematic errors are combined in quadrature. This corrected level of uncertainty is used in our analysis.
B.2 Key One-Step Measurements
B.2.1 TDCOSMO: Evolution of Strong Lensing Constraints
The Time-Delay Cosmography (TDCOSMO) collaboration has produced a series of Strong Lensing measurements that illustrate the critical importance of lens mass modeling.
H0LiCOW (Millon et al. 2020): The original analysis yielded km s-1 Mpc-1, assuming power-law mass profiles. This result was in tension with Planck.
TDCOSMO IV (Birrer et al. 2020): [Birrer2020] Incorporating stellar kinematics to constrain the mass-sheet degeneracy (MSD) shifted the result to km s-1 Mpc-1—now consistent with Planck. This dramatic shift demonstrated that power-law profiles alone cannot break the MSD.
Recent results (2023–2025): Subsequent analyses using hierarchical modeling, additional lens systems, and improved kinematics yield in the range – km s-1 Mpc-1, depending on modeling assumptions [Shajib2023, Birrer2025].
We exclude the original Millon et al. (2020) result from our compilation because its methodology is now recognized as insufficient for robust inference.
We exclude the original Pesce et al. measurement from our sample, as the more recent Barua et al. analysis uses the same underlying data while addressing the known systematic uncertainties of the earlier measurement
B.2.2 Kelly et al. (2023): SN Refsdal Time Delays
Supernova Refsdal, the first multiply-imaged supernova [Kelly2015], provides unique advantages for time-delay cosmography: the supernova light curve allows unambiguous identification of multiple images, unlike quasar variability.
The Kelly et al. [Kelly2023] analysis uses time delays from the “Einstein Cross” configuration plus the predicted reappearance of image “SX.” The result, km s-1 Mpc-1, is consistent with Planck.
Follow-up analyses with different cluster mass models yield consistent results: Li et al. [Li2024a] find km s-1 Mpc-1. The consistency across mass models provides confidence in the robustness of this constraint.
B.2.3 Piratova-Moreno et al.:FRBs
Piratova-Moreno et al. (2025) [PiratovMoreno2025] employed Fast Radio Bursts (FRBs) as cosmological probes to measure the Hubble constant using a catalog of 98 localized FRBs. Using the Maximum Likelihood Estimate (MLE) method with confirmed FRBs, they obtained km s-1 Mpc-1 (statistical), while their mock FRB catalog yielded km s-1 Mpc-1. However, this method is subject to substantial systematic uncertainties that are not fully quantified in the reported error budget. The primary systematic error arises from the decomposition of the observed dispersion measure (DM) into its constituent components: the Milky Way interstellar medium and halo (), the intergalactic medium (), and the host galaxy (). The uncertainty in depends on the Galactic electron density model used (NE2001 vs. YMW16), with differences of 20% at low Galactic latitudes, contributing 1.5–2 km s-1 Mpc-1 to the uncertainty. The host galaxy contribution, , is highly uncertain and varies dramatically between FRBs—from 30 pc cm-3 for typical hosts to over 900 pc cm-3 for extreme cases like FRB 20190520B. Studies suggest uncertainties of 50–100 pc cm-3, which translate to systematic errors of 2–3 km s-1 Mpc-1 in . Additionally, the baryon fraction in the IGM () introduces further systematic uncertainty; recent estimates range from to , corresponding to 1–2 km s-1 Mpc-1 uncertainty in . The choice of DM– relation (linear vs. power-law) also significantly impacts results, with Piratova-Moreno et al. reporting values ranging from km s-1 Mpc-1 (linear) to km s-1 Mpc-1 (power-law), indicating substantial model-dependent systematic uncertainties of order 5–15 km s-1 Mpc-1. Combining these systematic contributions in quadrature yields a total systematic error of approximately 6–8 km s-1 Mpc-1 for the MLE result, dominating over the statistical uncertainty. A more realistic error budget would therefore be km s-1 Mpc-1, or equivalently km s-1 Mpc-1 when combined in quadrature, reflecting the current limitations of FRB-based cosmology. This is the value used in our analysis.
B.2.4 Vogl et al. (2024): Tailored EPM
The Vogl et al. [Vogl2024] tailored Expanding Photosphere Method represents the most precise model-independent measurement supporting the high- solution outside the traditional distance ladder.
Method: EPM determines distances geometrically by combining photospheric angular size (from flux and temperature) with expansion velocity (from Doppler shifts). The “tailored” approach uses detailed NLTE radiative transfer models customized to individual SNe II in the redshift range , rather than empirical dilution factors.
Result: km s-1 Mpc-1, fully consistent with the Distance Ladder.In this measurement, the authors do not explicitly report the systematic uncertainties in their work. However, they estimate that the systematic errors are approximately of the same magnitude as the statistical ones, and we have adopted this estimate in our analysis.
Significance: This measurement is independent of:
SNe Ia physics (uses SNe II instead)
Cepheid/TRGB calibrators (direct geometric distance)
CDM assumptions (no integration required)
The agreement with the Distance Ladder suggests either a common systematic affecting both methods or that the high reflects true local physics.
Caveats: The method relies on supernova atmosphere modeling; current sample size is limited; further validation with independent codes is needed.
B.2.5 Barua et al. (2025): Megamaser Cosmology Project
The Megamaser Cosmology Project measures geometric distances to galaxies with circumnuclear water maser disks. The Barua et al. [Barua2025] frequentist analysis yields km s-1 Mpc-1.
Method: VLBI observations resolve the maser disk structure, providing angular diameter distances via Keplerian dynamics. The method is independent of stellar physics.
Classification note: We assign this measurement to Category 2 (CDM–depended) because it relies on the CDM prediction for the angular diameter distance. Earlier MCP analyses that adopted SH0ES flow models for peculiar velocity corrections would be classified differently.
Evolution: Earlier measurements (Kuo 2013: ; Gao 2015: ) favored lower values with larger uncertainties.
B.2.6 Pogosian et al. (2024): Sound-Horizon-Free BAO
The Pogosian et al. [Pogosian2024] analysis demonstrates that can be constrained from BAO data without assuming the CMB Sound Horizon.
Method: Instead of using as a standard ruler, this analysis treats as a free parameter and constrains through:
The shape of from BAO peak positions
Big Bang Nucleosynthesis constraints on
The matter-radiation equality scale
Results: DESI BAO yields km s-1 Mpc-1; pre-DESI BAO yields km s-1 Mpc-1.
Significance: These measurements anchor the low- solution in Category 4. Their high precision (%) means they dominate the weighted mean of this category, contributing to the very low .
B.2.7 Gravitational Wave standard sirens
GW standard sirens provide constraints independent of electromagnetic distance ladders:
Bright sirens (GW170817): The binary neutron star merger with optical counterpart yields km s-1 Mpc-1 [Palmese2023], limited by the peculiar velocity of NGC 4993 and viewing angle degeneracy.
Dark sirens (GWTC-3): Statistical analysis of 47 events without counterparts yields km s-1 Mpc-1 [Abbott2021], consistent with Planck.
The weighted mean of all GW measurements ( km s-1 Mpc-1) supports the low- solution, though with larger uncertainties than other probes.
B.2.8 Fast Radio Burst Measurements
FRB measurements use the dispersion measure–redshift relation to probe the cosmic baryon distribution:
Challenges: Significant systematic uncertainties arise from:
Host galaxy DM contribution (poorly constrained)
IGM density fluctuations
Sample selection effects
Results: Published values span – km s-1 Mpc-1. Maximum-likelihood approaches [PiratovMoreno2025] tend to favor lower values ( km s-1 Mpc-1).
As FRB samples grow and host DM distributions become better characterized, this probe may provide competitive constraints.
Appendix C Correlation Matrix Estimates
A fundamental assumption of the weighted mean analysis in Section III is that measurements within each category are statistically independent. This assumption is violated to varying degrees due to shared calibration data, common analysis frameworks, and overlapping samples. Here we assess the impact of correlations on our conclusions.
C.1 Sources of Correlation
C.1.1 Distance Ladder Correlations
Category 1 measurements share several inputs that induce positive correlations:
Geometric anchors: Most Cepheid measurements use NGC 4258, LMC, and/or MW parallaxes. The anchor uncertainties propagate coherently.
SNe Ia samples: Hubble-flow SNe Ia are predominantly from Pantheon+ [Brout2022], with substantial overlap. standardization parameters (, , ) are often shared.
Period–luminosity relations: Cepheid analyses adopt similar functional forms and metallicity corrections.
Crowding corrections: HST analyses use similar artificial star tests.
We estimate an effective correlation coefficient – for typical pairs of Distance Ladder measurements.
C.1.2 One-Step Correlations
Categories 2, 3, and 4 exhibit weaker correlations due to methodological diversity, but some correlations exist:
Strong lensing: Multiple analyses of the same system (e.g., SN Refsdal) share time delay data and are highly correlated.
Gravitational waves: Analyses using overlapping GW events share strain data and waveform models.
FRBs: Later analyses often include events from earlier samples.
Cosmological priors: Many analyses assume flat CDM with , inducing weak correlations.
We estimate – for the One-Step sample.
C.2 Methodology for Correlation-Adjusted Estimates
For correlated measurements, the generalized weighted mean and its uncertainty are:
where is the covariance matrix with elements .
For simplicity, we adopt a uniform correlation structure within each category:
C.3 Impact on Combined Constraints
C.3.1 Distance Ladder (Category 1)
With :
compared to km s-1 Mpc-1 under independence.
The central value shifts by only km s-1 Mpc-1 (within statistical uncertainty), but the uncertainty increases by %.
C.3.2 Combined One-Step (Categories 2+3+4)
With :
compared to km s-1 Mpc-1 under independence.
The central value shifts by km s-1 Mpc-1 , and the uncertainty increases by %.
C.3.3 Tension Between Categories
The correlation-adjusted tension is:
This is reduced from under independence but remains highly significant.
C.4 Sensitivity Analysis
Table 11 shows how the tension varies with assumed correlation strength.
| Tension | ||||
| 0.0 | 0.0 | 0.40 | 0.46 | |
| 0.2 | 0.1 | 0.61 | 0.60 | |
| 0.4 | 0.15 | 0.58 | 0.60 | |
| 0.5 | 0.2 | 0.54 | 0.59 | |
| 0.6 | 0.25 | 0.49 | 0.58 |
Notes: Uncertainties in km s-1 Mpc-1. Central values are relatively stable across correlation assumptions.
C.5 Conclusions from Correlation Analysis
Key findings:
1.
Central values are robust: The weighted means shift by km s-1 Mpc-1 across plausible correlation assumptions.
2.
Uncertainties increase: Accounting for correlations increases uncertainties by –%, depending on the assumed correlation strength.
3.
Tension remains significant: For reasonable correlation estimates (–, –), the tension remains at the – level.
4.
Qualitative conclusions unchanged: The characterization of the Hubble tension as “Distance Ladder vs. The Rest” is robust to correlation modeling.
A fully rigorous treatment would require constructing the complete covariance matrix from detailed assessment of shared data and methodology—a task beyond the scope of this work but warranted for future precision analyses.

### v2

@@
-Dissecting the Hubble tension: Insights from a diverse set of Sound Horizon-free measurements
+Dissecting the Hubble tension: Insights from a diverse set of Sound-Horizon-Free measurements
 Abstract
 The Hubble tension is commonly framed as a discrepancy between local, late-time measurements favoring km s-1 Mpc-1 and early-time, Sound-Horizon-based measurements favoring km s-1 Mpc-1.
-We challenge this viewpoint by analyzing 83 Sound-Horizon-independent measurements, categorized into four classes: Distance Ladder measurements using local calibrators; Local One-Step CDM measurements assuming the standard expansion history; Pure Local One-Step measurements independent of shape; and CMB Sound Horizon free measurements using CMB data without the Sound Horizon scale.
-Our analysis reveals that the 29 Distance Ladder measurements yield km s-1 Mpc-1 (), while the 54 One-Step measurements collectively yield km s-1 Mpc-1 (), a tension exceeding the Planck–SH0ES discrepancy.
-This tension remains significant at after accounting for correlations.
-Among One-Step categories, Local One-Step CDM measurements favor the lowest value ( km s-1 Mpc-1, ), Pure Local One-Step yield an intermediate value ( km s-1 Mpc-1, ), and CMB Sound Horizon Free measurements give km s-1 Mpc-1 ().
+We challenge this viewpoint by analyzing 88 Sound Horizon Free measurements, categorized into four classes: Distance Ladder measurements using local calibrators; Local CDM measurements assuming the standard expansion history; Pure Local measurements independent of shape; and CMB Sound–Horizon–Free measurements using CMB data without the Sound Horizon scale.
+Our analysis reveals that the 30 Distance Ladder measurements yield km s-1 Mpc-1 (), while the 58 Distance Ladder-Independent/Sound Horizon Free measurements collectively yield km s-1 Mpc-1 (), a tension exceeding the Planck–SH0ES discrepancy.
+This tension remains significant at a minimum value of after accounting for correlations.
+Among categories, Local CDM measurements favor the lowest value ( km s-1 Mpc-1), Pure Local yield an intermediate value ( km s-1 Mpc-1), and CMB Sound Horizon Free measurements give km s-1 Mpc-1.
 We conclude that the Hubble tension is better characterized as a discrepancy between the Distance Ladder and all other methodologies, rather than an early-vs-late-time split.
-We also identify a internal tension among One-Step measurements: analyses assuming CDM systematically recover lower values by approximately 3.2 km s-1 Mpc-1 compared to model-independent methods.
-This suggests either unrecognized systematics/new physics in the Distance Ladder or deviations from CDM in the late-time Universe.
+We also identify a internal tension among Distance Ladder Independent/Sound Horizon Free measurements (noting its sensitivity to categorization choices): analyses assuming CDM systematically recover lower values by 3.4 km s-1 Mpc-1 compared to cosmological-model-independent methods.
+This suggests either unrecognized systematics in the Distance Ladder or deviations from CDM or both.
 I Introduction
@@
 The standard view frames this tension as a dichotomy between “early” and “late” times.
-The Planck CMB analysis, assuming CDM, yields [Planck2018], an inference dependent on the Sound Horizon scale at the drag epoch ().
-The local Distance Ladder, anchored by geometric measurements and extended via Cepheids to SNe Ia, favors [Riess2022], reinforced by JWST observations [Riess2025] and extended network analyses [Casertano2025].
-If both are accurate, CDM is incomplete, potentially necessitating new physics [DiValentino2021review, Knox2020, Abdalla:2022yfr, Perivolaropoulos:2021jda, Alestas2020].
-Recent DESI hints of dynamical dark energy evolving across the phantom divide [Efstratiou2025, Ozulker:2025ehg, Scherer:2025esj] further complicate this picture, with phantom crossings previously proposed to address both the Hubble tension and the SNe Ia absolute magnitude discrepancy [Alestas2021].
+The Planck CMB analysis, assuming CDM, yields [141], an inference dependent on the Sound Horizon scale at the drag epoch ().
+The local Distance Ladder, anchored by geometric measurements and extended via Cepheids to SNe Ia, favors [150], reinforced by JWST observations [149] and extended network analyses [31].
+If both are accurate, CDM is incomplete, potentially necessitating new physics [51, 91, 3, 131, 9, 179, 33, 135].
+Various theoretical approaches have been proposed to address the Hubble tension, which can be broadly classified into three categories based on the cosmic epoch at which they introduce new physics [51, 91].
+Early-time models attempt to resolve the tension by modifying conditions during the recombination epoch, thereby reducing the Sound Horizon scale that serves as the standard ruler for CMB-based inference.
+Prominent examples include Early Dark Energy (EDE) scenarios and modified gravity theories.
+However, these approaches face significant challenges: they typically fail to fully resolve the tension, often exacerbate the tension by enhancing matter clustering, and require considerable fine-tuning of model parameters [153, 168, 146, 169, 170, 82, 181, 73].
+Late-time models propose deformations in the Hubble expansion history at low redshifts , often through dynamical dark energy with an evolving equation of state [143, 35, 158, 124, 162, 59, 49, 50, 97, 187, 61].
+To understand the implications, it is essential to first outline the theoretical classification of Dark Energy dynamics.
+The evolution of dark energy is characterized by its equation of state parameter, , defined as the ratio of its pressure to its energy density ().
+In the standard CDM model, this parameter corresponds to a Cosmological Constant and is strictly fixed at .
+However, dynamical dark energy models are generally categorized into two distinct regimes based on deviations from this value: Quintessence (), which is typically modeled by a scalar field rolling down a potential, and the Phantom regime (), where the fluid exhibits super-negative pressure, leading to an increasing energy density over time [120, 166]
+A dynamical evolution in which the parameter traverses the boundary is referred to as Phantom Crossing.
+This transition presents significant theoretical challenges; within the framework of simple perfect fluids or single scalar field models, crossing this boundary typically triggers gradient instabilities or introduces ghost degrees of freedom.
+Consequently, viable phantom crossing scenarios usually necessitate frameworks beyond standard General Relativity, such as Modified Gravity theories or Effective Field Theory (EFT) approaches, to stabilize the transition [57, 118, 166, 160].
+Recent analyses incorporating DESI Data Release 2 (DR2) [4] combined with CMB observations favor a Dynamical Dark Energy scenario characterized by a specific phantom crossing behavior.
+Contrary to simple quintessence models, the data indicate a transition from a phantom-like phase () in the past to a quintessence-like phase () at late times, crossing the phantom divide at approximately , with a present-day value of .
+Crucially, regarding the Hubble tension, this result presents a paradox: due to the geometric anti-correlation between and , a present-day value of implies a lower inferred value for the Hubble constant.
+Consequently, rather than bridging the gap towards the Distance Ladder value ( km s-1 Mpc-1), this specific form of dynamical dark energy risks exacerbating the discrepancy with the high Distance Ladder value.
+[126, 58, 114, 30, 100, 38, 151, 166, 19, 85].
+However, it aligns remarkably well with intermediate values: recent work by Zhang et al.
+using an interacting dark energy model finds from CMBDESI DR2 data (without SH0ES priors), a value consistent with intermediate estimates between Planck and SH0ES [194].
+If the DESI evidence for dynamical dark energy is confirmed, the observed shift between CDM-dependent (Category 2) and model-independent (Category 3) measurements transforms from a statistical curiosity into a physical necessity.
+In this scenario, the Pure Local measurements (Category 3), yielding km s-1 Mpc-1, would emerge as the only methodologically sound Distance Ladder Independent/Sound Horizon Free determinations.
+Conversely, the lower values found in Category 2 would be interpreted as an artifact of imposing an invalid CDM expansion history on local data.
+[57, 120, 160, 10, 165, 71, 118, 152, 159, 2].
+Ultralate-time models propose environmental or physical changes at very low redshifts that could affect the calibration of distance ladder indicators.
+The G-step model [112, 154, 128, 133], which posits a transition in the effective gravitational constant at , represents a well-studied example.
+Related approaches involving screened fifth forces have been explored by Desmond et al.
+[44, 45, 76], who showed that such modifications could systematically affect Cepheid and TRGB distance indicators differently.
+Alternative explanations within this category include local void scenarios, which posit that our local environment is underdense relative to the cosmic average, leading to an apparent acceleration of the local expansion rate [89, 109, 27, 32, 29, 111, 84, 15].
+While these models offer mechanisms to reconcile distance ladder and other measurements without modifying early-universe physics, they require fine-tuning and lack complete theoretical justification for the proposed transitions.
 However, this “Early vs.
 Late” characterization may be oversimplified.
-It assumes all local measurements favor high values, an assumption challenged by “One-Step” measurements independent of the Cepheid/SN Ia calibration chain.
-Our recent analysis [Perivolaropoulos2024] demonstrated a stark bifurcation: Distance Ladder measurements clustered around , while independent One-Step methods statistically favored lower values closer to Planck.
-This suggests the discrepancy may lie between the distance ladder methodology and other probes, rather than between early and late Universe physics [Huang:2024].Indeed, the tension has been shown to have a multidimensional character, with degeneracies between , , and other parameters playing a crucial role [Pedrotti:2024kpn].
-In this work, we present an up-to-date compilation of 83 measurements published through late-2025, all independent of the CMB Sound Horizon scale.
-We introduce a four-category classification based on methodology and model dependence:
-1.
-Category 1 – Distance Ladder (29 measurements): Traditional three-rung approach using geometric anchors to calibrate standard candles (Cepheids, TRGB, JAGB, Miras) that standardize SNe Ia in the Hubble flow.
-2.
-Category 2 – Local One-Step CDM (33 measurements): Local measurements bypassing the Distance Ladder but assuming CDM to infer (e.g., Gravitational Wave sirens, Strong Lensing time delays, FRB dispersion measures).
-3.
-Category 3 – Pure Local One-Step (15 measurements): Model-independent local measurements without assuming a specific shape (e.g., geometric Megamaser distances, Cosmic Chronometers, cosmographic lensing analyses).
-4.
-Category 4 – CMB Sound Horizon Free (6 measurements): CMB-based measurements decoupled from , using alternative scales such as the horizon at matter-radiation equality ().
+It assumes all local measurements favor high values, an assumption challenged by Distance Ladder Independent/Sound Horizon Free measurements independent of the Cepheid/SN Ia calibration chain.
+Our recent analysis [134] demonstrated a stark bifurcation: Distance Ladder measurements clustered around , while independent methods statistically favored lower values closer to Planck.
+This suggests the discrepancy may lie between the distance ladder methodology and other probes, rather than between early and late Universe physics [80].
+Indeed, the tension has been shown to have a multidimensional character, with degeneracies between , , and other parameters playing a crucial role [127].
+In this work, we present an up-to-date compilation of 88 H0 measurements published through late-2025, all independent of the CMB Sound Horizon scale.
+We introduce a four-category classification based on methodology and model dependence: (1) Distance Ladder (30 measurements), (2) Local CDM (33 measurements), (3) Pure Local (cosmological-model-independent, (16 measurements), and (4) CMB Sound Horizon Free (CMB–SHF) (9 measurements).
+The detailed definitions and classification criteria are presented in Section II.2.
 The classification follows a systematic decision process (Figure 1): measurements using multi-rung calibration chains are assigned to Category 1; those utilizing CMB data without to Category 4; remaining measurements are assigned to Category 2 or 3 based on whether they assume CDM.
-The same observational technique can yield measurements in different categories depending on analysis methodology—for example, Megamaser observations with Keplerian models belong to Category 3 [Kuo2013], while those adopting CDM priors belong to Category 2 [Barua2025].
-This classification isolates two distinct tensions: the external tension between the Distance Ladder and all other methods, and an internal tension within One-Step measurements examining whether CDM assumptions systematically shift inferred values.A comprehensive overview of observational tensions and potential resolutions involving both systematics and fundamental physics has been presented in the CosmoVerse White Paper [CosmoVerseNetwork:2025alb].
+The same observational technique can yield measurements in different categories depending on analysis methodology—for example, Megamaser observations with Keplerian models belong to Category 3 [95], while those adopting CDM priors belong to Category 2 [17].
+This classification isolates two distinct tensions: the external tension between the Distance Ladder and all other methods, and an internal tension within these measurements examining whether CDM assumptions systematically shift inferred values.
+A comprehensive overview of observational tensions and potential resolutions involving both systematics and fundamental physics has been presented in the CosmoVerse White Paper [52].
 This paper is structured as follows: Section II details our data compilation and measurement tables.
@@
 To dissect the Hubble tension beyond the traditional “Early vs.
-Late” dichotomy, we have compiled a comprehensive dataset of 83 independent measurements of the Hubble constant, updated through late-2025.
-This compilation builds upon the catalog presented in our previous work [Perivolaropoulos2024], significantly expanded to include recent results from James Webb Space Telescope (JWST) calibration campaigns, the third generation of gravitational wave catalogs (GWTC-3), and the latest releases from the Dark Energy Spectroscopic Instrument (DESI) [DESI:2025zgx].
+Late” dichotomy, we have compiled a comprehensive set of 88 independent measurements of the Hubble constant, updated through late-2025.
+This compilation builds upon the catalog presented in our previous work [134], significantly expanded to include recent results from James Webb Space Telescope (JWST) calibration campaigns, the third generation of gravitational wave catalogs (GWTC-3), and the latest releases from the Dark Energy Spectroscopic Instrument (DESI) [4].
 II.1 Selection Criteria
@@
 3.
-Independence of Estimates: Where multiple analyses utilize the exact same dataset (e.g., multiple re-analyses of the same supernova catalog with identical calibration), we select the most recent or representative result to avoid double-counting, unless the methodologies differ significantly in their systematic assumptions.
+Independence of Estimates: Where multiple analyses utilize the exact same sample (e.g., multiple re-analyses of the same supernova catalog with identical calibration), we select the most recent or representative result to avoid double-counting, unless the methodologies differ significantly in their systematic assumptions.
 4.
@@
 If not, we determine whether it utilizes CMB data (leading to Category 4 if Sound-Horizon-Free).
-For the remaining local One-Step measurements, we examine whether the analysis assumes a CDM expansion history (Category 2) or proceeds model-independently (Category 3).
+For the remaining local measurements, we examine whether the analysis assumes a CDM expansion history (Category 2) or proceeds cosmological-model-independently (Category 3).
 It is important to emphasize that the same observational technique can yield measurements in different categories depending on the analysis approach adopted by the authors.
 Table 1 provides illustrative examples of how specific techniques are distributed across categories based on their methodological implementation.
-| Technique | Category 2 (CDM) | Category 3 (Model-Indep.) |
+Throughout this work, we use the term “cosmological-model-independent” (Category 3) to denote measurements that do not assume a specific functional form for the expansion history —in particular, they do not assume the standard CDM expansion history.
+This should be understood as independence from cosmological model assumptions rather than complete model independence, as all measurements necessarily rely on certain physical assumptions (e.g., standard astrophysics, the FLRW metric, general relativity).
+Similarly, we define Distance Ladder-Independent/Sound Horizon Free (DLI–SHF) measurements (Categories 2, 3, and 4) as any determination that bypasses the traditional three-rung Distance Ladder calibration chain (geometric anchors Cepheids/TRGB/JAGB/Miras SNe Ia in the Hubble flow), regardless of whether it probes early or late cosmic times.
+Some DLI–SHF methods may involve their own calibration procedures (e.g., using cosmic chronometers to calibrate the HII galaxy Hubble diagram), but these do not rely on the standard candle–supernova chain that defines the Distance Ladder.
+| Technique | Category 2 (CDM) | Category 3 (Cosmological-Model-Independent) |
 | Megamasers | Barua (2025): uses CDM | Kuo (2013): Keplerian disk only |
@@
 II.3 Category 1: Distance Ladder Measurements
-This category comprises 29 measurements that employ the traditional multi-rung calibration method.
+This category comprises 30 measurements that employ the traditional multi-rung calibration method.
 These measurements anchor the absolute magnitude of Type Ia supernovae (SNe Ia) in the Hubble flow using local standard candles calibrated by geometric distances.
@@
 The three-rung structure proceeds as follows: (1) geometric anchors (Gaia parallaxes, NGC 4258 Megamaser distance, LMC eclipsing binaries) establish absolute distances to nearby calibrators; (2) these calibrators (Cepheids, TRGB, JAGB, Miras) are observed in galaxies that also host SNe Ia, transferring the distance scale; (3) the calibrated SNe Ia absolute magnitude is applied to supernovae in the Hubble flow () to infer .
-Our compilation (Table 2) includes Cepheid-based measurements and recent JWST updates [Riess2025], TRGB-based measurements including CCHP results [Freedman2024], and alternative calibrators such as JAGB stars, Mira variables, Surface Brightness Fluctuations (SBF), and the Tully-Fisher relation.
-We also include the galaxy pairwise velocity method [Zhang2024], which, while methodologically distinct, relies on calibration from distance ladder measurements.
+Our compilation (Table 2) includes Cepheid-based measurements and recent JWST updates [149], TRGB-based measurements including CCHP results [62], and alternative calibrators such as JAGB stars, Mira variables, Surface Brightness Fluctuations (SBF), and the Tully-Fisher relation.
+We also include the galaxy pairwise velocity method [193], which, while methodologically distinct, relies on calibration from distance ladder measurements [178, 56, 157].
 | Author | Year | Method | ||
-| Huang [Huang2020] | 2019 | 73.3 | 4.0 | Mira |
-| de Jaeger [deJaeger2020] | 2020 | 75.8 | 5.1 | Cepheids/TRGB + SNe II |
-| Khetan [Khetan2021] | 2021 | 70.5 | 4.1 | SBF |
-| Freedman [Freedman2021] | 2021 | 69.8 | 1.7 | TRGB (CCHP) |
-| Kenworthy [Kenworthy2022] | 2022 | 73.1 | 2.5 | Two-rung ladder |
-| Scolnic [Scolnic2024] | 2023 | 73.22 | 2.06 | TRGB (SH0ES calibrators) |
-| Li [Li2024b] | 2024 | 74.7 | 3.1 | JAGB + SNe Ia |
-| Lee [Lee2024] | 2024 | 67.8 | 2.7 | JAGB + SNe Ia |
-| Jensen [Jensen2025] | 2025 | 73.8 | 2.4 | TRGB |
-| Riess [Riess2025] | 2025 | 73.49 | 0.93 | Cepheids (JWST) |
-| Bhardwaj [Bhardwaj2025] | 2025 | 73.06 | 2.6 | Mira |
-| Newman [Newman2025] | 2025 | 75.3 | 2.9 | TRGB + SNe Ia |
-We note that the Lee [Lee2024] JAGB measurement ( km s-1 Mpc-1) represents a notable outlier within this category, yielding a value consistent with Planck rather than the typical Distance Ladder result.
-This discrepancy with other JAGB analyses (e.g., Li [Li2024b]) appears to originate from differences in the JAGB zero-point calibration and photometric selection criteria, highlighting that even within a single indicator class, methodological choices can significantly impact the inferred .
-II.4 Category 2: Local One-Step (CDM Assumption)
-This category includes 33 measurements that bypass the Distance Ladder (“One-Step”) but infer by explicitly assuming the standard CDM expansion history:
+| Huang [78] | 2019 | 73.3 | 4.0 | Mira |
+| de Jaeger [42] | 2020 | 75.8 | 5.1 | Cepheids/TRGB + SNe II |
+| Khetan [90] | 2021 | 70.5 | 4.1 | SBF |
+| Freedman [63] | 2021 | 69.8 | 1.7 | TRGB (CCHP) |
+| Kenworthy [88] | 2022 | 73.1 | 2.5 | Two-rung ladder |
+| Scolnic [161] | 2023 | 73.22 | 2.06 | TRGB (SH0ES calibrators) |
+| Li [98] | 2024 | 74.7 | 3.1 | JAGB + SNe Ia |
+| Lee [96] | 2024 | 67.8 | 2.7 | JAGB + SNe Ia |
+| Lapuente [155] | 2025 | 72.61 | 1.69 | TRGB+JAGB+Ceph.+SNe Ia (twins) |
+| Jensen [83] | 2025 | 73.8 | 2.4 | TRGB |
+| Riess [149] | 2025 | 73.49 | 0.93 | Cepheids (JWST) |
+| Bhardwaj [20] | 2025 | 73.06 | 2.6 | Mira |
+| Newman [117] | 2025 | 75.3 | 2.9 | TRGB + SNe Ia |
+We note that the Lee [96] JAGB measurement ( km s-1 Mpc-1) represents a notable outlier within this category, yielding a value consistent with Planck rather than the typical Distance Ladder result.
+This discrepancy with other JAGB analyses (e.g., Li, Lapuente [98, 155]) appears to originate from differences in the JAGB zero-point calibration and photometric selection criteria, highlighting that even within a single indicator class, methodological choices can significantly impact the inferred .
+A critical recent addition to this category is the analysis by Ruiz-Lapuente et al.
+[155], which utilizes the “SNe Ia twins” method calibrated via both TRGB and JAGB.
+Notably, their specific calibration using JAGB yields km s-1 Mpc-1, while their TRGB determination gives km s-1 Mpc-1.
+In our compilation (Table 2), we have adopted the value km s-1 Mpc-1, which represents the weighted average across the different methodological combinations presented in their work.
+This independent confirmation of high values using alternative calibrators is particularly significant as it directly challenges previous lower estimates ( km s-1 Mpc-1) historically associated with the CCHP program.
+It suggests that the previously reported discrepancies between Cepheids and other local calibrators of the Distance Ladder may have been overstated, implying that unrecognized systematics in the specific calibration or sample selection of earlier CCHP analyses—rather than the indicators themselves—may account for the lower values.
+Consequently, this result reinforces the internal consistency of Category 1, further isolating the Distance Ladder methodology from the rest of the cosmological probes.
+II.4 Category 2: Local (CDM Assumption)
+This category includes 33 measurements that bypass the Distance Ladder but infer by explicitly assuming the standard CDM expansion history:
 While these methods probe the local Universe (typically ), they rely on this model to map observables—such as time delays, luminosity distances, dispersion measures, or angular diameter distances—to .
@@
 This category includes:
-Gravitational Wave Standard Sirens: Both bright sirens (with EM counterparts, e.g., GW170817 [Palmese2023]) and dark sirens (using galaxy catalogs for redshift inference [Abbott2021, Beirnaert2025]), where the luminosity distance is computed assuming CDM.
-Strong Lensing Time Delays: Analyses where the cosmological model enters both the time-delay distance calculation and often the lens mass modeling [Shajib2023, Kelly2023, Birrer2025].
-Fast Radio Bursts: Measurements using the dispersion measure–redshift relation interpreted under CDM assumptions for the intergalactic medium [Wu2022, PiratovMoreno2025].
 | Author | Year | Method | ||
-| Wu [Wu2022] | 2021 | 68.8 | 4.7 | FRBs |
-| Shajib [Shajib2023] | 2023 | 77.1 | 7.2 | Lensing (TDCOSMO) |
-| Napier [Napier2023] | 2023 | 74.1 | 8.0 | Cluster Lenses |
-| Kelly [Kelly2023] | 2023 | 66.6 | 3.7 | Lensed SN Refsdal |
-| Liu [liu2023h] | 2023 | 59.1 | 3.6 | Cluster Lensed Quasar |
-| Martinez [Martinez2023] | 2023 | 74.0 | 11.2 | Time-Delay Cosmography |
-| Pascale [Pascale2024] | 2024 | 71.8 | 8.7 | Lensed SN H0pe (JWST) |
-| Yang [Yang2024] | 2024 | 74.0 | 7.4 | FRBs w/ Scattering |
-| Piratova [PiratovMoreno2025] | 2025 | 65.1 | 7.4 | FRBs (MLE) |
-| Beirnaert [Beirnaert2025] | 2025 | 79.0 | 9.0 | GW Dark Sirens |
-| Birrer [Birrer2025] | 2025 | 71.6 | 3.6 | TDCOSMO Strong Lensing |
-| Pierel [Pierel2025] | 2025 | 66.9 | 9.8 | SN Encore Time Delays |
-II.5 Category 3: Pure Local One-Step (Model Independent)
-This category represents the cleanest test of the Distance Ladder, as it comprises 15 measurements that determine geometrically or via direct measurement of the expansion rate, without assuming a specific functional form for .
+| Wu [189] | 2021 | 68.8 | 4.7 | FRBs |
+| Shajib [163] | 2023 | 77.1 | 7.2 | Lensing (TDCOSMO) |
+| Napier [116] | 2023 | 74.1 | 8.0 | Cluster Lenses |
+| Kelly [87] | 2023 | 66.6 | 3.7 | Lensed SN Refsdal |
+| Liu [104] | 2023 | 59.1 | 3.6 | Cluster Lensed Quasar |
+| Martinez [113] | 2023 | 74.0 | 11.2 | Time-Delay Cosmography |
+| Pascale [125] | 2024 | 71.8 | 8.7 | Lensed SN H0pe (JWST) |
+| Yang [190] | 2024 | 74.0 | 7.4 | FRBs w/ Scattering |
+| Piratova [140] | 2025 | 65.1 | 7.4 | FRBs (MLE) |
+| Beirnaert [18] | 2025 | 79.0 | 9.0 | GW Dark Sirens |
+| Birrer [175] | 2025 | 71.6 | 3.6 | TDCOSMO Strong Lensing |
+| Pierel [138, 174] | 2025 | 66.9 | 9.8 | SN Encore Time Delays |
+II.5 Category 3: Pure Local (Cosmological-Model Independent)
+This category represents the cleanest test of the Distance Ladder, as it comprises 16 measurements that determine geometrically or via direct measurement of the expansion rate, without assuming a specific functional form for .
 These measurements are either:
-Purely geometric: Relying on Keplerian dynamics or light-travel-time arguments that are independent of the global cosmological model (e.g., Megamaser disk modeling [Kuo2013, Gao2015]).
-Direct measurements: Cosmic chronometers measure directly from the differential aging of massive, passively evolving galaxies, providing model-independent determinations of the expansion rate at various redshifts [Moresco2023].
-To extract , two main approaches have been employed: (i) Gaussian Process reconstruction of combined with a cosmographic expansion of the luminosity distance in the Hubble flow (), (ii) using cosmic chronometers to calibrate the H ii galaxies Hubble diagram, which in turn anchors Type Ia supernovae at low redshift via Hubble’s law [Zhang2022].
+Direct measurements: Cosmic chronometers measure directly from the differential aging of massive, passively evolving galaxies, providing cosmological-model-independent determinations of the expansion rate at various redshifts [115].
+To extract , two main approaches have been employed: (i) Gaussian Process reconstruction of combined with a cosmographic expansion of the luminosity distance in the Hubble flow (), (ii) using cosmic chronometers to calibrate the H ii galaxies Hubble diagram, which in turn anchors Type Ia supernovae at low redshift via Hubble’s law [192].
 Both methods avoid assumptions about the underlying cosmological model beyond the FLRW metric.
-Cosmographic approaches: Analyses that expand or in Taylor series around , avoiding assumptions about the energy content of the Universe [Du2023].
-Expanding photosphere methods: Geometric distance determinations from supernovae using photospheric velocity and angular size evolution [Sneppen2023, Vogl2024].
+Cosmographic approaches: Analyses that expand or in Taylor series around , avoiding assumptions about the energy content of the Universe [55].
 The key distinction from Category 2 is that these analyses do not assume the validity of Eq.
@@
 | Author | Year | Method | ||
-| Kuo [Kuo2013] | 2013 | 68.0 | 9.0 | Megamasers (MCP) |
-| Gao [Gao2015] | 2015 | 66.0 | 6.0 | Megamasers |
-| Zhang [Zhang2022] | 2022 | 65.9 | 3.0 | CC + HII |
-| Du [Du2023] | 2023 | 71.5 | 3.8 | TD Lenses + GRBs |
-| Sneppen [Sneppen2023] | 2023 | 67.0 | 3.6 | EPM + Kilonovae |
-| Li [Li2024a] | 2024 | 66.3 | 3.7 | Lensing Bias Correction |
-| Vogl [Vogl2024] | 2024 | 74.9 | 2.7 | SNe II (Tailored EPM) |
-| Song [Song2025] | 2025 | 70.4 | 6.9 | GW + SGL Calibration |
-| Favale [Favale2025] | 2025 | 68.8 | 3.0 | Cosmic Chronometers |
+| Kuo [95] | 2013 | 68.0 | 9.0 | Megamasers (MCP) |
+| Gao [66] | 2015 | 66.0 | 6.0 | Megamasers |
+| Zhang [192] | 2022 | 65.9 | 3.0 | CC + HII |
+| Du [55] | 2023 | 71.5 | 3.8 | TD Lenses + GRBs |
+| Sneppen [171] | 2023 | 67.0 | 3.6 | EPM + Kilonovae |
+| Li [101] | 2024 | 66.3 | 3.7 | Lensing Bias Correction |
+| Vogl [182] | 2024 | 74.9 | 2.7 | SNe II (Tailored EPM) |
+| Song [172] | 2025 | 70.4 | 6.9 | GW + SGL Calibration |
+| Favale [60] | 2025 | 68.8 | 3.0 | Cosmic Chronometers |
 We highlight that this category exhibits significant internal heterogeneity.
-Notably, the Vogl [Vogl2024] tailored EPM measurement ( km s-1 Mpc-1) strongly favor the Distance Ladder value, while most cosmic chronometer measurements favor lower values.
+Notably, the Vogl [182] tailored EPM measurement ( km s-1 Mpc-1) strongly favor the Distance Ladder value, while most cosmic chronometer measurements favor lower values.
 This heterogeneity will be discussed further in Section IV.
-II.6 Category 4: CMB Sound Horizon Free
-The final category includes 6 measurements that utilize early-Universe data (primarily CMB) but specifically avoid the Sound Horizon as a standard ruler.
+II.6 Category 4: CMB Sound-Horizon-Free (CMB–SHF)
+The final category includes 9 measurements that utilize early-Universe data (primarily CMB) but specifically avoid the Sound Horizon as a standard ruler.
 Instead, these methods rely on alternative physical scales that are calibrated independently of pre-recombination acoustic physics:
-Matter-radiation equality scale (): The comoving wavenumber corresponding to the horizon size at matter-radiation equality provides a standard ruler that depends on and but is independent of [Philcox2022, Bahr-Kalus2025].
-BAO without priors: Analyses that use BAO peak positions but treat as a free parameter, constraining through the shape of rather than absolute calibration [Pogosian2024].
-Sunyaev-Zel’dovich (SZ) effect: Combining SZ decrements with X-ray observations of galaxy clusters to infer angular diameter distances independent of [Reese2003, Colaco2023].
+BAO without priors: Analyses that use BAO peak positions but treat as a free parameter, constraining through the shape of rather than absolute calibration [144].
 | Author | Year | Method | ||
-| Reese [Reese2003] | 2003 | 61.0 | 18.0 | SZ Effect |
-| Colaco [Colaco2023] | 2023 | 67.2 | 6.1 | SZ + X-ray |
-| Pogosian [Pogosian2024] | 2024 | 69.5 | 0.94 | BAO-DESI (No ) |
-| Pogosian [Pogosian2024] | 2024 | 68.0 | 0.94 | BAO-pre DESI (No ) |
-| Bahr-Kalus [Bahr-Kalus2025] | 2025 | 65.2 | 5.6 | DESI Turnover Scale |
-| Krolewski [Krolewski2025] | 2025 | 69.0 | 2.5 | DESI + CMB (No ) |
-We note that this category includes two entries from Pogosian et al.
-[Pogosian2024], corresponding to analyses using DESI and pre-DESI BAO data respectively.
-While these share a common methodology, they utilize substantially different datasets and thus provide partially independent constraints.
-However, the very low reduced chi-squared () for this category suggests that uncertainties may be conservative or that some correlation exists between measurements.
-We address this in the correlation analysis of Appendix C.
+| Reese [147] | 2003 | 61.0 | 18.0 | SZ Effect |
+| Philcox [137] | 2022 | 64.8 | BAO scale,Matter–rad. eq. | |
+| Colaco [36] | 2023 | 67.2 | 6.1 | SZ + X-ray |
+| Pogosian [144] | 2024 | 68.05 | 0.94 | BAO-pre DESI (No ) |
+| Pogosiana [145] | 2025 | 69.37 | 0.65 | BAO-DESI DR2 (No ) |
+| Bahr-Kalus [13] | 2025 | 65.2 | 5.6 | DESI Turnover Scale |
+| Garcia [70] | 2025 | 70.03 | 0.97 | BAO-DESI DR2 + (No ) |
+| Zaborowski [191] | 2025 | 70.8 | BAO-DESI + (No ) | |
+| Krolewski [93] | 2025 | 69.0 | 2.5 | DESI + CMB (No ) |
+a
+Private communication.
 II.7 Summary of Classification
@@
 | Cat. | Description | Uses Ladder? | Assumes CDM? | |
-| 1 | Distance Ladder | 29 | Yes | No |
-| 2 | Local One-Step (CDM) | 33 | No | Yes |
-| 3 | Pure Local One-Step | 15 | No | No |
-| 4 | CMB Sound Horizon Free | 6 | No | Yes∗ |
+| 1 | Distance Ladder | 30 | Yes | No |
+| 2 | Local (CDM) | 33 | No | Yes |
+| 3 | Pure Local | 16 | No | No |
+| 4 | CMB–SHF | 9 | No | Yes∗ |
 ∗Category 4 assumes CDM for shape but avoids the Sound Horizon .
@@
 We utilize the inverse-variance weighted mean to determine the central value for each category and the reduced chi-squared statistic () to assess internal consistency.
-We explicitly quantify the tension between the Distance Ladder (Category 1) and the various subsets of Sound-Horizon-independent measurements, concluding with a comparison against the combined ensemble of all One-Step methods.
-We also compare our results with our previous analysis [Perivolaropoulos2024] to assess the evolution of the tension landscape.
+We explicitly quantify the tension between the Distance Ladder (Category 1) and the various subsets of Sound-Horizon-Free measurements, concluding with a comparison against the combined ensemble of all Distance Ladder Independent/Sound Horizon Free (DLI–SHF) methods.
+We also compare our results with our previous analysis [134] to assess the evolution of the tension landscape.
 III.1 Methodology
-For a dataset of independent measurements , the weighted mean and its standard error are calculated as:
+For a set of independent measurements , the weighted mean and its standard error are calculated as:
 where the weights are given by . This estimator minimizes the variance for Gaussian-distributed errors.
 To evaluate whether the variation within a category is consistent with the reported uncertainties, we compute the reduced chi-squared:
+In cases of asymmetric errors, the following Equation is adopted for computations involving the mean or a unified error value:
 The interpretation of provides important diagnostic information:
@@
 The statistical results for the four categories defined in Section II are summarized in Table 7.
-For comparison, we also include the results from our previous analysis [Perivolaropoulos2024], which employed a binary classification scheme (Distance Ladder vs.
-One-Step).
+For comparison, we also include the results from our previous analysis [134], which employed a binary classification scheme (Distance Ladder vs.
+Distnance Ladder Independent/Sound Horizon Free).
 | Category | Tension | |||
-| 1. Distance Ladder | 29 | 0.74 | — | |
-| 2. Local One-Step (CDM) | 33 | 0.85 | ||
-| 3. Pure Local One-Step | 15 | 0.78 | ||
-| 4. CMB Sound Horizon Free | 6 | 0.36 | ||
-| Combined (2+3+4) | 54 | 0.85 | ||
-| Corr.-adjusted (2+3+4) | 54 | — | ||
-| Comparison with Perivolaropoulos (2024) [Perivolaropoulos2024]: | ||||
+| 1. Distance Ladder | 30 | 0.72 | — | |
+| 2. Local (CDM) | 33 | 0.80 | ||
+| 3. Pure Local | 16 | 0.77 | ||
+| 4. CMB–SHF | 9 | 0.89 | ||
+| Combined (2+3+4) | 58 | 0.95 | ||
+| Corr.-adjusted (2+3+4) | 58 | — | ||
+| Comparison with Perivolaropoulos (2024) [134]: | ||||
 | 2024 Distance Ladder | 20 | 0.51 | — | |
-| 2024 One-Step | 33 | 0.95 | ||
+| 2024 DLI–SHF | 33 | 0.95 | ||
 To visualize the distinct behaviors of these categories, we present the Gaussian probability distributions in Figure 2.
@@
 III.2.1 Category 1: Distance Ladder ()
-The 29 measurements comprising the Distance Ladder category yield a weighted mean of km s-1 Mpc-1.
+The 30 measurements comprising the Distance Ladder category yield a weighted mean of km s-1 Mpc-1.
 As shown in Figure 2, this category forms a sharp, narrow peak that is clearly separated from the other distributions.
 The reduced chi-squared indicates good internal consistency, with the scatter being slightly smaller than expected from the reported uncertainties.
-This value is notably higher than our 2024 analysis (), suggesting that the expanded dataset—which now includes measurements with greater methodological diversity (e.g., JAGB, Miras, pairwise velocities)—exhibits more realistic scatter.
+This value is notably higher than our 2024 analysis (), suggesting that the expanded data—which now includes measurements with greater methodological diversity (e.g., JAGB, Miras, pairwise velocities)—exhibits more realistic scatter.
 Nevertheless, may indicate that: (i) reported uncertainties are somewhat conservative, (ii) positive correlations exist between measurements sharing common calibration anchors, or (iii) there is a tendency for published results to cluster around previously established values.
-The high precision of this combined estimate is driven by the recent JWST Cepheid measurements [Riess2025] and compatible results from other calibrators including TRGB [Jensen2025], Miras [Bhardwaj2025], and SBF [Blakeslee2021].
+The high precision of this combined estimate is driven by the recent JWST Cepheid measurements [149] and compatible results from other calibrators including TRGB [83], Miras [20], and SBF [22].
 We note that the Lee et al.
-[Lee2024] JAGB measurement () represents a outlier within this category, as discussed in Section II.
-III.2.2 Category 2: Local One-Step CDM ()
+[96] JAGB measurement () represents a outlier within this category, as discussed in Section II.
+III.2.2 Category 2: Local CDM ( km s-1 Mpc-1)
 The 33 measurements in this category, which utilize local data but assume the CDM expansion history, yield the lowest aggregate value of km s-1 Mpc-1.
@@
 This remarkable alignment between local CDM-dependent measurements and early-universe physics—achieved without using the Sound Horizon—is one of the central findings of this work.
-III.2.3 Category 3: Pure Local One-Step ()
-The 15 model-independent measurements (Megamasers, Cosmic Chronometers, EPM) yield an intermediate value of km s-1 Mpc-1.
+III.2.3 Category 3: Pure Local ()
+The 16 cosmological-model-independent measurements (Megamasers, Cosmic Chronometers, EPM) yield an intermediate value of km s-1 Mpc-1.
 Crucially, the removal of the CDM shape assumption results in a central value shift of km s-1 Mpc-1 relative to Category 2.
@@
 This category exhibits significant internal heterogeneity: the Vogl et al.
-[Vogl2024] tailored EPM measurement () strongly favors the Distance Ladder value, while most Cosmic Chronometer measurements favor values closer to Planck.
+[182] tailored EPM measurement () strongly favors the Distance Ladder value, while most Cosmic Chronometer measurements favor values closer to Planck.
 This heterogeneity suggests that Category 3 may itself contain distinct subpopulations, a possibility we explore further in Section IV.
-III.2.4 Category 4: CMB Sound Horizon Free ()
-The 6 measurements utilizing CMB data without the sound horizon yield km s-1 Mpc-1.
-This confirms that early-universe constraints on matter density (via the equality scale ) favor a low regardless of the calibration [Pogosian2024].
-The very low reduced chi-squared warrants careful interpretation.
-With only measurements, statistical fluctuations in are expected to be large (the standard deviation of scales as for ).
-Nevertheless, this low value likely reflects: (i) the dominance of the two high-precision Pogosian et al.
-[Pogosian2024] measurements, which share a common methodology and partially overlapping datasets; and (ii) the fact that all measurements in this category are ultimately constrained by similar early-universe physics.
-In Figure 2, this distribution overlaps significantly with Category 2, reinforcing the consistency between early-universe physics and local data when CDM is assumed.
-III.3 Comparison: Distance Ladder vs. Combined One-Step
+III.2.4 Category 4: CMB Sound Horizon Free (CMB–SHF)()
+The 9 measurements utilizing CMB data without the Sound Horizon yield km s-1 Mpc-1.
+This confirms that early-universe constraints on matter density (via the equality scale ) favor a low regardless of the calibration [144].
+The reduced chi-squared is very close to unity, indicating good internal consistency among the measurements in this category.
+However, given that several measurements share common methodologies and partially overlapping datasets—particularly the Pogosian et al.
+[144, 145] and García Escudero et al.
+[70] results, which all employ Sound-Horizon-Free BAO analyses with DESI data—we examine potential correlations in Appendix C.
+In Figure 2, this distribution overlaps significantly with CMB–SHF measurements, reinforcing the consistency between early-universe physics and local data when CDM is assumed.
+III.3 Comparison: Distance Ladder vs. Combined Distance Ladder Independent/Sound Horizon Free (DLI–SHF)
 To test the robustness of the “Early vs.
-Late” narrative, we combine all 54 Sound-Horizon-independent measurements from Categories 2, 3, and 4 into a single “All One-Step” ensemble.
+Late” narrative, we combine all 58 DLI–SHF from Categories 2, 3, and 4 into a single ensemble.
 This combined sample represents the entirety of the evidence against the Distance Ladder that does not rely on the standard ruler.
@@
 This distinction explains why the curves overlap visually even when the weighted means differ by several standard deviations.
-The tension between the Distance Ladder and the Combined One-Step sample is:
+The tension between the Distance Ladder and the Combined DLI–SHF sample is:
 Combining the uncertainties in quadrature ( km s-1 Mpc-1), the significance of the discrepancy is:
@@
 III.3.1 Comparison with Previous Results
-Table 7 includes a comparison with our 2024 analysis [Perivolaropoulos2024].
-Despite the significant expansion of the dataset (from 53 to 83 measurements) and the refinement of the classification scheme (from 2 to 4 categories), the core results remain consistent:
+Table 7 includes a comparison with our 2024 analysis [134].
+Despite the significant expansion of the sample (from 53 to 88 measurements) and the refinement of the classification scheme (from 2 to 4 categories), the core results remain consistent:
 The Distance Ladder weighted mean has shifted slightly from to km s-1 Mpc-1, with improved precision due to the inclusion of JWST measurements.
-The One-Step weighted mean has shifted from to km s-1 Mpc-1, remaining consistent within uncertainties.
+The DLI–SHF weighted mean has shifted from to km s-1 Mpc-1, remaining consistent within uncertainties.
 The tension has increased from to , primarily due to reduced uncertainties rather than a change in central values.
@@
 As detailed in Appendix C, accounting for realistic correlations within each category increases the combined uncertainties.
-Using correlation coefficients estimated from shared calibration data and common methodology (– for Distance Ladder; – for One-Step), we obtain:
+Using correlation coefficients estimated from shared calibration data and common methodology (– for Distance Ladder; – for DLI–SHF), we obtain:
 The correlation-adjusted tension is:
@@
 III.4 Internal Tension: Model Dependence
-Beyond the external tension with the Distance Ladder, we identify a notable internal tension among the One-Step measurements based on model assumptions.
-Comparing Category 2 (Local One-Step CDM) with Category 3 (Pure Local One-Step), we find:
+Beyond the external tension with the Distance Ladder, we identify a notable internal tension among the Distance Ladder Independent/Sound Horizon Free(DLI–SHF) measurements based on model assumptions.
+Comparing Category 2 (Local CDM) with Category 3 (Pure Local), we find:
 The statistical significance is:
@@
 3.
-Sample size: Category 3 contains only 15 measurements, making the weighted mean sensitive to individual high-precision entries.
+Sample size: Category 3 contains only 16 measurements, making the weighted mean sensitive to individual high-precision entries.
 III.5 Summary of Statistical Results
@@
 2.
-The combined One-Step measurements yield km s-1 Mpc-1, consistent with Planck.
+The combined DLI–SHF measurements yield km s-1 Mpc-1, slightly elevated compared to our previous results but closer to Planck.
 3.
@@
 4.
-Local measurements assuming CDM (Category 2) yield systematically lower than model-independent local measurements (Category 3), with a internal tension.
+Local measurements assuming CDM (Category 2) yield systematically lower than cosmological-model-independent local measurements (Category 3), with a internal tension.
 5.
@@
 IV The Impact of Model Assumptions
-The results of Section III reveal a striking feature of the Hubble tension landscape: the discrepancy is not merely between the Distance Ladder and the Early Universe, but also manifests as an internal tension within the domain of One-Step measurements.
+The results of Section III reveal a striking feature of the Hubble tension landscape: the discrepancy is not merely between the Distance Ladder and the Early Universe, but also manifests as an internal tension within the domain of Distance Ladder Independent/Sound Horizon Free (DLI–SHF) measurements.
 Specifically, the statistical compatibility of a measurement with the Planck value depends critically on whether the standard CDM expansion history is assumed during the inference process.
@@
 IV.1 Implicit Bias of the CDM Expansion History
-The most significant finding of our four-category classification is the tension between local One-Step measurements that assume a CDM cosmology (Category 2: km s-1 Mpc-1) and those that do not (Category 3: km s-1 Mpc-1).
-Measurements in Category 2, such as Gravitational Wave standard sirens [Abbott2021] and standard analyses of Strong Lensing time delays [Shajib2023], typically span the redshift range .
+The most significant finding of our four-category classification is the tension between local DLI–SHF measurements that assume a CDM cosmology (Category 2: km s-1 Mpc-1) and those that do not (Category 3: km s-1 Mpc-1).
+Measurements in Category 2, such as Gravitational Wave standard sirens [102] and standard analyses of Strong Lensing time delays [163], typically span the redshift range .
 To extract from the observed luminosity distances or time delays, these methods must integrate the Hubble parameter :
@@
 The inferred is then the normalization that best fits the data given this assumed shape.
-Our analysis shows that this CDM shape constraint systematically prefers lower values ( km s-1 Mpc-1), aligning almost perfectly with the Planck CMB inference [Planck2018].
-This demonstrates a remarkable self-consistency of the CDM model: if one assumes the CDM background evolution holds true at late times, local data (excluding the Distance Ladder) yields results consistent with early-universe physics, even without using the Sound Horizon.
+Our analysis shows that this CDM shape constraint systematically prefers lower values ( km s-1 Mpc-1), aligning almost perfectly with the Planck CMB inference [141].
+This demonstrates a remarkable self-consistency of the CDM model: if one assumes the CDM background evolution holds true at late times, local data (excluding the Distance Ladder) yields results consistent with early-universe physics, even without using the Sound Horizon [16].
 However, when this shape constraint is released (Category 3), the inferred rises by km s-1 Mpc-1 to km s-1 Mpc-1. This shift has two possible interpretations:
 1.
-CDM is correct, and Category 3 has larger systematics: Model-independent methods may suffer from unrecognized systematic errors that bias high.
+CDM is correct, and Category 3 has larger systematics: Cosmological-model-independent methods may suffer from unrecognized systematic errors that bias high.
 The lack of a shape constraint allows more freedom for systematics to propagate into the final result.
 2.
-The true deviates from CDM: If the actual expansion history differs from the CDM prediction at (e.g., due to a late-time gravitational transition, dynamical dark energy, or modified gravity [Perivolaropoulos2024]), forcing a CDM fit would bias the inferred downward to compensate for the shape mismatch.
+The true deviates from CDM: If the actual expansion history differs from the CDM prediction at (e.g., due to a late-time gravitational transition, dynamical dark energy, or modified gravity [134, 77, 142, 119, 123, 7]), forcing a CDM fit would bias the inferred downward to compensate for the shape mismatch.
 A related possibility is that the tension reflects not a global modification of , but rather a localized change in the physics governing distance calibrators.
-Refs [Desmond:2019ygn, Desmond:2020wep, Marra2021, Ruchika2025, Perivolaropoulos:2025gzo] investigated whether a local variation in fundamental physics—such as a change in the fine-structure constant or gravitational strength within the calibration volume—could reconcile the SH0ES measurement with other probes without requiring modifications to the global expansion history.
-The current data do not definitively distinguish between these scenarios.
+Refs [44, 45, 112, 154, 128] investigated whether a local variation in fundamental physics—such as a change in the fine-structure constant or gravitational strength within the calibration volume—could reconcile the SH0ES measurement with other probes without requiring modifications to the global expansion history.The current data do not definitively distinguish between these scenarios.
+A detailed comparison of late-transition versus smooth H(z) deformation models has been presented in [8].
+Earlier reanalyses of SH0ES data exploring additional degrees of freedom also found evidence for such model-dependent effects [130].
 IV.2 The Vogl et al. Tailored EPM: A Critical Test Case
 Among the Category 3 measurements, the Vogl et al.
-[Vogl2024] tailored Expanding Photosphere Method (EPM) result deserves particular attention.
-This measurement yields km s-1 Mpc-1 (with systematic uncertainties — by the authors own estimate comparable to the statistical errors — taken into account) is fully consistent with the Distance Ladder and in tension with the Planck value—yet it is completely independent of both the Distance Ladder calibration chain and cosmological model assumptions.
+[182] tailored Expanding Photosphere Method (EPM) result deserves particular attention.
+This measurement yields km s-1 Mpc-1 (with systematic uncertainties — by the authors’ own estimate comparable to the statistical errors — taken into account) is fully consistent with the Distance Ladder and in tension with the Planck value—yet it is completely independent of both the Distance Ladder calibration chain and cosmological model assumptions.
 The EPM determines distances to Type II supernovae geometrically by combining measurements of the photospheric angular size (from flux and temperature) with the photospheric velocity (from spectral line Doppler shifts).
@@
 The Barua et al.
-[Barua2025] frequentist analysis of the Megamaser Cosmology Project (MCP) sample yields km s-1 Mpc-1, again consistent with the Distance Ladder.
+[17] frequentist analysis of the Megamaser Cosmology Project (MCP) sample yields km s-1 Mpc-1, again consistent with the Distance Ladder.
 Megamaser distances are determined geometrically from the Keplerian dynamics of circumnuclear water maser disks, providing angular diameter distances that are in principle completely independent of stellar physics.
 However, Megamaser measurements are sensitive to the treatment of peculiar velocities, which become significant at the low redshifts () where most masers hosts reside.
-In our classification, we assigned the Barua et al [Barua2025] measurement to Category 2 because it relies primarily on Keplerian disk modeling without adopting CDM-based peculiar velocity priors but relies on the CDM prediction for the angular diameter distance.
+In our classification, we assigned the Barua et al [17] measurement to Category 2 because it relies primarily on Keplerian disk modeling without adopting CDM-based peculiar velocity priors but relies on the CDM prediction for the angular diameter distance.
 Indeed, when is left as a free parameter rather than fixed to the Planck value, the best-fit drops to km s-1 Mpc-1 with uncertainties approximately five times larger, highlighting the model dependence of the headline result.
 In contrast, earlier analyses that adopted peculiar velocity corrections from the SH0ES flow model from Pesce et al.
-[Pesce2020] would be more appropriately classified as having some distance ladder dependence.
+[136] would be more appropriately classified as having some distance ladder dependence.
 The earlier Megamaser measurements by Kuo et al.
-[Kuo2013] ( km s-1 Mpc-1) and Gao et al.
-[Gao2015] ( km s-1 Mpc-1) yielded lower values, though with larger uncertainties.
+[95] ( km s-1 Mpc-1) and Gao et al.
+[66] ( km s-1 Mpc-1) yielded lower values, though with larger uncertainties.
 The evolution toward higher in recent Megamaser analyses may reflect improved disk modeling, better peculiar velocity treatments, or statistical fluctuations in small samples.
-IV.4 Bimodality in the One-Step Distribution
-Although Figure 3 does not exhibit an obvious bimodal structure—the distribution of One-Step measurements appears reasonably concentrated around a single peak—a more detailed examination reveals a subtle but significant effect.
-Upon closer inspection, we found that removing even two high-precision measurements (the Pogosian et al.
-[Pogosian2024] DESI and pre-DESI results) substantially alters the overall shape of the distribution.
-These measurements, with uncertainties of km s-1 Mpc-1, are among the most precise in our compilation.
-Due to their small uncertainties, they exert disproportionate influence on the inverse-variance weighted distribution, effectively pulling the combined probability density upward at their location (– km s-1 Mpc-1).
-When these high-weight measurements are excluded, the inverse-variance weighted sum of Gaussians transforms into a distinctly bimodal curve shown in Figure 4.
-This suggests that the apparent unimodality in the raw distribution may be an artifact of a small number of highly precise measurements rather than a reflection of the true underlying structure.
-We emphasize that this analysis should be interpreted with caution.
-Selectively removing measurements to reveal bimodality could be seen as a form of data manipulation.
-However, our purpose here is not to argue that the Pogosian et al.
-measurements are incorrect, but rather to illustrate that the statistical structure of the distribution is sensitive to a small number of high-precision anchors.
-A more rigorous approach would involve mixture model analysis (e.g., Gaussian Mixture Models) to formally test for bimodality, which we leave for future work.
+IV.4 Bimodality in the Distance Ladder Independent/Sound Horizon Free (DLI–SHF) Distribution
+Although Figure 3 does not exhibit an obvious bimodal structure—the distribution of DLI–SHF measurements appears reasonably concentrated around a single peak—a more detailed examination reveals a subtle but significant effect.
+To further explore the underlying structure of the DLI–SHF measurements, demonstrated also in figure 2, we employ a weighted Kernel Density Estimation (KDE) approach.
+Unlike a simple inverse-variance weighted sum of Gaussians, where each measurement contributes a Gaussian with width equal to its uncertainty , the KDE method assigns a fixed bandwidth to all measurements while weighting their contributions according to their precision.
+The weighted KDE is defined as:
+where is the measured value, is the bandwidth (kernel width), and are the normalized inverse-variance weights:
+This approach ensures that high-precision measurements contribute proportionally through their weights, but do not dominate the distribution through both weight and narrow kernel width simultaneously—a problem that arises with the standard inverse-variance weighted sum of Gaussians.
+For this analysis, we exclude Category 4 measurements (CMB Sound Horizon Free), as these rely on CMB-derived data that provide exceptionally high precision.
+The three most precise Category 4 measurements—Pogosian et al.
+[144, 145] (DESI: km s-1 Mpc-1; pre-DESI: km s-1 Mpc-1) and Garcia et al.
+[70] ( km s-1 Mpc-1)—would alone account for approximately of the total inverse-variance weight, clustering near km s-1 Mpc-1 and obscuring any underlying structure in the remaining measurements.
+The bandwidth is a critical parameter controlling the smoothness of the KDE.
+Silverman’s rule of thumb [167], which minimizes the mean integrated squared error assuming a unimodal Gaussian distribution, yields km s-1 Mpc-1 for our data.
+However, this criterion is known to oversmooth multimodal distributions.
+Given the a priori distinction between Category 2 (CDM-dependent) and Category 3 (model-independent) measurements, a smaller bandwidth is justified to reveal potential substructure.
+We find that bimodal structure is clearly visible for bandwidths in the range km s-1 Mpc-1.
+For , additional peaks emerge due to over-resolution of statistical fluctuations; for , the two peaks merge into a single mode with a shoulder.
+We adopt km s-1 Mpc-1, an intermediate value within the robust range, to balance resolution and smoothing.
+The resulting distribution, shown in Figure 4, exhibits two distinct peaks corresponding to the weighted means of Category 2 ( km s-1 Mpc-1) and Category 3 ( km s-1 Mpc-1) measurements, with a separation of km s-1 Mpc-1.
+This bimodal structure suggests a systematic difference between CDM-dependent and cosmological-model-independent not Distance Ladder methods that warrants further investigation.
 IV.5 Physical Interpretation of the Bimodality
@@
 Category 2 measurements that assume CDM (FRBs, GW sirens, most Strong Lensing)
-Category 4 measurements anchored to early-universe physics (-based BAO)
 Cosmic Chronometer measurements in Category 3
@@
 - 2.
-The High- Peak ( km s-1 Mpc-1): Populated by:
-Specific measurements: Megamasers [Barua2025] and tailored EPM [Vogl2024]
-These methods are characterized by their reliance on local geometric or kinematic physics.
-This bifurcation pattern suggests a reformulation of the Hubble tension: the discrepancy may fundamentally be between model-dependent methods (which yield low ) and local astrophysics dependent or model-independent methods (which yield high ), rather than between early and late times.
+This bifurcation pattern suggests a reformulation of the Hubble tension: the discrepancy may fundamentally be between model-dependent methods (which yield low ) and local astrophysics dependent or cosmological-model-independent methods (which yield high ), rather than between early and late times.
 IV.6 The Role of High-Precision Anchors
@@
 High- anchors:
-Riess et al. [Riess2025] JWST Cepheids: km s-1 Mpc-1
-Zhang et al. [Zhang2024] pairwise velocities: km s-1 Mpc-1
-Vogl et al. [Vogl2024] tailored EPM: km s-1 Mpc-1
 Low- anchors:
-Pogosian et al. [Pogosian2024] DESI BAO (no ): km s-1 Mpc-1
-Piratova-Moreno et al. [PiratovMoreno2025] FRBs: km s-1 Mpc-1
-Kelly et al. [Kelly2023] SN Refsdal: km s-1 Mpc-1
+Multi-parameter transition analyses of SH0ES data have explored how calibrator parameters may vary across the distance ladder, potentially contributing to the observed bifurcation.
+Evidence for possible transitions in Cepheid-SNe Ia calibrator parameters has been previously identified in the SH0ES sample [129, 164].
+Recent work has also highlighted the subtle statistical effects in Distance Ladder analyses, including distance priors and selection effects, that could impact such comparisons [46].
 The weighted means of each category are heavily influenced by these anchor measurements.
@@
 IV.7 Category 3 as the Critical Adjudication Ground
-The “Pure Local One-Step” category (Category 3) serves as the critical test for distinguishing between the two interpretations of the Hubble tension.
+The “Pure Local category (Category 3) serves as the critical test for distinguishing between the two interpretations of the Hubble tension.
 Currently, this category yields an intermediate value ( km s-1 Mpc-1), but it is internally heterogeneous:
 High- subset: Vogl EPM (), Palmese GW170817 (75.46±5.4)
-Low- subset: Zhang CC+HII (), Moresco CC (), Li lensing ()
+Low- subset: Zhang CC+HII (), Li lensing ()
 The future evolution of Category 3 will be decisive:
@@
 If Category 3 converges toward Category 1 (Distance Ladder): This would strongly imply that the CDM model assumption used in Category 2 is the source of the bias.
-Such convergence would suggest that the true late-time expansion history deviates from CDM, requiring new physics (e.g., dynamical dark energy, modified gravity, or a late-time gravitational transition)such as the CDM model [akarsu2024, Akarsu2025]
+Such convergence would suggest that the true late-time expansion history deviates from CDM, requiring new physics (e.g., dynamical dark energy, modified gravity, or a late-time gravitational transition)such as the CDM model [6, 5]
 2.
@@
 Currently, the intermediate position of Category 3, combined with its internal heterogeneity, does not clearly favor either scenario.
-However, the existence of multiple independent model-independent measurements (EPM, Megamasers) that support the high- solution is suggestive.
+However, the existence of multiple independent cosmological-model-independent measurements (EPM, Megamasers) that support the high- solution is suggestive.
 IV.8 Summary: The Model-Dependence Effect
@@
 3.
-The One-Step distribution shows evidence for bimodality when high-precision anchoring measurements are removed, with peaks corresponding to model-dependent () and model-independent () methods.
+The distribution shows evidence for bimodality with peaks corresponding to model-dependent () and cosmological-model-independent () methods.
 4.
-Specific model-independent measurements (Vogl EPM) yield high values consistent with the Distance Ladder, suggesting the high value may reflect true local physics rather than distance ladder systematics.
+Specific cosmological-model-independent measurements (Vogl EPM) yield high values consistent with the Distance Ladder, suggesting the high value may reflect true local physics rather than distance ladder systematics.
 5.
@@
 The Rest” crisis, with a secondary “Model-Dependent vs.
-Model-Independent” tension that may hold the key to its ultimate resolution.
+Cosmological-model-Independent” tension that may hold the key to its ultimate resolution.
 V Discussion and Conclusions
 The “Hubble Tension” has long been framed as a cosmological crisis pitting the Early Universe (calibrated by the Sound Horizon) against the Late Universe (measured by the distance ladder).
-In this work, we have dissected this narrative by analyzing 83 measurements of that are strictly independent of the CMB Sound Horizon scale.
+In this work, we have dissected this narrative by analyzing 88 measurements of that are strictly independent of the CMB Sound Horizon scale.
 Our findings suggest that the crisis is not necessarily a temporal one, but rather a methodological one.
@@
 As detailed in Section III, the Distance Ladder (Category 1) yields a robust high value of km s-1 Mpc-1.
-In stark contrast, when we combine all 54 independent measurements from Categories 2, 3, and 4—comprising Gravitational Waves, Strong Lensing, Cosmic Chronometers, Megamasers, FRBs, and sound-horizon-free CMB data—we obtain a combined value of km s-1 Mpc-1.
-The resulting tension ( after correlation adjustment) between the Distance Ladder and this “All One-Step” ensemble exceeds the standard Planck vs.
+In stark contrast, when we combine all 58 independent measurements from Categories 2, 3, and 4—comprising Gravitational Waves, Strong Lensing, Cosmic Chronometers, Megamasers, FRBs, and Sound-Horizon-Free CMB data—we obtain a combined value of km s-1 Mpc-1.
+The resulting tension ( after correlation adjustment) between the Distance Ladder and the Distance Ladder Independent/Sound Horizon Free (DLI–SHF) ensemble exceeds the standard Planck vs.
 SH0ES discrepancy ().
@@
 It persists when correlations within categories are accounted for (Section III.3).
-It is consistent with our 2024 analysis [Perivolaropoulos2024], despite the significant expansion of the dataset.
+It is consistent with our 2024 analysis [134], despite the significant expansion of the sample.
 It holds regardless of which specific measurements are included or excluded (with the exception of complete removal of entire methodological classes).
-V.2 Local Physics Does Not Imply High
-A pervasive misconception in the field is that “local” measurements universally favor high values, thereby pointing to a breakdown of CDM at late times.
-Our analysis decisively refutes this assumption.
-We found that 33 measurements utilizing exclusively local data () but assuming the validity of the CDM expansion history (Category 2) yield a weighted mean of km s-1 Mpc-1.
-This value is in perfect agreement with the Planck 2018 result ( km s-1 Mpc-1) [Planck2018].
-Local probes such as Gravitational Wave standard sirens [Abbott2021], Strong Lensing time delays [Shajib2023, Kelly2023, Birrer2025], and FRBs [Wu2022, PiratovMoreno2025] do not intrinsically favor high .
-When analyzed under the standard model framework, local data supports the early-universe inference.
-Furthermore, Category 4 demonstrates that even without the Sound Horizon , CMB data constraints (driven by the matter-radiation equality scale ) favor a low km s-1 Mpc-1 [Pogosian2024, Bahr-Kalus2025].
-Thus, neither “Local” nor “Sound Horizon free” is synonymous with “High .” The high value may be specific to the Distance Ladder methodology.
-V.3 Internal Tension and the Shape of
-While the external tension with the Distance Ladder dominates the landscape, we have identified a secondary internal tension among the One-Step measurements that may hold the key to understanding the underlying physics.
-Measurements that assume the CDM shape of (Category 2) yield km s-1 Mpc-1.
-However, measurements that are free from this shape assumption (Category 3) yield a higher intermediate value of km s-1 Mpc-1.
-This km s-1 Mpc-1 shift accounts for more than half of the total Hubble tension.
-This systematic offset has two possible interpretations:
-1.
-CDM is correct: Model-independent methods (Category 3) suffer from larger uncontrolled systematics that bias high.
-2.
-CDM is incomplete: The true expansion history deviates from the CDM prediction at , and forcing a CDM fit biases low.
-The fact that specific model-independent measurements—particularly the Vogl et al.
-[Vogl2024] tailored EPM ()— yield values consistent with the Distance Ladder is intriguing.
-This method is independent of both the Cepheid/SNe Ia calibration chain and cosmological model assumptions, yet it supports the high- solution.
-This pattern is more naturally explained by interpretation (2) than by a conspiracy of unrelated systematics.
-The persistence of this internal tension reinforces earlier findings that the CDM model may face challenges beyond the Hubble tension alone.
-A comprehensive review of these challenges [Perivolaropoulos2022review] identifies correlated anomalies that may point toward a common origin in modified late-time physics.
-V.4 Implications for Proposed Solutions
+V.2 Robustness of Results
+To assess the sensitivity of our results to methodological choices in measurement categorization, we performed a comprehensive robustness analysis111We thank Dan Scolnic for suggesting the examination of the categorization sensitivity for these specific cases to test the robustness of the internal tension..
+We considered a more conservative approach to categorization by: (i) adopting alternative reported values for measurements with multiple H0 estimates (Liu et al.: 59.1 67.5 km s-1 Mpc-1; Dominguez et al.: 61.9 65.6 km s-1 Mpc-1), (ii) removing potentially redundant measurements from the same observational data (Kelly/Li Refsdal analyses), and (iii) excluding measurements whose categorization could be debated (Zhang et al.
+2022 cosmic chronometer analysis).
+Table 8 summarizes the impact of these changes on our tension estimates.
+| Scenario | Cat2 | Cat3 | Tension | Tension |
+| (DL-LI) | (C2-C3) | |||
+| Baseline (this work) | ||||
+| Liu: 59.167.5 | ||||
+| Dominguez: 61.965.6 | ||||
+| Remove Kelly (Refsdal) | ||||
+| Remove Li (Refsdal) | ||||
+| Remove Zhang 2022 | ||||
+| Liu + Dominguez | ||||
+| All changes (keep Li) | ||||
+| All changes (keep Kelly) |
+The analysis reveals that our main result—the significant tension between Distance Ladder and all others measurements—is robust to all tested variations in categorization methodology.
+Even under the most conservative scenario combining all alternative choices, the main tension remains at , well above the threshold for statistical significance.
+No individual change affects the main tension by more than .
+In contrast, the internal tension between Category 2 (CDM-dependent) and Category 3 (cosmological-model-independent) measurements is more sensitive to these methodological choices, varying from to depending on the specific categorization adopted.
+V.3 Implications for Proposed Solutions
 Our findings have significant implications for the major classes of proposed solutions to the Hubble tension:
-V.4.1 Early Dark Energy and Sound Horizon Modifications
-Models that resolve the tension by reducing the Sound Horizon at recombination (e.g., Early Dark Energy [Knox2020, Mirpoorian2025], extra relativistic species, early modified gravity) predict that Sound-Horizon-independent measurements should align with the Distance Ladder.
-Our analysis shows the opposite: Sound-Horizon-free measurements (Categories 2, 3, and 4) predominantly align with Planck, not with SH0ES.
+V.3.1 Early Dark Energy and Sound Horizon Modifications
+Models that resolve the tension by reducing the Sound Horizon at recombination (e.g., Early Dark Energy [91, 114, 173], extra relativistic species, early modified gravity) predict that Sound-Horizon-Free measurements should align with the Distance Ladder.
+Our analysis shows the opposite: Distance Ladder Independent/Sound Horizon Free (DLI–SHF) (Categories 2, 3, and 4) yields a value closer to Planck, not to SH0ES.
 This presents a significant challenge to early-universe solutions. If EDE or similar models were correct, we would expect:
@@
 Sound-Horizon-Free BAO to yield km s-1 Mpc-1 ✗ (observed: )
-The failure of most Sound-Horizon-independent probes to support the high- solution disfavors pure early-universe modifications as the complete explanation.
-V.4.2 Late-Time Modifications
+The failure of most Sound-Horizon-Free probes to support the high- solution disfavors pure early-universe modifications as the complete explanation.
+V.3.2 Late-Time Modifications
 Models that modify the expansion history at late times (e.g., dynamical dark energy, late-time gravitational transitions, interacting dark energy) predict that the inferred should depend on whether CDM is assumed in the analysis.
@@
 They must be consistent with other late-time probes (SNe Ia Hubble diagrams, BAO, growth of structure) that show no dramatic deviation from CDM.
-Recent DESI results [DESI2024, Giare:2025pzu, Gao:2024ily] suggesting evolving dark energy (, ) may provide partial support, but the statistical significance remains modest and the robustness across different parameterizations has been questioned [Colgain:2024mtg].
+Recent DESI results [43, 72, 68] suggesting evolving dark energy (, ) provide evidence for deviations from CDM at late times.
+However, due to the strong anti-correlation between and [40], such models with actually predict lower values, potentially exacerbating rather than alleviating the tension with the Distance Ladder.
+Moreover, the statistical significance of the DESI signal remains modest (–) and its robustness across different parameterizations has been questioned [39].
 Theoretical frameworks for such behavior include models where the effective gravitational constant undergoes a rapid transition at late times, which can arise naturally in scalar-tensor theories.
@@
 A particularly intriguing perspective recasts the Hubble tension as an crisis.
-Theodoropoulos & Perivolaropoulos demonstrated that if the intrinsic luminosity of SNe Ia in the Hubble flow differs systematically from that in calibrator galaxies, the inferred would be biased without requiring modifications to fundamental cosmology [Marra2021, Perivolaropoulos:2025gzo, Theodoropoulos2021] .
-V.4.3 Distance Ladder Systematics
-The isolation of the Distance Ladder as a outlier against 54 independent measurements points toward unrecognized systematics within the calibration chain. Potential systematic effects include:
+Theodoropoulos & Perivolaropoulos demonstrated that if the intrinsic luminosity of SNe Ia in the Hubble flow differs systematically from that in calibrator galaxies, the inferred would be biased without requiring modifications to fundamental cosmology [112, 128, 177].
+V.3.3 Distance Ladder Systematics
+The isolation of the Distance Ladder as a outlier against 58 independent measurements points toward unrecognized systematics within the calibration chain. Potential systematic effects include:
 Cepheid systematics: Metallicity dependence of the period-luminosity relation, crowding and blending effects, or extinction corrections.
-SNe Ia systematics: Environmental dependencies (host galaxy mass, local star formation), evolution with redshift, or standardization procedure biases.
 Calibrator physics: Differences between calibrators observed in anchor galaxies vs. SNe Ia hosts.
@@
 However, the “systematics” interpretation faces its own challenges:
-Multiple independent calibrators (Cepheids, TRGB, JAGB, Miras, SBF) yield consistent high- values.
-JWST observations have largely confirmed HST Cepheid photometry [Riess2025].
-Model-independent methods like the tailored EPM [Vogl2024] support the high- solution.
-V.4.4 Hybrid Scenarios
+V.3.4 Hybrid Scenarios
 Given the complexity of the evidence, a hybrid scenario involving both partial distance ladder systematics and deviations from CDM dynamics may be required. For example:
@@
 Such a scenario is constrained but not excluded by current data.
-V.5 Falsifiability and Future Tests
+V.4 Falsifiability and Future Tests
 A robust scientific framework requires clear predictions that can be tested by future observations. We outline the key tests that would distinguish between the competing interpretations:
-V.5.1 What Would Confirm Distance Ladder Systematics?
+V.4.1 What Would Confirm Distance Ladder Systematics?
 The “systematics” hypothesis would be strongly supported if:
@@
 Independent geometric distance methods (e.g., detached eclipsing binaries in external galaxies) yield distances inconsistent with Cepheids.
-V.5.2 What Would Confirm New Physics?
+V.4.2 What Would Confirm New Physics?
 The “new physics” hypothesis would be strongly supported if:
@@
 2.
-Multiple independent model-independent methods (EPM, Megamasers, future techniques) consistently yield km s-1 Mpc-1.
+Multiple independent cosmological-model-independent methods (EPM, Megamasers, future techniques) consistently yield km s-1 Mpc-1.
 3.
@@
 A theoretical model is developed that simultaneously explains the Hubble tension, the Category 2 vs. Category 3 offset, and passes all other cosmological tests.
-V.6 Observational Priorities
+V.5 Observational Priorities
 Based on our analysis, we identify the following observational priorities for resolving the Hubble tension:
 1.
-Expand Category 3 (Pure Local One-Step): This category currently contains only 15 measurements with significant internal heterogeneity.
-Additional model-independent measurements are crucial for determining whether the intermediate km s-1 Mpc-1 persists or converges toward one of the two solutions.
+Expand Category 3 (Pure Local): This category currently contains only 16 measurements with significant internal heterogeneity.
+Additional cosmological-model-independent measurements are crucial for determining whether the intermediate km s-1 Mpc-1 persists or converges toward one of the two solutions.
 2.
 Improve tailored EPM precision: The Vogl et al.
-[Vogl2024] result is currently the strongest model-independent support for the high- solution outside the traditional Distance Ladder.
+[182] result is currently the strongest cosmological-model-independent support for the high- solution outside the traditional Distance Ladder.
 Larger SNe II samples observed with JWST spectroscopy could significantly reduce uncertainties.
@@
 4.
-Accumulate GW standard sirens: The LIGO-Virgo-KAGRA O4 and O5 [LIGOScientific:2025jau] runs will significantly increase the sample of both bright and dark sirens.
+Accumulate GW standard sirens: The LIGO-Virgo-KAGRA O4 and O5 [1] runs will significantly increase the sample of both bright and dark sirens.
 Crucially, these should be analyzed both with and without CDM assumptions to test for model-dependent biases.
 5.
-Model-independent Strong Lensing: Future time-delay measurements should emphasize cosmographic or model-independent analyses alongside standard CDM fits.
+Cosmological-model-independent Strong Lensing: Future time-delay measurements should emphasize cosmographic or cosmological-model-independent analyses alongside standard CDM fits.
 6.
 Cross-calibration tests: Direct comparisons between different calibrators (Cepheids vs. TRGB vs. JAGB) in the same galaxies can identify potential systematics.
-V.7 Concluding Remarks
+V.6 Concluding Remarks
 The Hubble tension remains one of the most significant challenges to the standard cosmological model.
-Our analysis of 83 Sound-Horizon-independent measurements reveals that the tension is more accurately characterized as a “Distance Ladder vs.
+Our analysis of 88 Sound-Horizon-Free measurements reveals that the tension is more accurately characterized as a “Distance Ladder vs.
 The Rest” crisis rather than an “Early vs.
@@
 1.
-The Distance Ladder yields km s-1 Mpc-1, while all other Sound-Horizon-free methods combined yield km s-1 Mpc-1, a ( correlation-adjusted) tension.
+The Distance Ladder yields km s-1 Mpc-1, while all other Distance Ladder Independent/Sound Horizon Free(DLI–SHF) methods combined yield km s-1 Mpc-1, a ( correlation-adjusted) tension.
 2.
@@
 3.
-A internal tension exists between CDM-dependent (Category 2) and model-independent (Category 3) local measurements, with the CDM assumption reducing inferred by km s-1 Mpc-1.
+A internal tension exists between CDM-dependent (Category 2) and cosmological-model-independent (Category 3) local measurements, with the CDM assumption reducing inferred by km s-1 Mpc-1.
 4.
-Specific model-independent measurements (tailored EPM,) support the high- solution, complicating the pure “systematics” interpretation.
+Specific cosmological-model-independent measurements (e.g., tailored EPM) support the high- solution, complicating the pure “systematics” interpretation.
 5.
@@
 The four-category classification scheme introduced in this work provides a framework for tracking progress toward this resolution as new measurements become available.
-Whatever the ultimate explanation, the Hubble tension has already achieved a valuable scientific outcome: it has motivated an unprecedented scrutiny of both our measurement techniques and our theoretical assumptions, pushing cosmology toward greater rigor and potentially toward new discoveries.Recent results from the Atacama Cosmology Telescope DR6 provide additional CMB constraints that will be crucial for further testing the robustness of our conclusions [AtacamaCosmologyTelescope:2025blo, AtacamaCosmologyTelescope:2025nti]
+Whatever the ultimate explanation, the Hubble tension has already achieved a valuable scientific outcome: it has motivated an unprecedented scrutiny of both our measurement techniques and our theoretical assumptions, pushing cosmology toward greater rigor and potentially toward new discoveries.Recent results from the Atacama Cosmology Telescope DR6 provide additional CMB constraints that will be crucial for further testing the robustness of our conclusions [108, 28]
 Acknowledgements.
-This research was supported by COST Action CA21136 - Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse), supported by COST (European Cooperation in Science and Technology).Appendix A Complete Data Tables
+This research was supported by COST Action CA21136 – Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse), supported by COST (European Cooperation in Science and Technology).
+We thank Adam Riess and Dan Scolnic for their detailed and constructive comments on measurement categorization, which helped us strengthen and clarify our analysis.
+We are grateful to Levon Pogosian for providing the unpublished DESI DR2 Sound-Horizon-Free result and to Helena García Escudero and Seyed Hamidreza Mirpoorian for sharing their work prior to publication.
+We thank Pilar Ruiz-Lapuente for useful correspondence regarding the JAGB calibration debate and for providing her “SNe Ia twins” measurement prior to its inclusion in our compilation.
+We also thank Yi-Ying Wang for pointing out relevant FRB and gravitational wave measurements, and Yi-min Zhang for sharing results on interacting dark energy models.Appendix A Complete Data Tables
 In this appendix, we provide expanded versions of the data tables presented in the main text, including additional metadata and methodological details for each measurement.
@@
 | Category | Range | ||||
-| 1. Distance Ladder | 29 | 72.74 | 0.40 | 0.74 | 67.8–76.9 |
-| 2. Local One-Step (CDM) | 33 | 67.18 | 0.90 | 0.85 | 59.1–86.2 |
-| 3. Pure Local One-Step | 15 | 70.38 | 1.00 | 0.78 | 65.9–75.5 |
-| 4. CMB Sound Horizon Free | 6 | 68.71 | 0.63 | 0.36 | 61–69.5 |
-| Combined One-Step (2+3+4) | 54 | 68.67 | 0.46 | 0.85 | 59.1–86.2 |
+| 1. Distance Ladder | 30 | 72.73 | 0.39 | 0.72 | 67.8–77.0 |
+| 2. Local (CDM) | 33 | 67.61 | 0.96 | 0.80 | 54.6–86.2 |
+| 3. Pure Local | 16 | 71.03 | 0.69 | 0.77 | 65.9–75.5 |
+| 4. CMB–SHF | 9 | 69.07 | 0.44 | 0.89 | 61–70.8 |
+| Combined DLI–SHF (2+3+4) | 58 | 69.37 | 0.34 | 0.95 | 54.6–86.2 |
 Notes: is the inverse-variance weighted mean in km s-1 Mpc-1. Range shows min–max of central values.
 A.1 Distance Ladder Measurements: Extended Information
-Table 9 provides extended information for all 29 distance-ladder-dependent measurements.
+Table 10 provides extended information for all 30 Distance-Ladder-dependent measurements.
 We include the geometric anchor(s) employed, the primary distance indicator, the secondary distance indicator (if applicable), and the reported value with uncertainty.
 | # | Author (Year) | Anchor | Primary | Secondary | range | |
-| 1 | Huang (2019) [Huang2020] | NGC 4258, MW, LMC | Mira | SNe Ia | ||
-| 2 | de Jaeger (2020) [deJaeger2020] | LMC, MW | Cepheids/TRGB | SNe II | ||
-| 3 | Kourkchi (2020) [Kourkchi2020] | Cepheids, TRGB | Tully–Fisher | — | ||
-| 4 | Khetan (2021) [Khetan2021] | Cepheids, TRGB | SBF | SNe Ia | ||
-| 5 | Blakeslee (2021) [Blakeslee2021] | NGC 4258, LMC | SBF | SNe Ia | ||
-| 6 | Freedman (2021) [Freedman2021] | LMC (DEB), MW | TRGB | SNe Ia | ||
-| 7 | Dhawan (2022) [Dhawan2022] | LMC | TRGB | SNe Ia | ||
-| 8 | Kenworthy (2022) [Kenworthy2022] | NGC 4258, LMC | Cepheids | — | ||
-| 9 | Dhawan (2022) [Dhawan2023] | LMC | TRGB | SNe Ia | ||
-| 10 | Dhawan (2022) [Dhawan2023] | NGC 4258, LMC, MW | Cepheids | SNe Ia | ||
-| 11 | Scolnic (2023) [Scolnic2024] | LMC | TRGB | SNe Ia | ||
-| 12 | de Jaeger (2023) [dejaeger2023] | LMC, MW | Cepheids | SNe II | ||
-| 13 | Uddin (2023) [uddin2023] | Multiple | Combined | SNe Ia | ||
-| 14 | Uddin (2023) [uddin2023] | Multiple | Combined | SNe Ia | ||
-| 15 | Huang (2023) [Huang2024] | NGC 4258, LMC | Mira | SNe Ia | ||
-| 16 | Li (2024) [Li2024b] | LMC, SMC | JAGB | SNe Ia | ||
-| 17 | Chavez (2024) [chavez2025] | Cepheids | H II – | — | ||
-| 18 | Boubel (2024) [Boubel2024] | Cepheids, TRGB | Tully–Fisher | — | ||
-| 19 | Lee (2024) [Lee2024] | LMC | JAGB | SNe Ia | ||
-| 20 | Freedman (2024) [Freedman2024] | LMC (DEB), MW | Cepheids | SNe Ia | ||
-| 21 | Said (2024) [Said2025] | SBF | Fund. Plane | — | ||
-| 22 | Zhang (2024)a [Zhang2024] | — | Pairwise Vel. | — | ||
-| 23 | Jensen (2025) [Jensen2025] | LMC, MW | TRGB | — | ||
-| 24 | Wojtak (2025) [Wojtak2025] | NGC 4258, LMC, MW | Cepheids | SNe Ia | ||
-| 25 | Bhardwaj (2025) [Bhardwaj2025] | NGC 4258, LMC | Mira | SNe Ia | ||
-| 26 | Newman (2025) [Newman2025] | LMC | TRGB | SNe Ia | ||
-| 27 | Riess (2025)b [Riess2025] | NGC 4258, LMC, MW | Cepheids | SNe Ia | ||
-| 28 | Kudritzki (2025) [kudritzki2025] | FGLR | Tully–Fisher | — | ||
-| 29 | Wagner (2025)c [wagner2025] | — | TRGB | Infall |
+| 1 | Huang (2019) [78] | NGC 4258, MW, LMC | Mira | SNe Ia | ||
+| 2 | de Jaeger (2020) [42] | LMC, MW | Cepheids/TRGB | SNe II | ||
+| 3 | Kourkchi (2020) [92] | Cepheids, TRGB | Tully–Fisher | — | ||
+| 4 | Khetan (2021) [90] | Cepheids, TRGB | SBF | SNe Ia | ||
+| 5 | Blakeslee (2021) [22] | NGC 4258, LMC | SBF | SNe Ia | ||
+| 6 | Freedman (2021) [63] | LMC (DEB), MW | TRGB | SNe Ia | ||
+| 7 | Dhawan (2022) [47] | LMC | TRGB | SNe Ia | ||
+| 8 | Kenworthy (2022) [88] | NGC 4258, LMC | Cepheids | — | ||
+| 9 | Dhawan (2022) [48] | LMC | TRGB | SNe Ia | ||
+| 10 | Dhawan (2022) [48] | NGC 4258, LMC, MW | Cepheids | SNe Ia | ||
+| 11 | Scolnic (2023) [161] | LMC | TRGB | SNe Ia | ||
+| 12 | de Jaeger (2023) [41] | LMC, MW | Cepheids | SNe II | ||
+| 13 | Uddin (2023) [180] | Multiple | Combined | SNe Ia | ||
+| 14 | Uddin (2023) [180] | Multiple | Combined | SNe Ia | ||
+| 15 | Huang (2023) [79] | NGC 4258, LMC | Mira | SNe Ia | ||
+| 16 | Li (2024) [98] | LMC, SMC | JAGB | SNe Ia | ||
+| 17 | Chavez (2024) [34] | Cepheids | H II – | — | ||
+| 18 | Boubel (2024) [24] | Cepheids, TRGB | Tully–Fisher | — | ||
+| 19 | Lee (2024) [96] | LMC | JAGB | SNe Ia | ||
+| 20 | Freedman (2024) [62] | LMC (DEB), MW | Cepheids | SNe Ia | ||
+| 21 | Said (2024) [156] | SBF | Fund. Plane | — | ||
+| 22 | Zhang (2024)a [193] | — | Pairwise Vel. | — | ||
+| 23 | Jensen (2025) [83] | LMC, MW | TRGB | — | ||
+| 24 | Wojtak (2025) [186] | NGC 4258, LMC, MW | Cepheids | SNe Ia | ||
+| 25 | Bhardwaj (2025) [20] | NGC 4258, LMC | Mira | SNe Ia | ||
+| 26 | Newman (2025) [117] | LMC | TRGB | SNe Ia | ||
+| 27 | Riess (2025)b [149] | NGC 4258, LMC, MW | Cepheids | SNe Ia | ||
+| 28 | Kudritzki (2025) [94] | FGLR | Tully–Fisher | — | ||
+| 29 | Wagner (2025)c [183] | — | TRGB | Infall | ||
+| 30 | Lapuente (2025) [155] | — | M 101, NGC 5643, NGC 7250, M 66 | SNe Ia (twins) |
 Notes: DEB = Detached Eclipsing Binaries; FGLR = Flux-weighted Gravity–Luminosity Relation; SBF = Surface Brightness Fluctuations.
@@
 b
-We adopt this measurement for all numerical values in this work. For comparisons, we use the SH0ES 2022 result [Riess2022] as reference baseline, given its established presence in the literature.
+We adopt this measurement for all numerical values in this work. For comparisons, we use the SH0ES 2022 result [150] as reference baseline, given its established presence in the literature.
 c
@@
 Other: 2 measurements (pairwise velocities, HII regions)
-A.2 One-Step Measurements: Extended Information
-Table 10 provides extended information for all 54 One-Step measurements from Categories 2, 3, and 4.
+A.2 Distance Ladder Independent/Sound Horizon Free (DLI–SHF) Measurements: Extended Information
+Table 11 provides extended information for all 58 DLI–SHF measurements from Categories 2, 3, and 4.
 We include the physical observable, the cosmological probe, the model assumptions (which determine the category assignment), and the reported value with uncertainty.
@@
 CDM: Category 2 — Analysis explicitly assumes flat CDM expansion history
-Independent: Category 3 — Model-independent analysis (geometric, cosmographic, or direct )
+Independent: Category 3 — Cosmological-model-independent analysis (geometric, cosmographic, or direct )
 CMB–NSH: Category 4 — Uses CMB data but avoids Sound Horizon (No Sound Horizon)
 The measurements span diverse physical probes:
-Strong Gravitational Lensing: 15 measurements (time delays from quasars, SNe, and clusters)
-Gravitational Waves: 8 measurements (bright and dark sirens)
-Fast Radio Bursts: 8 measurements (dispersion measure–redshift relation)
+Strong Gravitational Lensing: 18 measurements (time delays from quasars, SNe, and clusters)
+Gravitational Waves: 9 measurements (bright and dark sirens)
+Fast Radio Bursts: 7 measurements (dispersion measure–redshift relation)
 Cosmic Chronometers: 4 measurements (differential galaxy ages)
-Megamasers: 2 measurements (geometric disk distances)
-CMB-based (no ): 6 measurements (SZ effect, scale)
-Other: 11 measurements (EPM, kilonovae, gamma-ray attenuation, etc.)
+Megamasers: 3 measurements (geometric disk distances)
+CMB-based (no ): 9 measurements (SZ effect, scale, etc.)
+Other: 8 measurements (EPM, kilonovae, gamma-ray attenuation, etc.)
 | # | Cat. | Author (Year) | Observable | Probe | Model | z range | |
-| Category 2: Local One-Step (CDM Assumption) — 33 measurements | |||||||
-| 1 | 2 | Wu (2021) [Wu2022] | Dispersion measure, | FRBs | CDM | 0.0039–0.695 | |
-| 2 | 2 | Abbott (2021) [Abbott2021] | GW strain amplitude | CBC dark sirens | CDM | ||
-| 3 | 2 | Philcox (2022) [Philcox2022] | BAO scale, shape | Matter–rad. eq. | CDM | 0.38–0.61 | |
-| 4 | 2 | Shajib (2023) [Shajib2023] | Time delays, dynamics | Lensed quasars | CDM | 0.657 | |
-| 5 | 2 | Napier (2023) [Napier2023] | Time delays | Cluster lenses | CDM | 1.734–2.805 | |
-| 6 | 2 | Kelly (2023) [Kelly2023] | Time delays | SN Refsdal | CDM | 1.49 | |
-| 7 | 2 | Dominguez (2023) [Dominguez2023] | -ray opacity | Blazar spectra | CDM | ||
-| 8 | 2 | Gao (2023) [gao2023] | DM– + SNe Ia | FRBs | CDM | 0.039–2.3 | |
-| 9 | 2 | Liu (2023) [liu2023h] | Time delays | Cluster lensed QSO | CDM | ||
-| 10 | 2 | Fung (2023) [Fung2023] | GW strain only | NSBH population | CDM | ||
-| 11 | 2 | Martinez (2023) [Martinez2023] | Time delays | Lensed quasars | CDM | 1.734 | |
-| 12 | 2 | Ballard (2023) [Ballard2023] | GW strain + galaxy | Dark siren | CDM | ||
-| 13 | 2 | Moresco (2023) [Moresco2023] | Spectroscopic ages | Passive galaxies | CDM | 0.07–1.965 | |
-| 14 | 2 | Alfradique (2024) [Alfradique2023] | GW strain + galaxy | Dark sirens | CDM | ||
-| 15 | 2 | Grillo (2024) [Grillo2024] | Time delays | HFF clusters | CDM | 1.240–3.703 | |
-| 16 | 2 | Pascale (2024) [Pascale2024] | Time delays | H0pe SN (JWST) | CDM | 1.78 | |
-| 17 | 2 | Hernandez (2024) [Hernandez2024] | GW + BGP | Spectral sirens | CDM | ||
-| 18 | 2 | Bom (2024) [Bom2024] | GW + kilonova | O4a events | CDM | 0.1–0.6 | |
-| 19 | 2 | TDCOSMO XVI (2024) [TDCOSMO2024] | Time delays | WGD2038 | CDM | 0.777 | |
-| 20 | 2 | Gao (2024) [Gao2025a] | DM– relation | Localized FRBs | CDM | 0.0008–1.016 | |
-| 21 | 2 | Li (2024) [Li2024c] | Time delays (corrected) | Lensed quasars | CDM | ||
-| 22 | 2 | Yang (2024) [Yang2024] | DM + scattering | FRBs | CDM | 0.034–1.016 | |
-| 23 | 2 | Piratova-Moreno (2025) [PiratovMoreno2025] | DM– relation | FRBs (MLE) | CDM | 0.0008–1.354 | |
-| 24 | 2 | Barua (2025) [Barua2025] | Maser dynamics | MCP sample | CDM | 0.002–0.034 | |
-| 25 | 2 | Zhang (2025) [Zhang2025] | PRS association | FRBs | CDM | 0.03–0.25 | |
-| 26 | 2 | Gao (2025) [Gao2025] | RM–PRS correlation | FRBs | CDM | 0.098–0.241 | |
-| 27 | 2 | Loubser (2025) [Loubser2025] | Spectroscopic ages | BCGs | CDM | 0.3–0.7 | |
-| 28 | 2 | Beirnaert (2025) [Beirnaert2025] | GW strain | Dark sirens (avg) | CDM | ||
-| 29 | 2 | Bahr-Kalus (2025) [Bahr-Kalus2025] | Turnover scale + BAO | DESI | CDM | 0.4–3.1 | |
-| 30 | 2 | Birrer (2025) [Birrer2025] | Time delays + kinematics | Multiple lenses | CDM | 0.082–0.745 | |
-| 31 | 2 | Liu (2025) [liu2025a] | Strong grav. lensing | Multiple lenses | CDM | 1.49 | |
-| 32 | 2 | Pierel (2025) [Pierel2025] | Time delays | SN Encore | CDM | 1.95 | |
-| 33 | 2 | Paic (2025) [paic2025tdcosmoxxivmeasurementhubble] | Strong Lensing | SMARTS | CDM | 2.32 | |
-| Category 3: Pure Local One-Step (Model Independent) — 15 measurements | |||||||
-| 34 | 3 | Kuo (2013) [Kuo2013] | Maser positions, velocities | Megamaser disk | Independent | 0.034 | |
-| 35 | 3 | Gao (2015) [Gao2015] | Maser positions, velocities | Megamaser disk | Independent | 0.028 | |
-| 36 | 3 | Bulla (2022) [bulla2022] | GW + kilonova light curve | BNS merger | Independent | 0.01 | |
-| 37 | 3 | Zhang (2022) [Zhang2022] | Spectroscopic ages | Passive galaxies | Independent | 0.009–2.545 | |
-| 38 | 3 | Du (2023) [Du2023] | Time delays + GRBs | Lensed quasars | Independent | 0.654–2.375 | |
-| 39 | 3 | Palmese (2023) [Palmese2023] | GW + afterglow | GW170817 | Independent | 0.01 | |
-| 40 | 3 | Sneppen (2023) [Sneppen2023] | Photospheric velocity | Kilonova | Independent | 0.0986 | |
-| 41 | 3 | Liu (2023) [liu2023a] | Lensing + SNe Ia | Combined | Independent | 0.01–2.3 | |
-| 42 | 3 | Gonzalez (2024) [Gonzalez2024] | Gas fraction + SNe Ia | Clusters | Independent | 0.01–2.3 | |
-| 43 | 3 | Li (2024) [Li2024a] | Time delays | SN Refsdal | Independent | 0.001–2.261 | |
-| 44 | 3 | Jaiswal (2024) [Jaiswal2024] | RM lag | NGC 5548 | Independent | 0.017 | |
-| 45 | 3 | Vogl (2024) [Vogl2024] | Photospheric expansion | SNe II | Independent | 0.01–0.04 | |
-| 46 | 3 | Song (2025) [Song2025] | GW + lensing | Combined | Independent | ||
-| 47 | 3 | Colaco (2025) [Colaco2025] | Lensing + SNe Ia | Joint | Independent | 0.001–2.261 | |
-| 48 | 3 | Favale (2025) [Favale2025] | Ages + SNe Ia | CC + Pantheon+ | Independent | 0.07–1.965 | |
-| Category 4: CMB Sound Horizon Free — 6 measurements | |||||||
-| 49 | 4 | Reese (2003) [Reese2003] | SZ decrement, X-ray | Galaxy clusters | CMB–NSH | 0.02–0.83 | |
-| 50 | 4 | Colaco (2023) [Colaco2023] | SZ + X-ray + SNe Ia | Clusters | CMB–NSH | 0.01–2.3 | |
-| 51 | 4 | Pogosian (2024) [Pogosian2024] | BAO peak + BBN | DESI BAO | CMB–NSH | 0.1–4.16 | |
-| 52 | 4 | Pogosian (2024) [Pogosian2024] | BAO peak + BBN | Pre-DESI BAO | CMB–NSH | 0.1–4.16 | |
-| 53 | 4 | Bahr-Kalus (2025) [Bahr-Kalus2025] | Turnover scale + SNe | DESI + Pantheon+ | CMB–NSH | 0.4–3.1 | |
-| 54 | 4 | Krolewski (2025) [Krolewski2025] | DESI DR1 + CMB | No sound horizon | CMB–NSH | 0.1–2.33 | |
+| Category 2: Local (CDM Assumption) — 33 measurements | |||||||
+| 1 | 2 | Wu (2021) [189] | Dispersion measure, | FRBs | CDM | 0.0039–0.695 | |
+| 2 | 2 | Abbott (2021) [102] | GW strain amplitude | CBC dark sirens | CDM | ||
+| 3 | 2 | Shajib (2023) [163] | Time delays, dynamics | Lensed quasars | CDM | 0.657 | |
+| 4 | 2 | Napier (2023) [116] | Time delays | Cluster lenses | CDM | 1.734–2.805 | |
+| 5 | 2 | Kelly (2023) [87] | Time delays | SN Refsdal | CDM | 1.49 | |
+| 6 | 2 | Dominguez (2023) [53] | -ray opacity | Blazar spectra | CDM† | ||
+| 7 | 2 | Gao (2023) [67] | DM– + SNe Ia | FRBs | CDM | 0.039–2.3 | |
+| 8 | 2 | Liu (2023) [104] | Time delays | Cluster lensed QSO | CDM | ||
+| 9 | 2 | Fung (2023) [64] | GW strain only | NSBH population | CDM | ||
+| 10 | 2 | Martinez (2023) [113] | Time delays | Lensed quasars | CDM | 1.734 | |
+| 11 | 2 | Ballard (2023) [14] | GW strain + galaxy | Dark siren | CDM | ||
+| 12 | 2 | Moresco (2023) [115] | Spectroscopic ages | Passive galaxies | CDM | 0.07–1.965 | |
+| 13 | 2 | Alfradique (2024) [11] | GW strain + galaxy | Dark sirens | CDM | ||
+| 14 | 2 | Grillo (2024) [75] | Time delays | HFF clusters | CDM | 1.240–3.703 | |
+| 15 | 2 | Pascale (2024) [125] | Time delays | H0pe SN (JWST) | CDM | 1.78 | |
+| 16 | 2 | Hernandez (2024) [110] | GW + BGP | Spectral sirens | CDM | ||
+| 17 | 2 | Bom (2024) [23] | GW + kilonova | O4a events | CDM | 0.1–0.6 | |
+| 18 | 2 | TDCOSMO XVI (2024) [188] | Time delays | WGD2038 | CDM | 0.777 | |
+| 19 | 2 | Gao (2024) [65] | DM– relation | Localized FRBs | CDM | 0.0008–1.016 | |
+| 20 | 2 | Li (2024) [99] | Time delays (corrected) | Lensed quasars | CDM | ||
+| 21 | 2 | Yang (2024) [190] | DM + scattering | FRBs | CDM | 0.034–1.016 | |
+| 22 | 2 | Piratova-Moreno (2025) [140] | DM– relation | FRBs (MLE) | CDM | 0.0008–1.354 | |
+| 23 | 2 | Barua (2025) [17] | Maser dynamics | MCP sample | CDM | 0.002–0.034 | |
+| 24 | 2 | Zhang (2025) [195] | PRS association | FRBs | CDM | 0.03–0.25 | |
+| 25 | 2 | Gao (2025) [69] | RM–PRS correlation | FRBs | CDM | 0.098–0.241 | |
+| 26 | 2 | Loubser (2025) [107] | Spectroscopic ages | BCGs | CDM | 0.3–0.7 | |
+| 27 | 2 | Beirnaert (2025) [18] | GW strain | Dark sirens (avg) | CDM | ||
+| 28 | 2 | Bahr-Kalus (2025) [13] | Turnover scale + BAO | DESI | CDM | 0.4–3.1 | |
+| 29 | 2 | Birrer (2025) [175] | Time delays + kinematics | Multiple lenses | CDM | 0.082–0.745 | |
+| 30 | 2 | Liu (2025) [105] | Strong grav. lensing | Multiple lenses | CDM | 1.49 | |
+| 31 | 2 | Pierel (2025) [138] | Time delays | SN Encore | CDM | 1.95 | |
+| 32 | 2 | Paic (2025) [121] | Strong Lensing | SMARTS | CDM | 2.32 | |
+| 33 | 2 | Oliveira (2026) [12] | 3×2pt + Standard Sirens | Shear/positions + GW strain | CDM | 0.1-2 | |
+| Category 3: Pure Local (Cosmological-Model Independent) — 16 measurements | |||||||
+| 34 | 3 | Kuo (2013) [95] | Maser positions, velocities | Megamaser disk | Independent | 0.034 | |
+| 35 | 3 | Gao (2015) [66] | Maser positions, velocities | Megamaser disk | Independent | 0.028 | |
+| 36 | 3 | Bulla (2022) [26] | GW + kilonova light curve | BNS merger | Independent | 0.01 | |
+| 37 | 3 | Zhang (2022) [192] | Spectroscopic ages | Passive galaxies | Independent | 0.009–2.545 | |
+| 38 | 3 | Du (2023) [55] | Time delays + GRBs | Lensed quasars | Independent | 0.654–2.375 | |
+| 39 | 3 | Palmese (2023) [122] | GW + afterglow | GW170817 | Independent | 0.01 | |
+| 40 | 3 | Sneppen (2023) [171] | Photospheric velocity | Kilonova | Independent | 0.0986 | |
+| 41 | 3 | Liu (2023) [103] | Lensing + SNe Ia | Combined | Independent | 0.01–2.3 | |
+| 42 | 3 | Gonzalez (2024) [74] | Gas fraction + SNe Ia | Clusters | Independent | 0.01–2.3 | |
+| 43 | 3 | Li (2024) [101] | Time delays | SN Refsdal | Independent | 0.001–2.261 | |
+| 44 | 3 | Jaiswal (2024) [81] | RM lag | NGC 5548 | Independent | 0.017 | |
+| 45 | 3 | Vogl (2024) [182] | Photospheric expansion | SNe II | Independent | 0.01–0.04 | |
+| 46 | 3 | Song (2025) [172] | GW + lensing | Combined | Independent | ||
+| 47 | 3 | Colaco (2025) [37] | Lensing + SNe Ia | Joint | Independent | 0.001–2.261 | |
+| 48 | 3 | Du (2025) [54] | CC+BAO+ SNe Ia | IDL+ MAPAge | Independent | 0.001–2.5 | |
+| 49 | 3 | Favale (2025) [60] | Ages + SNe Ia | CC + Pantheon+ | Independent | 0.07–1.965 | |
+| Category 4: CMB Sound Horizon Free (CMB–SHF) — 9 measurements | |||||||
+| 50 | 4 | Reese (2003) [147] | SZ decrement, X-ray | Galaxy clusters | CMB–NSH | 0.02–0.83 | |
+| 51 | 4 | Philcox (2022) [137] | BAO scale, shape | Matter–rad. eq. | CMB–NSH | 0.2–0.75 | |
+| 52 | 4 | Colaco (2023) [36] | SZ + X-ray + SNe Ia | Clusters | CMB–NSH | 0.01–2.3 | |
+| 53 | 4 | Pogosian (2024) [144] | BAO peak + BBN | Pre-DESI BAO | CMB–NSH | 0.1–4.16 | |
+| 54 | 4 | Pogosian (2025) [145] | BAO peak + BBN | DESI DR BAO | CMB–NSH | 0.1–4.16 | |
+| 55 | 4 | Bahr-Kalus (2025) [13] | Turnover scale + SNe | DESI + Pantheon+ | CMB–NSH | 0.4–3.1 | |
+| 56 | 4 | Garcia (2025) [70] | BAO+ +CMB lensing | DESI+APS+ DES 3×2pt | CMB–NSH | 0.1-4.2 | |
+| 57 | 4 | Zaborowski [191] | BAO+ | Clust.+ BAO+ Lym.- | CMB–NSH | 0.1-4.2 | |
+| 58 | 4 | Krolewski (2025) [93] | DESI DR1 + CMB | Energy densities | CMB–NSH | 0.1–2.33 | |
 Notes: CMB–NSH = CMB-based, No Sound Horizon; MLE = Maximum Likelihood Estimator; PRS = Persistent Radio Source; RM = Reverberation Mapping; BGP = Binary Gravitational-wave Population; HFF = Hubble Frontier Fields; BNS = Binary Neutron Star; NSBH = Neutron Star–Black Hole.
+† fixed.
 Includes systematic uncertainties not fully accounted for in the original analysis; see Appendix B.
@@
 Rung 1 (Geometric Anchors): Three independent anchors establish the absolute distance scale:
-Gaia EDR3 parallaxes to Milky Way Cepheids (corrected for zero-point offset)
-Megamaser geometric distance to NGC 4258 [Reid2019]
-Detached eclipsing binary distances to the LMC [Pietrzynski2019]
 Rung 2 (Cepheid Calibration): Cepheid variables observed with HST WFC3 in the F160W (H-band) filter provide period–luminosity relations in 37+ galaxies hosting Type Ia supernovae.
@@
 Rung 3 (Hubble Flow SNe Ia): The calibrated SNe Ia absolute magnitude is applied to the Pantheon+ sample in the redshift range .
-The 2022 HST result [Riess2022] yielded km s-1 Mpc-1 (1.4% precision).
-The 2025 JWST result [Riess2025] yields km s-1 Mpc-1 (1.3% precision), confirming the HST photometry and ruling out crowding-induced biases as the source of the tension.
+The 2022 HST result [150] yielded km s-1 Mpc-1 (1.4% precision).
+The 2025 JWST result [149] yields km s-1 Mpc-1 (1.3% precision), confirming the HST photometry and ruling out crowding-induced biases as the source of the tension.
 Systematic Error Budget: The SH0ES analyses include comprehensive systematic assessments for Cepheid photometry ( km s-1 Mpc-1), metallicity effects ( km s-1 Mpc-1), crowding corrections ( km s-1 Mpc-1), SNe Ia standardization ( km s-1 Mpc-1), and peculiar velocities ( km s-1 Mpc-1).
@@
 The TRGB method exploits the helium flash that terminates red giant branch evolution at a nearly constant luminosity (), producing a sharp discontinuity detectable via edge-detection algorithms (Sobel filter).
-The 2021 analysis [Freedman2021] yielded km s-1 Mpc-1, intermediate between Planck and SH0ES.
+The 2021 analysis [63] yielded km s-1 Mpc-1, intermediate between Planck and SH0ES.
 This sparked debate about whether TRGB measurements genuinely favor lower or whether methodological differences explain the offset from SH0ES.
-The 2024 JWST analysis [Freedman2024] found km s-1 Mpc-1 from Cepheids, now more consistent with SH0ES. The CCHP continues to analyze TRGB and JAGB calibrations from JWST data.
+The 2024 JWST analysis [62] found km s-1 Mpc-1 from Cepheids, now more consistent with SH0ES. The CCHP continues to analyze TRGB and JAGB calibrations from JWST data.
 B.1.3 Lee et al. (2024): The Low- JAGB Outlier
-The Lee et al. [Lee2024] JAGB analysis yields km s-1 Mpc-1, consistent with Planck and below other distance ladder measurements. This represents the most significant outlier within Category 1.
+The Lee et al. [96] JAGB analysis yields km s-1 Mpc-1, consistent with Planck and below other Distance Ladder measurements. This represents the most significant outlier within Category 1.
 The JAGB method uses carbon stars on the Asymptotic Giant Branch, which exhibit a narrow luminosity distribution in the J-band.
 The discrepancy with Li et al.
-[Li2024b] () appears to originate from differences in:
+[98] () appears to originate from differences in:
 JAGB zero-point calibration (LMC anchor treatment)
@@
 The measurement by Wagner et al.
-[wagner2025] (2025) reports km s-1 Mpc-1 using the local expansion parameter , based on the infall velocity relation .
+[183] (2025) reports km s-1 Mpc-1 using the local expansion parameter , based on the infall velocity relation .
 However, the choice of appears somewhat arbitrary; standard models of the local Hubble flow typically adopt .
 Since , adopting —a value intermediate between the theoretical expectation and the assumed value—would yield km s-1 Mpc-1, shifting the result toward the Planck value and closer to the Distance Ladder mean value.
-The systematic uncertainty in propagates directly into but this source of error was not included in the original analysis [wagner2025].
+The systematic uncertainty in propagates directly into but this source of error was not included in the original analysis [183].
 However, we do include this effect in our analysis and the data shown in Table 10.
@@
 Zhang et al.
-(2024) [Zhang2024] measured the Hubble constant using galaxy pairwise peculiar velocities from the Cosmicflows-4 catalog, obtaining km s-1 Mpc-1 (statistical).
-However, as discussed in [Zhang2024], several systematic uncertainties are not fully accounted for in the reported error budget.
+(2024) [193] measured the Hubble constant using galaxy pairwise peculiar velocities from the Cosmicflows-4 catalog, obtaining km s-1 Mpc-1 (statistical).
+However, as discussed in [193], several systematic uncertainties are not fully accounted for in the reported error budget.
 The distance calibration of CF-4 galaxies introduces uncertainties of 2–3%, translating to 1.5–2.3 km s-1 Mpc-1 in .
@@
 This corrected level of uncertainty is used in our analysis.
-B.2 Key One-Step Measurements
+B.2 Key DLI–SHF Measurements
 B.2.1 TDCOSMO: Evolution of Strong Lensing Constraints
@@
 TDCOSMO IV (Birrer et al.
-2020): [Birrer2020] Incorporating stellar kinematics to constrain the mass-sheet degeneracy (MSD) shifted the result to km s-1 Mpc-1—now consistent with Planck.
+2020): [21] Incorporating stellar kinematics to constrain the mass-sheet degeneracy (MSD) shifted the result to km s-1 Mpc-1—now consistent with Planck.
 This dramatic shift demonstrated that power-law profiles alone cannot break the MSD.
-Recent results (2023–2025): Subsequent analyses using hierarchical modeling, additional lens systems, and improved kinematics yield in the range – km s-1 Mpc-1, depending on modeling assumptions [Shajib2023, Birrer2025].
+Recent results (2023–2025): Subsequent analyses using hierarchical modeling, additional lens systems, and improved kinematics yield in the range – km s-1 Mpc-1, depending on modeling assumptions [163, 175].
 We exclude the original Millon et al. (2020) result from our compilation because its methodology is now recognized as insufficient for robust inference.
@@
 measurement from our sample, as the more recent Barua et al.
-analysis uses the same underlying data while addressing the known systematic uncertainties of the earlier measurement
-B.2.2 Kelly et al. (2023): SN Refsdal Time Delays
-Supernova Refsdal, the first multiply-imaged supernova [Kelly2015], provides unique advantages for time-delay cosmography: the supernova light curve allows unambiguous identification of multiple images, unlike quasar variability.
-The Kelly et al. [Kelly2023] analysis uses time delays from the “Einstein Cross” configuration plus the predicted reappearance of image “SX.” The result, km s-1 Mpc-1, is consistent with Planck.
-Follow-up analyses with different cluster mass models yield consistent results: Li et al.
-[Li2024a] find km s-1 Mpc-1.
-The consistency across mass models provides confidence in the robustness of this constraint.
-B.2.3 Piratova-Moreno et al.:FRBs
+analysis uses the same underlying data while addressing the known systematic uncertainties of the earlier measurement.
+B.2.2 Liu et al. (2023): SDSS J1004+4112 Cluster Lens
+The Liu et al.
+[104] analysis of the cluster-lensed quasar SDSS J1004+4112 explores 16 different mass models, yielding values spanning 55–90 km s-1 Mpc-1.
+Their combined constraint with equal weighting gives km s-1 Mpc-1.
+In our analysis, we adopt their more constrained result of km s-1 Mpc-1, which corresponds to their selection of the two mass models (m113 and m123) that best reproduce the observed shapes of the lensed quasar host galaxy and background galaxies (see their Figure 5 and Section V).
+The authors themselves emphasize this as a more physically motivated constraint, stating that these models “best match the observed shapes.”
+We acknowledge that this choice favors the lower value.
+However, selecting mass models based on their ability to reproduce independent observables (lensed galaxy morphologies) is more principled than equal weighting of all models regardless of their physical plausibility.
+Readers should note that the conservative combined result ( km s-1 Mpc-1) remains consistent with our conclusions given its large uncertainty.
+B.2.3 Kelly et al. (2023) and Li et al. (2024): SN Refsdal Time Delays
+Supernova Refsdal, the first multiply-imaged supernova [86], provides unique advantages for time-delay cosmography: the supernova light curve allows unambiguous identification of multiple images, unlike quasar variability.
+Both Kelly et al. [87] and Li et al. [101] analyze time delays from SN Refsdal, yet we classify them in different categories based on their methodological approaches:
+Kelly et al.
+(2023) — Category 2: This analysis uses time delays from the “Einstein Cross” configuration plus the predicted reappearance of image “SX.” The result, km s-1 Mpc-1, is derived using specific cluster mass models that assume CDM cosmology for the time-delay distance calculation and interprets results within the CDM framework.
+Li et al.
+(2024) — Category 3: This analysis explicitly develops a “cosmological-model-independent” approach, as emphasized in their title and methodology.
+They obtain km s-1 Mpc-1 using a combination of strong lensing and other probes to determine without assuming CDM.
+The distinction hinges on whether the original analysis explicitly states model-independence.
+While both measurements yield consistent values and analyze the same underlying data, their category assignments follow our classification criterion based on the authors’ stated assumptions.
+This example illustrates that identical data can yield measurements in different categories depending on the analysis methodology employed.
+B.2.4 Piratova-Moreno et al.:FRBs
 Piratova-Moreno et al.
-(2025) [PiratovMoreno2025] employed Fast Radio Bursts (FRBs) as cosmological probes to measure the Hubble constant using a catalog of 98 localized FRBs.
+(2025) [140] employed Fast Radio Bursts (FRBs) as cosmological probes to measure the Hubble constant using a catalog of 98 localized FRBs.
 Using the Maximum Likelihood Estimate (MLE) method with confirmed FRBs, they obtained km s-1 Mpc-1 (statistical), while their mock FRB catalog yielded km s-1 Mpc-1.
@@
 This is the value used in our analysis.
-B.2.4 Vogl et al. (2024): Tailored EPM
-The Vogl et al. [Vogl2024] tailored Expanding Photosphere Method represents the most precise model-independent measurement supporting the high- solution outside the traditional distance ladder.
+B.2.5 Vogl et al. (2024): Tailored EPM
+The Vogl et al.
+[182] tailored Expanding Photosphere Method represents the most precise cosmological-model-independent measurement supporting the high- solution outside the traditional distance ladder.
 Method: EPM determines distances geometrically by combining photospheric angular size (from flux and temperature) with expansion velocity (from Doppler shifts).
@@
 Caveats: The method relies on supernova atmosphere modeling; current sample size is limited; further validation with independent codes is needed.
-B.2.5 Barua et al. (2025): Megamaser Cosmology Project
-The Megamaser Cosmology Project measures geometric distances to galaxies with circumnuclear water maser disks. The Barua et al. [Barua2025] frequentist analysis yields km s-1 Mpc-1.
+B.2.6 Barua et al. (2025): Megamaser Cosmology Project
+The Megamaser Cosmology Project measures geometric distances to galaxies with circumnuclear water maser disks. The Barua et al. [17] frequentist analysis yields km s-1 Mpc-1.
 Method: VLBI observations resolve the maser disk structure, providing angular diameter distances via Keplerian dynamics. The method is independent of stellar physics.
@@
 Evolution: Earlier measurements (Kuo 2013: ; Gao 2015: ) favored lower values with larger uncertainties.
-B.2.6 Pogosian et al. (2024): Sound-Horizon-Free BAO
-The Pogosian et al. [Pogosian2024] analysis demonstrates that can be constrained from BAO data without assuming the CMB Sound Horizon.
+B.2.7 Pogosian et al. (2024): Sound-Horizon-Free BAO
+The Pogosian et al. [144] analysis demonstrates that can be constrained from BAO data without assuming the CMB Sound Horizon.
 Method: Instead of using as a standard ruler, this analysis treats as a free parameter and constrains through:
@@
 The matter-radiation equality scale
-Results: DESI BAO yields km s-1 Mpc-1; pre-DESI BAO yields km s-1 Mpc-1.
-Significance: These measurements anchor the low- solution in Category 4. Their high precision (%) means they dominate the weighted mean of this category, contributing to the very low .
-B.2.7 Gravitational Wave standard sirens
-GW standard sirens provide constraints independent of electromagnetic distance ladders:
-Bright sirens (GW170817): The binary neutron star merger with optical counterpart yields km s-1 Mpc-1 [Palmese2023], limited by the peculiar velocity of NGC 4993 and viewing angle degeneracy.
-Dark sirens (GWTC-3): Statistical analysis of 47 events without counterparts yields km s-1 Mpc-1 [Abbott2021], consistent with Planck.
+Results: DESI BAO DR2 yields ; pre-DESI BAO yields .
+The DESI DR2 result was presented by L.
+Pogosian during the CosmoVerse Seminar Series [145] and represents an update of [144] using DESI DR1 BAO data; publication is forthcoming.
+Significance: These measurements anchor the low- solution in Category 4. Their high precision (%) means they dominate the weighted mean of this category.
+B.2.8 Gravitational Wave standard sirens
+GW standard sirens provide constraints independent of electromagnetic distance ladder:
+Bright sirens (GW170817): The binary neutron star merger with optical counterpart yields km s-1 Mpc-1 [122], limited by the peculiar velocity of NGC 4993 and viewing angle degeneracy.
+Additionally, Wang et al.
+(2025) [185] find a consistent value of km s-1 Mpc-1.
+These measurements provide important complementary constraints that are independent of the Distance Ladder.
+Dark sirens (GWTC-3): Statistical analysis of 47 events without counterparts yields km s-1 Mpc-1 [102], consistent with Planck.
 The weighted mean of all GW measurements ( km s-1 Mpc-1) supports the low- solution, though with larger uncertainties than other probes.
-B.2.8 Fast Radio Burst Measurements
+B.2.9 Fast Radio Burst Measurements
 FRB measurements use the dispersion measure–redshift relation to probe the cosmic baryon distribution:
@@
 Sample selection effects
-Results: Published values span – km s-1 Mpc-1. Maximum-likelihood approaches [PiratovMoreno2025] tend to favor lower values ( km s-1 Mpc-1).
-As FRB samples grow and host DM distributions become better characterized, this probe may provide competitive constraints.
+Galactic electron density models
+Results: Published values span – km s-1 Mpc-1.
+Maximum-likelihood approaches [140] tend to favor lower values ( km s-1 Mpc-1).
+Also the estimation of using Fast Radio Bursts (FRBs) is highly sensitive to the model adopted for the Galactic electron density distribution.
+As highlighted by recent studies (e.g., Wang et al.
+[184]), switching between established models (such as NE2001 and YMW16) can shift the inferred value significantly (e.g., from to km s-1 Mpc-1), introducing a systematic uncertainty that rivals the statistical error.As FRB samples grow and host DM distributions become better characterized, this probe may provide competitive constraints.
 Appendix C Correlation Matrix Estimates
@@
 Geometric anchors: Most Cepheid measurements use NGC 4258, LMC, and/or MW parallaxes. The anchor uncertainties propagate coherently.
-SNe Ia samples: Hubble-flow SNe Ia are predominantly from Pantheon+ [Brout2022], with substantial overlap. standardization parameters (, , ) are often shared.
+SNe Ia samples: Hubble-flow SNe Ia are predominantly from Pantheon+ [25], with substantial overlap. standardization parameters (, , ) are often shared.
 Period–luminosity relations: Cepheid analyses adopt similar functional forms and metallicity corrections.
@@
 We estimate an effective correlation coefficient – for typical pairs of Distance Ladder measurements.
-C.1.2 One-Step Correlations
+C.1.2 DLI–SHF Correlations
 Categories 2, 3, and 4 exhibit weaker correlations due to methodological diversity, but some correlations exist:
@@
 Cosmological priors: Many analyses assume flat CDM with , inducing weak correlations.
-We estimate – for the One-Step sample.
+We estimate – for the sample.
 C.2 Methodology for Correlation-Adjusted Estimates
 For correlated measurements, the generalized weighted mean and its uncertainty are:
-where is the covariance matrix with elements .
-For simplicity, we adopt a uniform correlation structure within each category:
+where is the covariance matrix with elements . For simplicity, we adopt a uniform correlation structure within each category:
 C.3 Impact on Combined Constraints
@@
 With :
-compared to km s-1 Mpc-1 under independence.
-The central value shifts by only km s-1 Mpc-1 (within statistical uncertainty), but the uncertainty increases by %.
-C.3.2 Combined One-Step (Categories 2+3+4)
+compared to km s-1 Mpc-1 under independence. The central value shifts by only km s-1 Mpc-1 (within statistical uncertainty), but the uncertainty increases by %.
+C.3.2 Combined (Categories 2+3+4)
 With :
-compared to km s-1 Mpc-1 under independence.
-The central value shifts by km s-1 Mpc-1 , and the uncertainty increases by %.
+compared to km s-1 Mpc-1 under independence. The central value remains essentially unchanged, and the uncertainty increases by %.
 C.3.3 Tension Between Categories
@@
 | Tension | ||||
-| 0.0 | 0.0 | 0.40 | 0.46 | |
-| 0.2 | 0.1 | 0.61 | 0.60 | |
-| 0.4 | 0.15 | 0.58 | 0.60 | |
-| 0.5 | 0.2 | 0.54 | 0.59 | |
-| 0.6 | 0.25 | 0.49 | 0.58 |
+| 0.0 | 0.00 | 0.39 | 0.34 | |
+| 0.2 | 0.10 | 0.60 | 0.42 | |
+| 0.4 | 0.15 | 0.57 | 0.42 | |
+| 0.5 | 0.20 | 0.53 | 0.41 | |
+| 0.6 | 0.25 | 0.48 | 0.40 |
 Notes: Uncertainties in km s-1 Mpc-1. Central values are relatively stable across correlation assumptions.